### v1

# On the minimax optimality of Flow Matching through the connection to kernel density estimation

On the minimax optimality of Flow Matching through the connection to kernel density estimation
Lea Kunkel and Mathias Trabs [^1] We thank the participants of the Oberwolfach Mini-Workshop 2508b ”Statistical Challenges for Deep Generative Models” for inspiring discussions and helpful comments on this project.
(Karlsruhe Insitute of Technology)
Flow Matching has recently gained attention in generative modeling as a simple and flexible alternative to diffusion models, the current state of the art. While existing statistical guarantees adapt tools from the analysis of diffusion models, we take a different perspective by connecting Flow Matching to kernel density estimation. We first verify that the kernel density estimator matches the optimal rate of convergence in Wasserstein distance up to logarithmic factors, improving existing bounds for the Gaussian kernel. Based on this result, we prove that for sufficiently large networks, Flow Matching also achieves the optimal rate up to logarithmic factors, providing a theoretical foundation for the empirical success of this method. Finally, we provide a first justification of Flow Matching’s effectiveness in high-dimensional settings by showing that rates improve when the target distribution lies on a lower-dimensional linear subspace.
Keywords: Generative models, continuous normalizing flows, rate of convergence, kernel density estimator, Wasserstein distance, distribution estimation
MSC 2020: 62E17, 62G07, 68T07
1 Introduction
The goal of generative models is to learn a map $\psi$ that transforms a random variable $Z$ drawn from a simple and known latent distribution $\mathbb{U}$ into a new variable $\psi(Z)$ whose distribution closely approximates the unknown target distribution $\mathbb{P}^{*}$. Hence, generative models are naturally linked to the classical non-parametric distribution estimation problem, enhanced by the additional requirement of enabling fast and efficient sampling from the learned distribution.
Flow Matching, a generative model introduced by Lipman et al. [2023], has recently attained significant interest due to its considerably more straightforward construction in comparison to diffusion models (building up on Sohl-Dickstein et al. [2015]), which have been regarded as the state-of-the-art generative method. Instead of learning $\psi$ directly, like GANs (Goodfellow et al. [2014]), Flow Matching relies on learning a vector field $v$ that describes the direction of movement to transfer the probability mass from $\mathbb{U}$ to $\mathbb{P}^{*}$. This leads to the following ODE in time $t\in[0,1]$,
$\frac{d}{dt}\psi_{t}(x)=v_{t}(\psi_{t}(x)),\quad\psi_{0}(x)=x.$
The solution to this ODE yields a full path of probability densities via the push-forward distributions
$p_{t}=[\psi_{t}]_{\#}p_{0},$
where $p_{0}$ is the density of $\mathbb{U}$. The vector field $v$ should be chosen such that $p_{1}$ is close to the density of $\mathbb{P}^{*}$.
Originally, normalizing flows were used to transform a latent distribution into an approximation of an unknown distribution by an invertible map constructed according to the change of variable theorem. This method was defined by Tabak and Vanden-Eijnden [2010] and Tabak and Turner [2013], but popularized by Rezende and Mohamed [2015] and Dinh et al. [2015]. For an overview see Kobyzev et al. [2020]. Continuous normalizing flows (CNF, Chen et al. [2018]) use a neural ordinary differential equation (ODE) to construct this mapping via a maximum likelihood method. However, the training requires simulations of the ODE, which results in high computational cost, see Grathwohl et al. [2019]. The Flow Matching approach of Lipman et al. [2023] avoids this by using a regression objective to train a CNF, which avoids simulations during training,
$\Psi(\tilde{v})\coloneqq\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}\big{[}\left|\tilde{v}_{t}(X_{t})-v_{t}\left(X_{t}|Y\right)\right|^{2}\big{]},$ (1)
where $p_{t}(\cdot|Y)$ is a time dependent conditional density path associated to the conditional vector field $v_{t}(\cdot|Y)$ and both $p_{t}(\cdot|Y)$ and $v_{t}(\cdot|Y)$ are known by construction. This regression objective is in certain cases similar to the score matching objective (Hyvärinen [2005], Vincent [2011]), which is used to train diffusion models based on stochastic differential equations, introduced by Song et al. [2020].
Flow Matching algorithms have been successfully used in many different applications that benefit from efficient sampling, such as text-to-speech (Guo et al. [2024]) and text-to-image (Yang et al. [2024], Esser et al. [2024]) settings, the production of novel molecular or protein structures (Dunn and Koes [2024], Bose et al. [2024]) or the construction of surrogate models in high energy physics (Bieringer et al. [2024]). It has also been adjusted theoretically to different settings. Atanackovic et al. [2024] adapted Flow Matching to the case of interacting particles, Gat et al. [2024] explored the discrete setting. Chen and Lipman [2024] generalized the Euclidean setting to the Riemannian setting, allowing for more general geometries. Kerrigan et al. [2023] extended Flow Matching to function spaces. However, the statistical properties have only recently been studied by Fukumizu et al. [2024] in the Wasserstein $2$ distance. They do not use the exact setting of Lipman et al. [2023], but rather introduce stopping times that depend on the number of samples, enabling the transfer of methods known from the statistical analysis of diffusion models.
In this work, we demonstrate that Flow Matching in the setting of Lipman et al. [2023] is closely related to the classical kernel density estimation (KDE). This connection allows us to analyze Flow Matching from a new perspective. First, we show that the motivation of Flow Matching also holds for its empirical counterparts where we replace the unknown target density $p^{*}$ in (1) with the empirical measure of the observations. For sufficiently large network classes, the resulting generative algorithm coincides exactly with a kernel density estimator, where the kernel is given by the density of the latent distribution. We show convergence rates for the kernel density estimator in Wasserstein-$1$ distance in a setting that allows for the Gaussian kernel which is the most popular choice in Flow Matching applications. Our rates coincide with the minimax optimal rates from Niles-Weed and Berthet [2022] up to logarithmic factors. We then transfer our analysis to Flow Matching using neural network classes of parameterized vector fields.
Unlike Fukumizu et al. [2024], we do not build on the similarities to diffusion models. Separating the error of the kernel density estimator allows us to use empirical risk minimization without the need for Chernoff-type bounds. this avoids the problems pointed out by Yakovlev and Puchkin [2025] in the analysis of diffusion models. Overall, the analysis of Flow Matching is more delicate, since Girsanov’s theorem does not apply and thus the strategy of Chen et al. [2023c] cannot be used. Therefore, our bound depends exponentially on the Lipschitz constant of the vector field. This is one of the reasons for the difficulties in the proof of Fukumizu et al. [2024]. This analysis allows us to show convergence rates in Wasserstein-$1$ distance in the case where finite neural networks are used for the vector field. While these rates are minimax optimal up to a logarithmic constant, they suffer from the curse of dimensionality. In case the support of $\mathbb{P}^{*}$ is concentrated on a linear subspace, we improve our results in a way that the convergence rate depends only on the intrinsic dimension. This provides a first justification for the excellent performance of Flow Matching for high-dimensional data sets.
Related work
Lipman et al. [2023] use a fixed latent distribution. Similar approaches for flows between two possibly unknown distributions $\mathbb{P}$ and $\mathbb{Q}$ are studied by Tong et al. [2024], Liu et al. [2022] and Albergo and Vanden-Eijnden [2022]. Tong et al. [2024] also generalize the mentioned methods. Gao et al. [2024] prove a suboptimal rate of convergence for certain forms of Flow Matching not exactly conforming to the construction on Lipman et al. [2023] in the Wasserstein $2$ distance. Benton et al. [2024] analyzed Flow Matching excluding the approximation error by imposing assumptions on the covariance that lead to global Lipschitz bounds on the vector field. They also focus on different constructions than Lipman et al. [2023]. Gong et al. [2025] study the properties of ReLU networks to approximate a vector field corresponding to higher order trajectories.
As Lipman et al. [2023] point out, Flow Matching is closely related to diffusion models. For an overview of generative diffusion models, see Cao et al. [2024]. Indeed, even in cases that are not constructed to be consistent with diffusion models, the approximation of a score function has similar properties to the approximation of the Flow Matching vector field. Fukumizu et al. [2024] build up on this connection. The statistical properties of score matching are an area of ongoing research, see for example Chen et al. [2023a], Chen et al. [2023b], Chen et al. [2023c], Oko et al. [2023], Tang and Yang [2024], Azangulov et al. [2024], Zhang et al. [2024] , Yakovlev and Puchkin [2025]. Marzouk et al. [2023] study the statistical properties of CNFs trained by likelihood maximization.
The kernel density or Parzen-Rosenblatt estimator (Rosenblatt [1956], Parzen [1962]) is the classical method for estimating a smooth density. For an overview in the univariate case, see Tsybakov [2009], Chapter 1.2 or Devroye and Lugosi [2001] and in the multivariate case see Scott [1992]. Typically, the kernel density estimator is analyzed using the mean squared error or the $L_{1}$ distance. Despite its longstanding use, kernel density estimation is still the subject of ongoing research, for example in density estimation on unknown manifolds (Berry and Sauer [2017], Berenfeld and Hoffmann [2021], Divol [2022], Wu and Wu [2022]). In particular, Divol [2022] uses the Wasserstein metric to evaluate the performance of a kernel density estimator, however he needs to assume properties of the kernel that are not satisfied by standard choices such as the Gaussian kernel.
Outline
In Section 2 we introduce the problem setting mathematically and recall all the definitions needed to define the conditional Flow Matching objective from Lipman et al. [2023]. We introduce the empirical counterparts, validate the empirical conditional Flow Matching objective, and portray the connection to the kernel density estimator. In Section 3 we start by separating the error corresponding to the kernel density estimator and the vector field approximation. In Section 3 we first separate the error corresponding to the kernel density estimator and the vector field approximation. We then proceed to bound the former error in Section 3.1. Incorporating theory of network approximation, we bound the latter error and obtain a rate of convergence for the entire Flow Matching procedure in Section 3.2. Subsequently, we show in Section 3.3 that the rate can be improved if the unknown distribution is supported on a linear subspace. All proofs are deferred to Section 4.
1.1 Some notations
We write $X\sim p$ as a synonym for $X\sim\mathbb{P}$ in case $\mathbb{P}$ has a density $p$ with respect to the Lebesgue measure and $\mathbb{P}$ is the distribution of the random variable $X$. By $\lesssim$ we denote $a\leq c\cdot b,$ where $a,b\in\mathbb{R}$ and $c$ is a constant independent of $n$. For $d\in\mathbb{N}$, we use the Euclidean norm ||$\cdot$|| ⋅ | on $\mathbb{R}^{d}$. The supremum norm and the $L_{1}$-norm of a function $f\colon\Omega\rightarrow\mathbb{R}^{d^{\prime\prime}}$ with $d^{\prime},d^{\prime\prime}\in\mathbb{R}$ and $\Omega\subset\mathbb{R}^{d^{\prime}}$ is defined as
$\|f\|_{\infty,\Omega}\coloneqq\sup_{x\in\Omega}|f(x)|\quad\text{and}\quad\|f\|_{1,\Omega}\coloneqq\int_{\Omega}|f(x)|\;\mathrm{d}x.$
In case of $\Omega=\mathbb{R}^{d^{\prime}},$ we write $\|f\|_{\infty,\mathbb{R}^{d^{\prime}}}=\|f\|_{\infty}$ and $\|f\|_{1,\mathbb{R}^{d^{\prime}}}=\|f\|_{1}.$
For $d^{\prime\prime}=1$ and $\alpha\in(0,1],$ the Besov space $B_{1,\infty}^{\alpha}$ is defined as
$B_{1,\infty}^{\alpha}(\Omega)\coloneqq\big{\{}f\in L_{1}(\Omega)\colon|f|_{B_{1,\infty}^{\alpha}(\Omega)}\coloneqq\sup_{t>0}t^{-\alpha}\omega_{1}(f,t)_{1}<\infty\big{\}}$
where
$\omega_{1}(f,t)_{1}:=\sup_{0<|h|\leq t}\int|f(x)-f(x+h)|\;\mathrm{d}x,\quad t>0,x\in\Omega.$
The difference is set to zero if $x+h\notin\Omega.$ On $B_{1,\infty}^{\alpha}$ we define the norm $\|f\|_{B_{1,\infty}^{\alpha}(\Omega)}=\|f\|_{L_{1}(\Omega)}+|f|_{B_{1,\infty}^{\alpha}(\Omega)}$.
Finally, let $C^{k}(\Omega)$ denote the set of functions $f$ such that the coordinate functions are each $k$-times continuously differentiable and the supremum norm of the derivatives are finite. For $k\in\mathbb{N}_{0}^{d}$ and $|k|=\sum_{i=1}^{d}k_{i}$ set $D^{k}=\frac{\partial^{|k|}}{\partial x_{1}^{k_{1}}\ldots\partial x_{d}^{k_{d}}},$ where $\frac{\partial}{\partial x_{i}}$ denotes the weak partial derivative.
2 Empirical Flow Matching
Assume we observe an i.i.d. sample $X^{*}_{1},\dots,X^{*}_{n}$ from an unknown distribution $\mathbb{P}^{*}$ on a compact set $\mathcal{X}\subset[-1,1]^{d}.$ Further assume that $\mathbb{P}^{*}$ has a density $p^{*}$ with respect to the Lebesgue measure. The goal of generative modeling is to mimic the distribution $\mathbb{P}^{*}.$
For a time dependent vector field $v\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ we consider the flow $\psi\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ given as the solution to the ODE
$\frac{d}{dt}\psi_{t}(x)=v_{t}(\psi_{t}(x)),\quad\psi_{0}(x)=x.$ (2)
For a fixed latent distribution with Lebesgue density $p_{0}$, the vector field $v$ generates a probability density path $p\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{>0}$ with $\int p_{t}(x)\mathrm{d}x=1$ for all $t$ via the push-forward distributions
$p_{t}=[\psi_{t}]_{\#}p_{0},\quad\text{i.e. }\psi_{t}(Z)\sim p_{t}\text{ for }Z\sim p_{0}.$
The ODE (2) corresponds to the Lagrangian description (in terms of particle trajectories) of the conservation of mass formula [Villani, 2008, p.14]. The change of variables formula links it to the Eulerian description: A necessary and sufficient condition for $v_{t}$ to generate $p_{t}$ is
$\frac{d}{dt}p_{t}+\operatorname{div}(p_{t}v_{t})=0,$ (3)
see [Villani, 2008, p. 14].
Approximating a given vector field $v$ that generates a certain density path $p_{t}$ using a function $\tilde{v}$ leads to the Flow Matching objective
$\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1]\\ X_{t}\sim p_{t}\end{subarray}}\big{[}|v_{t}(X_{t})-\tilde{v}_{t}(X_{t})|^{2}\big{]}$ (4)
by Lipman et al. [2023]. They have consider a probability path of the form
$p_{t}(x)=\int p_{t}(x|y)p^{*}(y)\;\mathrm{d}y,$ (5)
where $p_{t}(\cdot|y)\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a conditional probability path generated by some vector field $v_{t}(\cdot|y)\colon\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ for $y\in\mathbb{R}^{d}$. The vector field generating (5) is then given by
[…]
In this setting Lipman et al. [2023] show that the Flow Matching objective (4) is equivalent to the conditional Flow Matching objective
$\Psi(\tilde{v})\coloneqq\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}\big{[}\left|\tilde{v}_{t}(X_{t})-v_{t}\left(X_{t}|Y\right)\right|^{2}\big{]}.$ (7)
As $p^{*}$ is unknown, in practice, the expectation in $Y\sim p^{*}$ is replaced by the empirical counterpart based on the observations $X_{1}^{*},\dots,X_{n}^{*}$. This leads to the empirical counterparts of (5) and (6) given by
$p^{n}_{t}(x)$ $=\frac{1}{n}\sum_{i=1}^{n}p_{t}(x|X_{i}^{*})\qquad\text{and}\qquad v^{n}_{t}(x)=\sum_{i=1}^{n}v_{t}(x|X_{i}^{*})\frac{p_{t}(x|X_{i}^{*})}{\sum_{j=1}^{n}p_{t}(x|X_{j}^{*})}.$ (8)
With this modification we recover the sufficient condition for $v^{n}_{t}$ to generate $p^{n}_{t}$ analogously to [Lipman et al., 2023, Theorem 1].
Lemma 2.1.
If $v_{t}(\cdot|X_{i}^{*})$ generates $p_{t}(\cdot|X_{i}^{*})$ for all $i=1,\dots,n,$ then $v^{n}_{t}$ generates $p^{n}_{t}.$
The motivation for the conditional Flow Matching objective (7) is the equivalence to the unconditioned Flow Matching objective with respect to the optimizing arguments [Lipman et al., 2023, Theorem 2]. Using the empirical counterparts, this still holds true.
Theorem 2.2.
Using $p^{n}_{t}$ and $v^{n}_{t}$ from (8), and a class of parameterized vector fields $\mathcal{M},$ which is constructed such that the minimal arguments exists, we have that
$\underset{\hat{v}\in\mathcal{M}}{\mathrm{argmin}}\int_{0}^{1}\mathbb{E}_{X_{t}\sim p^{n}_{t}}\big{[}|\hat{v}_{t}(X_{t})-v^{n}_{t}(X_{t})|^{2}\big{]}\;\mathrm{d}t=\underset{\hat{v}\in\mathcal{M}}{\mathrm{argmin}}\int_{0}^{1}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\tilde{X}_{t}\sim p_{t}(\cdot|X_{i}^{*})}\big{[}|\hat{v}_{t}(\tilde{X}_{t})-v_{t}(\tilde{X}_{t}|X_{i}^{*})|^{2}\big{]}\;\mathrm{d}t.$
Hence the empirical conditional Flow Matching objective
$\tilde{\Psi}(\tilde{v})\coloneqq\int_{0}^{1}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\tilde{X}_{t}\sim p_{t}(\cdot|X_{i}^{*})}\big{[}|\tilde{v}_{t}(\tilde{X}_{t})-v_{t}(\tilde{X}_{t}|X_{i}^{*})|^{2}\big{]}\;\mathrm{d}t$ (9)
is justified theoretically. Compared to Gao et al. [2024], we use $v_{t}(\cdot|X_{i}^{*})$ directly as proposed by Lipman et al. [2023] and stick to the entire time interval. Note that for simplicity we do not sample $t.$
From minimizing (9) in $\tilde{v}\in\mathcal{M},$ we obtain an optimal argument $\hat{v}.$ Solving the ODE (2) using $\hat{v}$, we obtain a flow $\hat{\psi}_{t}$, i.e. $\hat{\psi}$ is given by
$\frac{d}{dt}\hat{\psi}_{t}(x)=\hat{v}_{t}(\hat{\psi}_{t}(x)),\quad\hat{\psi}_{0}(x)=x,\qquad\text{for}\quad\hat{v}\in\operatorname*{argmin}_{\tilde{v}\in\mathcal{M}}\tilde{\Psi}(\tilde{v}).$ (10)
We use this flow to push forward the known, latent distribution to time $t=1$. In accordance with the goal of generative modeling, the distribution of this pushforward should mimic $\mathbb{P}^{*}.$
In order to apply Flow Matching, we have to construct a class of conditional probability paths. Let $Z\sim\mathbb{U}$ and let $K$ denote the density of $\mathbb{U}$. Consider the functions $\sigma\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{>0}$ and $\mu\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}.$ Set for $t\in[0,1]$ and a given $X_{i}^{*},i\in\{1,..,n\}$,
$\psi_{t}(Z|X_{i}^{*})\coloneqq\sigma_{t}(X_{i}^{*})Z+\mu_{t}(X_{i}^{*}).$ (11)
The density of $\psi_{t}(Z|X_{i}^{*})$ is by the transformation formula
$p_{t}(x|X_{i}^{*})=\frac{1}{\sigma^{d}_{t}(X_{i}^{*})}K\Big{(}\frac{x-\mu_{t}(X_{i}^{*})}{\sigma_{t}(X_{i}^{*})}\Big{)}.$
We call $p_{t}(\cdot|X_{i}^{*})$ the conditional kernel probability path. Setting
$\frac{d}{dt}\psi_{t}(x|X_{i}^{*})=v_{t}(\psi_{t}(x|X_{i}^{*})|X_{i}^{*}),$ (12)
we recover the same result like in [Lipman et al., 2023, Theorem 3]:
Lemma 2.3.
Let $p_{t}\left(x|X_{i}^{*}\right)$ be a conditional kernel probability path, and $\psi_{t}(\cdot|X_{i}^{*})$ its corresponding flow as in (11). Then, the unique vector field that defines $\psi_{t}(\cdot|X_{i}^{*})$ via (12) has the form:
$v_{t}\left(x|X_{i}^{*}\right)=\frac{\frac{\partial\sigma_{t}}{\partial t}\left(X_{i}^{*}\right)}{\sigma_{t}\left(X_{i}^{*}\right)}\big{(}x-\mu_{t}\left(X_{i}^{*}\right)\big{)}+\frac{\partial\mu_{t}}{\partial t}\left(X_{i}^{*}\right).$
Set $\sigma_{\min}>0.$ To flow from $p^{n}_{0}(x)=\frac{1}{n}\sum_{i=1}^{n}K(x)=K(x)$ to $p^{n}_{1}(x)=\frac{1}{n\sigma^{d}_{\mathrm{min}}}\sum_{i=1}^{n}K(\frac{x-X_{i}^{*}}{\sigma_{\mathrm{min}}}),$ we can choose any differentiable functions $\sigma_{t}$ and $\mu_{t}$ such that
$\mu_{0}=0,\quad\mu_{1}=1\quad\text{and}\quad\sigma_{0}=1,\quad\sigma_{1}=\sigma_{\min}.$
At time $t=1$ the distribution $\mathbb{P}^{\psi_{1}^{n}(Z)}$ then coincides with the kernel density estimator
$p^{n}_{1}(x)=\frac{1}{n\sigma^{d}_{\mathrm{min}}}\sum_{i=1}^{n}K\Big{(}\frac{x-X_{i}^{*}}{\sigma_{\mathrm{min}}}\Big{)},$ (13)
where the kernel is given by the latent distribution $\mathbb{U}$. Choosing $\mathbb{U}=\mathcal{N}_{d}(0,1)$, i.e., we consider the $d$-dimensional Gaussian kernel $K(x)=(2\pi)^{-d/2}\exp(-|x|^{2}/2)$, yields the proposed flow from Lipman et al. [2023, Section 4]. Moreover, considering general kernels is in line with methods by Tong et al. [2024], Liu et al. [2022] and Albergo and Vanden-Eijnden [2022], which transform an unknown distribution to another.
3 Rate of convergence
The aim of this work is to evaluate how well Flow Matching performs depending on the number of observations $n.$ To this end, we have to control the distance between $\mathbb{P}^{*}$ and $\mathbb{P}^{\hat{\psi}_{1}(Z)}$ with flow $\hat{\psi}$ from (10). As as evaluation metric, we use the Wasserstein-$1$ metric.
On the normed vector space $(\mathbb{R}^{d},|\cdot|)$ the Wasserstein-$1$ distance between two probability distributions $\mathbb{P}$ and $\mathbb{Q}$ on $\mathbb{R}^{d}$ is defined as
$\mathsf{W}_{1}(\mathbb{P},\mathbb{Q})$ $\coloneqq\Big{(}\inf_{\pi\in\Pi(\mathbb{P},\mathbb{Q})}\int_{\mathcal{X}}|x-y|\;\mathrm{d}\pi(x,y)\Big{)},$
where $\Pi$ is the set of all distribution whose marginal distributions correspond to $\mathbb{P}$ and $\mathbb{Q}$ respectively. By [Villani, 2008, Theorem 5.10(i)], the following duality holds:
$\mathsf{W}_{1}(\mathbb{P},\mathbb{Q})=\sup_{W\in\operatorname{Lip}(1)}\mathbb{E}_{\begin{subarray}{c}X\sim\mathbb{P}\\ Y\sim\mathbb{Q}\end{subarray}}[W(X)-W(Z)].$ (14)
On the space of probability measures with finite first moment, the Wasserstein-$1$ distance metrizes weak convergence. A comprehensive overview of the advantages of this distance over other measures that metrize weak convergence can be found in [Villani, 2008, p. 98 f.]. Due to its compatibility with dimension reduction settings such as the manifold hypothesis, the Wasserstein $1$ distance is a popular choice for the evaluation of generative models, see e.g. Schreuder et al. [2021], Liang [2021], Stéphanovitch et al. [2024] or Kunkel and Trabs [2025].
The following theorem provides a first comparison of the performance of $\mathbb{P}^{\hat{\psi}_{1}(Z)}$ and the empirical flow $\mathbb{P}^{\psi^{n}_{1}(Z)}$.
Theorem 3.1.
Assume that all functions in $\mathcal{M}$ are Lipschitz continuous for fixed $t$ with Lipschitz constant $\Gamma_{t}.$ Then for any $\tilde{v}\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d},$ with $\tilde{v}\in\mathcal{M}$ and $v^{n}$ from (8)
$\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{\psi}_{1}(Z)})\leq\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})+\sqrt{2e}e^{\int_{0}^{1}\Gamma_{t}\;\mathrm{d}t}\|v^{n}-\tilde{v}\|_{\infty}.$ (15)
The first term in (15) depends on the choice of the conditional densities $p_{t}(\cdot|X_{i}^{*})$ and on the number of observations, but is independent of the approximation class $\mathcal{M}$. The error term $\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})$ is a mixture of an approximation and a stochastic error and is determined by the choice of the flow $\psi^{n}_{1}$ or the distribution $p^{n}_{1}$, respectively. Due to the Flow Matching construction, the latter depends only on the conditional density path $p_{t}(\cdot|X_{i}^{*})$. Note that Lemma 2.3 demonstrates how $v_{t}(\cdot|X_{i}^{*})$ can be constructed based on a given $p_{t}(\cdot|X_{i}^{*})$. Using the construction from (11) to (13), $\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})$ is the estimation error of a kernel density estimator.
The second term $\|v^{n}-\hat{v}\|_{\infty}$ depends on the set $\mathcal{M}$ and its ability to approximate $v^{n}$. For a sufficiently rich class $\mathcal{M}$ this second term will be negligible in our analysis.
The proof of Theorem 3.1 uses Grönwall’s Lemma, which leads to the factor $e^{\int_{0}^{1}\Gamma_{t}\;\mathrm{d}t}$ in the second term in (15). This is standard in the analysis of ODEs, in context of flow-based generative models see Fukumizu et al. [2024] and Albergo and Vanden-Eijnden [2022]. Compared to the SDE setting of diffusion models, there is no such result like Girsanov’s theorem, which is typically used to study these models (following Chen et al. [2023c]) and leads to the avoidance of the exponential dependence on the Lipschitz constant. This greatly complicates the analysis of Flow Matching. Given this dependence, a Lipschitz regularization in the Flow Matching objective seems to be theoretically beneficial. This could be a possible avenue for future research.
Remark 3.2.
A bound for arbitrary Lipschitz $\tilde{v}_{t}$ cannot be better than $\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})$. If $v^{n}_{t}$ is contained in $\mathcal{M},$ then the Picard-Lindelöf theorem yields $\hat{\psi}_{t}=\psi^{n}_{t}$ and hence $\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{\psi_{1}}})=\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}}).$
Due to Theorem 3.1, we can bound bothnm error terms individually. We start with the error corresponding to the kernel density estimator in Section 3.1. Subsequently, we bound the second error term for approximation classes $\mathcal{M}$ consisting of feedforward neural networks in Section 3.2 to obtain our final result.
3.1 Rate of convergence of the KDE in Wasserstein distance
To bound $\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi_{1}^{n}(Z)})$ in Theorem 3.1, we have to analyze the corresponding kernel density estimator in the Wasserstein-$1$ distance. For compact kernels, existing results for the $L_{1}$-error of the KDE can be exploited. However, this can only lead to the suboptimal convergence rate $n^{-\frac{\alpha}{2\alpha+d}}$. Indeed, the Lipschitz regularity of the test function in (14) can be exploited to obtain a faster rate. Additionally, to allow for latent distributions with unbounded support, especially the Gaussian distribution used by Lipman et al. [2023], we need to investigate kernels with unbounded support.
Theorem 3.3.
Assume $d\geq 2,p^{*}\in B_{1,\infty}^{\alpha},\alpha\in(0,1]$ with $\|p^{*}\|_{B_{1,\infty}^{\alpha}}\leq M$ and $\operatorname{supp}(p^{*})$ bounded. Assume $K$, used in the Flow Matching setting via (13), is a nonnegative $d$-dimensional kernel such that
$\int yK(y)\;\mathrm{d}y=0,\quad\text{and}\quad\int|y|^{1+\alpha}K(y)\;\mathrm{d}y<\infty.$
Further assume $K\in C^{\frac{d+2}{2}}$ with $\|D^{k}K\|_{1}\leq C$ for $k\in\mathbb{N}_{0}^{d}$ such that $|k|\leq\frac{d+2}{2}$ and some $C>0$. Then there are $C_{1}$ and $C_{2}$ such that for $v^{n}$ from (8),
$\mathbb{E}\big{[}\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})\big{]}\leq C_{1}\sigma_{\min}^{1+\alpha}+\frac{C_{2}}{\sqrt{\sigma^{d-2}}}\frac{\log n}{\sqrt{n}}.$
For $\sigma_{\operatorname{min}}\sim(n/\log^{2}n)^{-\frac{1}{2\alpha+d}}$ we obtain $\mathbb{E}\big{[}\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})\big{]}=\mathcal{O}((n/\log^{2}n)^{-\frac{1+\alpha}{2\alpha+d}}).$
The convergence rate $(n/\log^{2}n)^{-\frac{1+\alpha}{2\alpha+d}}$ coincides up to the logarithmic factor with the lower bound by Niles-Weed and Berthet [2022, Theorem 3] and thus the above rate is minimax optimal up to the logarithm. Divol [2022] also considers rates of convergence for kernel density estimators. In case of $d\neq 2,$ he obtains the optimal rate for distributions bounded away from zero. However, the kernels he considers must be smooth radial functions with support bounded in $(0,1)^{d}$, [Divol, 2022, Condition A]. Hence, his result does not apply to the Gaussian kernel.
Remark 3.4.
  1. 1.
    The assumption $\|D^{k}K\|_{1}\leq C$ for $k\in\mathbb{N}_{0}^{d}$ and $C>0$ is satisfied by the Gaussian kernel.
  2. 2.
    For kernels that to not satisfy the differentiability condition in Theorem 3.3, we show in Theorem A.1 a rate of convergence of $n^{-\frac{1+\alpha}{2\alpha+2+d}}.$ For nonzero $\alpha$ and $d$ big enough, this rate decays faster than $n^{-\frac{1}{d}},$ the rate of convergence of the empirical measure (Dudley [1969], Boissard and Gouic [2014]). This result replaces the assumption that $\operatorname{supp}(p^{*})$ is bounded by a moment assumption on $p^{*}.$
  3. 3.
    The second term in the proof of Theorem 3.3 is the expected value of $\mathsf{W}_{1}(K_{\sigma_{\min}}\ast\mathbb{P}_{n},K_{\sigma_{\min}}\ast\mathbb{P}^{*}).$ Results on the convergence of the smooth empirical measure have been discussed in case of the Gaussian kernel. Goldfeld et al. [2020, Proposition 1] obtain a rate in $\mathcal{O}\big{(}\frac{1}{\sqrt{n\sigma_{\min}^{d}}}\big{)}$ for subgaussian distributions $\mathbb{P}^{*}$, which coincides with the rate obtained in Theorem A.1. Weed [2018] looks at distributions on $[-1,1]$ and obtains a rate in $\mathcal{O}\big{(}\frac{1}{\sqrt{n}\sigma_{\min}^{d}}\big{)}.$ Hence the results above improves these results for distributions in $B^{\alpha}_{1,\infty}$ on bounded support.
3.2 Flow Matching with parameterized vector field
If the empirical vector field $v^{n}$ is contained in $\mathcal{M}$, then the second error term in Theorem 3.1 vanishes and the Flow Matching procedure inherits the optimal rate of convergence from Theorem 3.3. In practice, (9) is optimized over a class of neural networks to find a good approximate of $v^{n}$. We thus have to take this approximation error into account.
To analyze the ability of a network to approximate $v^{n},$ it is necessary to know the properties of $v^{n}$ for all $t\in[0,1].$ Hence we are going to specify $\sigma$, $\mu$ and a latent distribution for the subsequent analysis.
Assumption 3.5.
  1. 1.
    We consider the following choices of $\sigma\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{>0}$ and $\mu\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}:$
    $\sigma_{t}(X_{i}^{*})=1-(1-\sigma_{\mathrm{min}})t\quad\text{and}\quad\mu_{t}(X_{i}^{*})=tX_{i}^{*}.$
  2. 2.
    We choose $\mathbb{U}=\mathcal{N}(0,I_{d})$ for the latent distribution.
The choice of $\sigma_{t}$ and $\mu_{t}$ in 3.5 coincides with the setting of Lipman et al. [2023] in Example II. This is not a unique choice, for other choices used in the literature in similar settings, see [Tong et al., 2024, Table 1]. To avoid losing information, if $\mathbb{U}=\mathcal{N}(0,I_{d})$, then the choice of $\mu_{t}$ should satisfy $|\mu_{t}(X_{i}^{*})|\leq|X_{i}^{*}|$. $\sigma_{t}$ should map to $[0,1]$ to avoid unnecessarily increasing the Lipschitz constant $\Gamma_{t}$ of the vector field. Generalizing the linear choice of $\sigma_{t}$ to
$\sigma_{t}(X_{i}^{*})=(1-(1-\sigma_{\mathrm{min}}^{\frac{1}{k}}{})^{k}$
for a $k>0$ has only a negligible effect. The curvature cancels out in the construction of $v_{t}(\cdot|X_{i}^{*}),$ see Lemma 2.3.
In order to obtain quantitative results for the neural network approximation, we further need to specify the neural networks. We use fully connected feedforward ReLU networks, which enjoy nice approximation properties, see for example Yarotsky [2017], Gühring et al. [2020], Kohler and Langer [2021], Schmidt-Hieber [2020], Suzuki [2019].
Let us fix some notation: define the Rectified Linear Unit (ReLU) activation function, $\phi\colon\mathbb{R}\rightarrow\mathbb{R}$ $\phi(x)=\max(0,x)$. For a vector $v=\left(v_{1},\ldots,v_{p}\right)\in\mathbb{R}^{p}$, we define the shifted ReLU activation function $\phi_{v}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{p}$ as
$\phi_{v}(x)=\left(\phi\left(x_{1}-v_{1}\right),\ldots,\phi\left(x_{p}-v_{p}\right)\right),\quad x=\left(x_{1},\ldots,x_{p}\right)\in\mathbb{R}^{p}.$
For a number $L\in\mathbb{N}$ of hidden layers and a vector $\mathcal{A}=\left(p_{0},p_{1},\ldots,p_{L+1}\right)\in$ $\mathbb{N}^{L+2}$, a neural network of depth $L+1$ and architecture $\mathcal{A}$ is a function of the form
$f:\mathbb{R}^{p_{0}}\rightarrow\mathbb{R}^{p_{L+1}},\quad f(x)=W_{L}\circ\phi_{v_{L}}\circ W_{L-1}\circ\phi_{v_{L-1}}\circ\cdots\circ W_{1}\circ\phi_{v_{1}}\circ W_{0}\circ x,$ (16)
where $W_{i}\in\mathbb{R}^{p_{i+1}\times p_{i}}$ are weight matrices and $v_{i}\in\mathbb{R}^{p_{i}}$ are shift vectors.
Now we want to combine the approximation properties of ReLU networks with our above analysis to evaluate the performance of the Flow Matching mechanism from (10) where $\mathcal{M}=\mathsf{NN}(L,M,\Gamma)$ is a set of ReLU networks $v\colon[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d},(t,x)\mapsto v_{t}(x)$ with a fixed maximal number of layers $L$, at most $M$ nonzero weights and Lipschitz constant of $v_{t}$ of at most $\Gamma$ for any $t$. We obtain the following rate:
Theorem 3.6.
Grant Assumption 3.5 and assume $\operatorname{supp}(p^{*})\subset[-1,1]^{d}$ and $p^{*}\in B^{\alpha}_{1,\infty}$ with $\|p^{*}\|_{B^{\alpha}_{1,\infty}}\leq M<\infty,\alpha\in(0,1]$. Set $\sigma_{\min}=n^{-\frac{1}{2\alpha+d}}.$ Then there are sequences $L_{n},M_{n},\Gamma_{n}\in\mathbb{N}$ such that for $n$ big enough
$\mathbb{E}\big{[}\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{\psi}_{1}})\big{]}\lesssim n^{-\frac{1+\alpha}{2\alpha+d}}\log^{2}(n),$
where $\hat{\psi}$ is given by (10) with $\mathcal{M}=\mathsf{NN}(L_{n},M_{n},\Gamma_{n})$.
Remark 3.7.
  1. 1.
    The Lipschitz constraint in Theorem 3.6 is typically not enforced in practice. However, we only require a bound on the Lipschitz constant of order $\Gamma_{n}=\frac{2d+1}{\sigma_{t}^{3}}+\frac{1}{2}$ which is only a mild restriction for $\sigma_{\min}\to 0$ as $n\to\infty$.
  2. 2.
    Compared to the literature on diffusion models and the first results on Flow Matching, we do not use early stopping times, but rather set $\sigma_{\min}>0$ according to $n.$ In terms of variance, both procedures are interchangeable: calculating $\sigma_{t_{*}}$ in [Fukumizu et al., 2024, Theorem 9] leads to exactly the same variance. However, early stopping leads to a bias induced by the mean function $\mu$, which is not $1$ at a time $t<1.$ This could be fixed using an adapted mean function, which is $1$ at the time of the early stopping, but this leads to much more complicated structures of $\mu.$
Theorem 3.6 shows that Flow Matching achieves minimax optimal rates (up to logarithmic factors) for certain classes of unknown densities, giving some justification of their the empirical success. There are several limitations and potential avenues for further research in the results of this work.
First of all the networks used for Theorem 3.6 are very large and do not correspond to the networks used in practice. While theoretical worst-case error bounds will always lead to networks that are not exactly small, the work of Yakovlev and Puchkin [2025] gives hope for improvement. Up to now it is an open question, whether a smaller network would come at the cost of a non-optimal rate. In addition, the size of the network and the overall result could benefit from the theoretical use of more sophisticated networks, like the U-net (Ronneberger et al. [2015]) construction in Lipman et al. [2023].
We have focused on the statistical analysis of Flow Matching, neglecting the optimization process as a source of additional error as well as the numerical error for solving the ODE (10). Including these errors would be more in line with the real world. The same applies for including the sampling of $t\sim\mathbb{U}$ in the Flow Matching objective.
3.3 Density estimation on a linear subspace
The rate in Theorem 3.6 depends on the dimension $d$ and therefore exhibits the classical curse of dimensionality. In this section, we analyze a case, in which the curse of dimensionality can be circumvented and the rate depends on some intrinsic dimension $d^{\prime}<d$. Assume that the support of $\mathbb{P}^{*}$ is a subset of an $d^{\prime}$ dimensional linear subspace of $\mathbb{R}^{d}$. Let the linear subspace $V$ be defined by
$V\coloneqq\{x\in\mathbb{R}^{d}\colon\exists y\in\mathbb{R}^{d^{\prime}}\text{ s.t. }x=Ay\},$
where $A$ is a $\mathbb{R}^{d\times d^{\prime}}$ matrix with normalized orthogonal columns. Let $p^{*}_{d^{\prime}}$ be the density with respect to the Lebesgue measure on $\mathbb{R}^{d^{\prime}}$, such that for $Y\sim p^{*}_{d^{\prime}}$ we have that $AY\sim p^{*}$, where $p^{*}$ is the density of $\mathbb{P}^{*}$ with respect to the Lebesgue measure restricted on $V$. This setting has been studied in context of diffusions by Chen et al. [2023b] and Oko et al. [2023].
In this case we can improve the rate of convergence, such that the corresponding rate depends only on the intrinsic dimension $d^{\prime}$ and not on the ambient dimension $d.$ We focus on the case of the Gaussian kernel.
Theorem 3.8.
Assume $d^{\prime}\geq 2,p^{*}\in B_{1,\infty}^{\alpha},\alpha\in(0,1]$ with $\|p^{*}\|_{B_{1,\infty}^{\alpha}}\leq M$, $\operatorname{supp}(p^{*})$ bounded and $\mathbb{U}=\mathcal{N}(0,I_{d})$. Then there are $C_{1}$ and $C_{2}$ such that for $v^{n}$ from (8)
$\mathbb{E}\big{[}\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})\big{]}\leq C_{1}\sigma_{\min}^{1+\alpha}+\frac{C_{2}}{\sqrt{\sigma^{d^{\prime}-2}}}\frac{\log n}{\sqrt{n}}.$
For $\sigma_{\operatorname{min}}\sim(n/\log^{2}n)^{-\frac{1}{2\alpha+d^{\prime}}}$ we obtain $\mathbb{E}\big{[}\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})\big{]}=\mathcal{O}((n/\log^{2}n)^{-\frac{1+\alpha}{2\alpha+d^{\prime}}}).$
The above result is in line with the similar results for kernel density estimators for distributions on manifolds Berenfeld and Hoffmann [2021], Divol [2022]. Note however that in these articles, the kernel is normalized with respect to the lower dimension, i.e. they consider $h^{-d^{\prime}}K(\cdot/h)$ for bandwidth $h>0$, while our generative setting enforces us to prove Theorem 3.8 for kernels which are normalized with respect to the ambient space dimension, i.e. $h^{-d}K(\cdot/h)$ (and $h=\sigma_{\min}$ in our notation).
Analogous to Theorem 3.6, we obtain a rate of convergence in case ReLU networks are used for the vector field.
Theorem 3.9.
Grant Assumption 3.5 and assume $\operatorname{supp}(p^{*})\subset[-1,1]^{d}$ and $p^{*}\in B^{\alpha}_{1,\infty}$ with $\|p^{*}\|_{B^{\alpha}_{1,\infty}}\leq M<\infty,\alpha\in(0,1]$. Set $\sigma_{\min}=n^{-\frac{1}{2\alpha+d^{\prime}}}.$ Then there are sequences $L_{n},M_{n},\Gamma_{n}\in\mathbb{N}$ such that for $n$ big enough
$\mathbb{E}\big{[}\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{\psi}_{1}})\big{]}\lesssim n^{-\frac{1+\alpha}{2\alpha+d^{\prime}}}\log^{2}(n),$
where $\hat{\psi}$ is given by (10) with $\mathcal{M}=\mathsf{NN}(L_{n},M_{n},\Gamma_{n})$.
Theorem 3.8 and Theorem 3.9 demonstrate that Flow Matching can benefit from lower intrinsic dimensions. Note that the method automatically adapts to the linear subspace, only the dimension $d^{\prime}$ is required for the choice of $\sigma_{\min}$. However, the smaller error bound comes at the cost of even larger networks to ensure that the approximation error of the network remains negligible compared to the faster rate of convergence in the dominating first term in decomposition (15). Therefore, the above results can only serve as first step to an analysis in a dimension reduction setting. A natural next step would be a generalization from linear subspaces to $d^{\prime}$-dimensional sub-manifolds. Moreover, it would be desirable to achieve the same rate with a smaller network, ideally even benefiting from the smaller dimension.
4 Proofs
Before we begin with the proofs, let us introduce some additional notation: The spectral norm of $A$ is denoted by $\|A\|_{2}.$ We define the ceiling of $x\in\mathbb{R}$ as $\lceil x\rceil:=\min\{k\in\mathbb{Z}|k\geq x\}.$ For $\alpha\in(0,1]$, we define the Hölder space $\mathcal{H}^{\alpha}$ by
$\mathcal{H}^{\alpha}(\Omega)\coloneqq\big{\{}f\colon\Omega\rightarrow\mathbb{R}^{d}\colon\|f\|_{\mathcal{H}^{\alpha}(\Omega)}<\infty\big{\}},\quad\text{ where }\quad\|f\|_{\mathcal{H}^{\alpha}(\Omega)}\coloneqq\sup_{x,y\in\Omega}\frac{|f(x)-f(y)|}{|x-y|^{\alpha}}.$
For $\alpha=1,$ we obtain the space of Lipschitz functions. We abbreviate $\operatorname{Lip}(f)=\|f\|_{\mathcal{H}^{1}(\Omega)}.$ Again, in case of $\Omega=\mathbb{R}^{d},$ we omit $\Omega.$
4.1 Proofs of Section 2
Proof of Lemma 2.1
Proof.
A necessary and sufficient condition for $v_{t}$ to generate $p_{t}$ is given in (3). We thus verify
$\frac{d}{dt}p^{n}_{t}(x)$ $=\frac{1}{n}\sum_{i=1}^{n}\frac{d}{dt}p_{t}(x|X_{i}^{*})=-\frac{1}{n}\sum_{i=1}^{n}\operatorname{div}(p_{t}(x|X_{i}^{*})v_{t}(x|X_{i}^{*}))$
$=-\operatorname{div}\Big{(}\frac{1}{n}\sum_{i=1}^{n}p_{t}(x|X_{i}^{*})\sum_{i=1}^{n}\frac{p_{t}(x|X_{i}^{*})v_{t}(x|X_{i}^{*})}{\sum_{i=1}^{n}p_{t}(x|X_{i}^{*})}\Big{)}=-\operatorname{div}(p^{n}_{t}(x)v^{n}_{t}(x)).\qed$
Proof of Theorem 2.2
Proof.
For fixed $t\in[0,1]$ we have
$|\hat{v}_{t}(x)-v^{n}_{t}(x)|^{2}$ $=|\hat{v}_{t}(x)|^{2}-2\langle\hat{v}_{t}(x),v^{n}_{t}(x)\rangle+|v^{n}_{t}(x)|^{2},$
$|\hat{v}_{t}(x)-v_{t}(x|X_{i}^{*})|^{2}$ $=|\hat{v}_{t}(x)|^{2}-2\langle\hat{v}_{t}(x),v_{t}(x|X_{i}^{*})\rangle+|v_{t}(x|X_{i}^{*})|^{2}.$
The last term does not influence the minimal argument in $\hat{v}.$ For the first two we have
$\mathbb{E}_{X_{t}\sim p^{n}_{t}}[|\hat{v}_{t}(X_{t})|^{2}]=\int|\hat{v}_{t}(x)|^{2}p^{n}_{t}(x)\;\mathrm{d}x=\frac{1}{n}\sum_{i=1}^{n}\int|\hat{v}_{t}(x)|^{2}p_{t}(x|X_{i}^{*})\;\mathrm{d}x=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\tilde{X}_{t}\sim p_{t}(\cdot|X_{i}^{*})}[|\hat{v}_{t}(\tilde{X}_{t})|^{2}],$
and
$\mathbb{E}_{X_{t}\sim p_{t}}[\langle\hat{v}_{t}(X_{t}),v^{n}_{t}(X_{t})\rangle]$ $=\int\langle\hat{v}_{t}(x),v^{n}_{t}(x)\rangle p_{t}(x)\;\mathrm{d}x$
$=\int\Big{\langle}\hat{v}_{t}(x),\sum_{i=1}^{n}v_{t}(x|X_{i}^{*})\frac{p_{t}(x|X_{i}^{*})}{\sum_{i=1}^{n}p_{t}(x|X_{i}^{*})}\Big{\rangle}\frac{1}{n}\sum_{i=1}^{n}p_{t}(x|X_{i}^{*})\;\mathrm{d}x$
$=\frac{1}{n}\sum_{i=1}^{n}\int\langle\hat{v}_{t}(x),v_{t}(x|X_{i}^{*})\rangle p_{t}(x|X_{i}^{*})\;\mathrm{d}x$
$=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\tilde{X}_{t}\sim p_{t}(\cdot|X_{i}^{*})}[\langle\hat{v}_{t}(\tilde{X}_{t}),v_{t}(\tilde{X}_{t}|X_{i}^{*})\rangle].\qed$
Proof of Lemma 2.3
Proof.
As $\sigma_{t}(X_{i}^{*})>0$ for all $t,X_{i}^{*},$ we have that
$\psi_{t}^{-1}(x|X_{i}^{*})=\frac{x-\mu_{t}(X_{i}^{*})}{\sigma_{t}(X_{i}^{*})}.$
Hence
$\frac{\partial\psi_{t}}{\partial t}(z|X_{i}^{*})\Big{|}_{z=\psi_{t}^{-1}(x|X_{i}^{*})}=v_{t}(x|X_{i}^{*}),$
and
$\frac{\partial\psi_{t}}{\partial t}(x|X_{i}^{*})=\frac{\partial\sigma_{t}}{\partial t}(X_{i}^{*})x+\frac{\partial\mu_{t}}{\partial t}(X_{i}^{*}).$
Thus, we get
$v_{t}\left(x|X_{i}^{*}\right)=\frac{\frac{\partial\sigma_{t}}{\partial t}\left(X_{i}^{*}\right)}{\sigma_{t}\left(X_{i}^{*}\right)}\big{(}x-\mu_{t}\left(X_{i}^{*}\right)\big{)}+\frac{\partial\mu_{t}}{\partial t}\left(X_{i}^{*}\right).\qed$
4.2 Proofs and extended results of Section 3
Proof of Theorem 3.1
Proof.
As $\mathsf{W}_{1}$ and $\mathsf{W}_{2}$ satisfy the triangle inequality (see Villani [2008, p. 94]),
$\mathsf{W}_{q}(\mathbb{P}^{*},\mathbb{P}^{\hat{\psi}_{1}(Z)})\leq\mathsf{W}_{q}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})+\mathsf{W}_{q}(\mathbb{P}^{\psi^{n}_{1}(Z)},\mathbb{P}^{\hat{\psi}_{1}(Z)}).$
Using [Villani, 2008, Remark 6.5] we get
$\mathsf{W}_{1}(\mathbb{P}^{\psi^{n}_{1}(Z)},\mathbb{P}^{\hat{\psi}_{1}(Z)})$ $\leq\mathsf{W}_{2}(\mathbb{P}^{\psi^{n}_{1}(Z)},\mathbb{P}^{\hat{\psi}_{1}(Z)})$
$\leq\sqrt{\mathbb{E}[|\psi^{n}_{1}(Z)-\hat{\psi}_{1}(Z)|^{2}]}$
$=\Big{(}\int|\psi^{n}_{1}(x)-\hat{\psi}_{1}(x)|^{2}p_{0}(x)\;\mathrm{d}x\Big{)}^{1/2}.$
Like in [Albergo and Vanden-Eijnden, 2022, Proposition 3] set
$Q_{t}=\int|\psi^{n}_{t}(x)-\hat{\psi}_{t}(x)|^{2}p_{0}(x)\;\mathrm{d}x.$
Since $p^{*},\psi,$ $\hat{\psi}$ and $v_{t}^{n}$ are integrable with respect to the Lebesgue measure, we can use the dominated convergence theorem and interchange integration and differentiation to obtain
$\frac{d}{dt}Q_{t}$ $=2\int(\psi^{n}_{t}(x)-\hat{\psi}_{t}(x))^{\top}(v^{n}_{t}(\psi^{n}_{t}(x))-\hat{v}_{t}(\hat{\psi}_{t}(x)))p_{0}(x)\;\mathrm{d}x$
$=2\int(\psi^{n}_{t}(x)-\hat{\psi}_{t}(x))^{\top}(v^{n}_{t}(\psi^{n}_{t}(x))-\hat{v}_{t}(\psi^{n}_{t}(x)))p_{0}(x)\;\mathrm{d}x$
$\quad+2\int(\psi^{n}_{t}(x)-\hat{\psi}_{t}(x))^{\top}(\hat{v}_{t}(\psi^{n}_{t}(x))-\hat{v}_{t}(\hat{\psi}_{t}(x)))p_{0}(x)\;\mathrm{d}x.$
As $0\leq|\psi^{n}_{t}(x)-\hat{\psi}_{t}(x)-(v^{n}_{t}(\psi^{n}_{t}(x))-\hat{v}_{t}(\psi^{n}_{t}(x)))|^{2},$ we get
$2(\psi^{n}_{t}(x)-\hat{\psi}_{t}(x))^{\top}(v^{n}_{t}(\psi^{n}_{t}(x))-\hat{v}_{t}(\psi^{n}_{t}(x)))\leq|\psi^{n}_{t}(x)-\hat{\psi}_{t}(x)|^{2}+|v^{n}_{t}(\psi^{n}_{t}(x))-\hat{v}_{t}(\psi^{n}_{t}(x))|^{2},$
and since for fixed $t$ the vector field $\hat{v}_{t}$ is $\Gamma_{t}$ Lipschitz continuous
$2(\psi^{n}_{t}(x)-\hat{\psi}_{t}(x))^{\top}(\hat{v}_{t}(\psi^{n}_{t}(x))-\hat{v}_{t}(\hat{\psi}_{t}(x)))\leq 2\Gamma_{t}|\psi^{n}_{t}(x)-\hat{\psi}_{t}(x)|^{2}.$
Therefore
$\frac{d}{dt}Q_{t}\leq(1+2\Gamma_{t})Q_{t}+2\int|v^{n}_{t}(\psi^{n}_{t}(x))-\hat{v}_{t}(\psi^{n}_{t}(x))|^{2}p_{0}(x)\;\mathrm{d}x.$
Now we use the formulation of Grönwall’s lemma in Walter [1970] and use $Q_{0}=0$ (this holds since $\psi_{0}(x)=\hat{\psi}_{0}(x)=x$). This leads to
$Q_{1}\leq e^{\int_{0}^{1}1+2\Gamma_{t}\;\mathrm{d}t}2\int_{0}^{1}\int|v^{n}_{t}(\psi^{n}_{t}(x))-\hat{v}_{t}(\psi^{n}_{t}(x))|^{2}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t.$ (17)
By Theorem 2.2 the vector field $\hat{v}$ is chosen such that it minimizes $\int_{0}^{1}\mathbb{E}_{X_{t}\sim p_{t}}\big{[}|\hat{v}_{t}(X_{t})-v^{n}_{t}(X_{t})|^{2}\big{]}\;\mathrm{d}t$ and $X_{t}=\psi_{t}(Z),$ thus we get for every $\tilde{v}\in\mathcal{M}$
$Q_{1}$ $\leq 2e^{\int_{0}^{1}1+2\Gamma_{t}\;\mathrm{d}t}\int_{0}^{1}\int|v^{n}_{t}(\psi^{n}_{t}(x))-\tilde{v}_{t}(\psi^{n}_{t}(x))|^{2}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t$
$\leq 2e^{\int_{0}^{1}1+2\Gamma_{t}\;\mathrm{d}t}\int_{0}^{1}\|v^{n}_{t}-\tilde{v}_{t}\|^{2}_{\infty}p_{0}(x)\;\mathrm{d}t,$
$\leq 2e^{\int_{0}^{1}1+2\Gamma_{t}\;\mathrm{d}t}\|v^{n}-\tilde{v}\|^{2}_{\infty}.$
Taking the square root leads to the result. ∎
Proof of Theorem 3.3
Proof.
Set $\mu_{n}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}}.$ We denote $K_{\sigma_{\min}}\coloneqq\frac{1}{\sigma_{\min}^{d}}K(\frac{\cdot}{\sigma_{\min}}).$ By the triangle inequality of the supremum we get
$\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})$ $=\sup_{f\in\mathrm{Lip}(1)}\int f(x)(p^{*}(z)-(K_{\sigma_{\min}}\ast\mu_{n})(z))\;\mathrm{d}z$
$\leq\sup_{f\in\mathrm{Lip}(1)}\int f(z)(p^{*}(z)-(K_{\sigma_{\min}}\ast p^{*})(z))\;\mathrm{d}z$
$\quad+\sup_{f\in\mathrm{Lip}(1)}\int f(z)((K_{\sigma_{\min}}\ast p^{*})(z)-(K_{\sigma_{\min}}\ast\mu_{n})(z))\;\mathrm{d}z.$
The first term is the bias of the kernel density estimator and the second term refers to the stochastic error.
First term: Using Fubini’s theorem and $\int K_{\sigma_{\min}}(y)\;\mathrm{d}y=1$, we rewrite the first term as
$\sup_{f\in\mathrm{Lip}(1)}\int f(z)(p^{*}(z)$ $-(K_{\sigma_{\min}}\ast p^{*})(z))\;\mathrm{d}z$
$=\sup_{f\in\mathrm{Lip}(1)}\int f(z)\Big{(}\int K_{\sigma_{\min}}(y)p^{*}(z)\;\mathrm{d}y-\int K_{\sigma_{\min}}(y)p^{*}(z-y)\;\mathrm{d}y\Big{)}\;\mathrm{d}z$
$=\sup_{f\in\mathrm{Lip}(1)}\int K_{\sigma_{\min}}(y)\Big{(}\int f(z)p^{*}(z)\;\mathrm{d}z-\int f(z)p^{*}(z-y)\;\mathrm{d}z\Big{)}\;\mathrm{d}y$
$=\sup_{f\in\mathrm{Lip}(1)}\int K_{\sigma_{\min}}(y)\big{(}(f\ast p^{*}(-\cdot))(0)-(f\ast p^{*}(-\cdot))(y)\big{)}\;\mathrm{d}y.$
Then, since $|p^{*}|_{B^{\alpha}_{1,\infty}}\leq M,$ we have that
$|\nabla(f\ast p^{*}(-\cdot))(x)-\nabla(f\ast p^{*}(-\cdot))(y)|$ $=\Big{|}\int\nabla f(z)p^{*}(x-z)\;\mathrm{d}z-\int\nabla f(z)p^{*}(y-z)\;\mathrm{d}z\Big{|}$
$\leq\sqrt{d}M|x-y|^{\alpha},$
where $\nabla$ denotes the vector of weak partial derivatives and we use that $|\nabla f|\leq\sqrt{d}$ for $f\in\operatorname{Lip}(1)$.
Therefore by the generalized mean value theorem, for every $y$ there exists a $\xi_{y}$ and a $C_{1}^{\prime}>0$ such that
$\sup_{f\in\mathrm{Lip}(1)}\int$ $K_{\sigma_{\min}}(y)\Big{(}\int f(z)p^{*}(z)\;\mathrm{d}z-\int f(z)p^{*}(z-y)\;\mathrm{d}z\Big{)}\;\mathrm{d}y$
$=\sup_{f\in\mathrm{Lip}(1)}\int K_{\sigma_{\min}}(y)(\nabla(f\ast p^{*}(-\cdot))(\xi_{y}))^{\top}y\;\mathrm{d}y$
$=\sup_{f\in\mathrm{Lip}(1)}\int K_{\sigma_{\min}}(y)(\nabla(f\ast p^{*}(-\cdot))(\xi_{y})-\nabla(f\ast p^{*}(-\cdot))(0))^{\top}y\;\mathrm{d}y$
$\quad+\int K_{\sigma_{\min}}(y)(\nabla(f\ast p^{*}(-\cdot))(0))^{\top}y\;\mathrm{d}y$
$\leq\sup_{f\in\mathrm{Lip}(1)}\int|K_{\sigma_{\min}}(y)||\nabla(f\ast p^{*}(-\cdot))(\xi_{y})-\nabla(f\ast p^{*}(-\cdot))(0)||y|\;\mathrm{d}y$
$\quad+\int K_{\sigma_{\min}}(y)(f\ast p^{*}(-\cdot))(0))^{\top}y\;\mathrm{d}y$
$\leq C_{1}^{\prime}\int|K_{\sigma_{\min}}(y)||y|^{1+\alpha}\;\mathrm{d}y+\int K_{\sigma_{\min}}(y)(\nabla(f\ast p^{*}(-\cdot))(0))^{\top}y\;\mathrm{d}y.$
For the first term we have that
$\int K_{\sigma_{\min}}(y)|y|^{1+\alpha}\;\mathrm{d}y$ $=\frac{1}{\sigma_{\min}^{d}}\int K(\frac{y}{\sigma_{\min}})|y|^{1+\alpha}\;\mathrm{d}y$
$=\sigma_{\min}^{1+\alpha}\int K(u)|u|^{1+\alpha}\;\mathrm{d}u.$
For the last term we use the assumption $\int K_{\sigma_{\min}}(y)y\;\mathrm{d}y=0.$
Hence we get
$\sup_{f\in\mathrm{Lip}(1)}\int f(z)(p^{*}(z)-(K_{\sigma_{\min}}\ast p^{*})(z))\;\mathrm{d}z\leq\sigma_{\min}^{1+\alpha}C_{1}^{\prime}\int K(u)|u|^{1+\alpha}\;\mathrm{d}u.$
By assumption the last term is finite. Setting $C_{1}\coloneqq C_{1}^{\prime}\int K(u)|u|^{1+\alpha}\;\mathrm{d}u$ yields the first term in the final bound.
Second term: We rewrite the second term as
$\sup_{f\in\mathrm{Lip}(1)}\int f(z)((K_{\sigma_{\min}}\ast p^{*})(z)$ $-(K_{\sigma_{\min}}\ast\mu_{n})(z))\;\mathrm{d}z$
$=\sup_{f\in\mathrm{Lip}(1)}\int\int f(z)K_{\sigma_{\min}}(z-x)(p^{*}(x)-\mu_{n}(x))\;\mathrm{d}x\;\mathrm{d}z$
$=\sup_{f\in\mathrm{Lip}(1)}\int(K_{\sigma_{\min}}(-\cdot)\ast f)(x)p^{*}(x)\;\mathrm{d}x-\int(K_{\sigma_{\min}}(-\cdot)\ast f)(x)\mu_{n}(x)\;\mathrm{d}x$
$=\sup_{\begin{subarray}{c}g=K_{\sigma_{\min}}(-\cdot)\ast f\\ f\in\mathrm{Lip}(1)\end{subarray}}\int g(x)(p^{*}(x)-\mu_{n}(x))\;\mathrm{d}x.$
Using Youngs inequality and the properties of convolution, we obtain for $k\in\mathbb{N}^{d}_{0}$
$\|D^{k}(K_{\sigma_{\min}}\ast f)\|_{\infty}\leq\|D^{1}f\|_{\infty}\|D^{k-1}K_{\sigma_{\min}(-\cdot)}\|_{1}\leq 1\cdot\|D^{k-1}K_{\sigma_{\min}}\|_{1}.$
By assumption $\|D^{k-1}K\|_{1}\leq C$ and hence $\|D^{k-1}K_{\sigma_{\min}}\|_{1}\leq C\sigma_{\min}^{-k+1}.$ Now we can use [Schreuder, 2020, Theorem 4]. For even $d,$ set $k=\frac{d}{2}.$ Then there exists a constant $C^{\prime}_{2}>0$ such that
$\sup_{\begin{subarray}{c}g=K_{\sigma_{\min}}\ast f\\ f\in\mathrm{Lip}(1)\end{subarray}}\int g(x)(p^{*}(x)-\mu_{n}(x))\;\mathrm{d}x\leq\frac{C^{\prime}_{2}}{\sigma_{\min}^{\frac{d}{2}-1}}n^{-1/2}\log(n).$
For odd $d>1,$ we know that
$\|D^{\frac{d-1}{2}}(f\ast K_{\sigma_{\min}})\|_{\infty}\leq\|D^{1}f\|_{\infty}\|D^{\frac{d-3}{2}}K_{\sigma_{\min}}\|_{1}\leq C^{\prime}_{2}\sigma_{\min}^{-\frac{d-3}{2}},$
and
$\frac{D^{\frac{d-1}{2}}(f\ast K_{\sigma_{\min}})(x)-D^{\frac{d-1}{2}}(f\ast K_{\sigma_{\min}})(y)}{|x-y|^{1/2}}\leq 2C^{\prime}_{2}\sigma^{-\frac{d-3}{2}}\leq 2C^{\prime}_{2}\sigma^{-\frac{d}{2}+1}.$
Then we can use [Schreuder, 2020, Theorem 4] with $\alpha=\frac{d}{2}$ again. Therefore, there is a $C^{\prime\prime}_{2}>0$ such that
$\sup_{\begin{subarray}{c}g=K_{\sigma_{\min}}\ast f\\ f\in\mathrm{Lip}(1)\end{subarray}}\int g(x)(p^{*}(x)-\mu_{n}(x))\;\mathrm{d}x\leq\frac{C^{\prime\prime}_{2}}{\sigma^{\frac{d}{2}-1}}n^{-1/2}\log(n).$
Combining both terms we conclude that there are constants $C_{1}$ and $C_{2}$ with
$\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})\leq C_{1}\sigma_{\min}^{1+\alpha}+\frac{C_{2}}{\sigma^{\frac{d}{2}-1}}n^{-1/2}\log(n).\qed$
Proof of Theorem 3.6
The proof of Theorem 3.6 requires some additional results. In order to control the error $e^{\int_{0}^{1}\Gamma_{t}\;\mathrm{d}t}\|v^{n}-\hat{v}\|_{\infty}$ (where we omitted constants) we need to approximate the function $v^{n}$ and control the Lipschitz constant of the apprximation $\hat{v}_{t}$ for fixed $t.$ However, we cannot expect to approximate a non-tivial function on $\mathbb{R}^{d}$ with respect to the supremum norm. Additionally, $v^{n}_{t}$ as constructed by Lipman et al. [2023] is only locally Lipschitz in $x$ for fixed $t$ and the bound of the Lipschitz constant grows for $\sigma_{\min}\rightarrow 0$ (see Lemma 4.2). Hence we cannot expect to construct a network $\hat{v}$ that approximates $v^{n}$ and its Lipschitz constant on $\mathbb{R}^{d}$ with a small absolute error for every $x\in\mathbb{R}.$
In the given setting, a rapid decay of the kernel function results in less significant approximation errors outside of the support of $p^{*}.$ We first show that in case of rapidly decreasing kernels the error $\mathsf{W}_{1}(\mathbb{P}^{\psi(Z)},\mathbb{P}^{\hat{\psi}(Z)})$ can be decomposed into two approximation errors, one depending on the approximation of $v$ on a compact domain and one corresponding to the area outside with little probability mass.
Lemma 4.1.
Assume that $\operatorname{supp}(p^{*})\subset(-1,1)^{d}$ and $\sigma_{\min}\leq 1.$ Additionally, assume that $K(x)=\prod_{i=1}^{d}\varphi(x_{i}),$ where $\varphi$ is a one dimensional symmetric density that decays faster than $x\mapsto e^{-\lambda x}$ for $|x|>2$ and some $\lambda>0$. Then for every $a>1$ there is a network $\tilde{v},$ which is cutoff outside $(-a,a)^{d}$ such that
$\mathsf{W}_{1}(\mathbb{P}^{\psi_{1}^{n}(Z)},\mathbb{P}^{\hat{\psi}_{1}(Z)})\leq\sqrt{2e}e^{\int_{0}^{1}\Gamma_{t}\;\mathrm{d}t}\Big{(}\int_{0}^{1}\int_{(-a,a)^{d}}|v^{n}_{t}(\psi^{n}_{t}(x))-\tilde{v}_{t}(\psi^{n}_{t}(x))|^{2}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t+c\frac{a^{2d+2}e^{-da}}{\sigma_{\min}^{2}}\Big{)}^{1/2}.$
Proof.
First, recall the proof from Theorem 3.1. For any $a>0$ we can bound the term from (17)
$Q_{1}$ $\leq\sqrt{e}e^{2\int_{0}^{1}\Gamma_{t}\;\mathrm{d}t}\int_{0}^{1}\int|v^{n}_{t}(\psi^{n}_{t}(x))-\tilde{v}_{t}(\psi^{n}_{t}(x))|^{2}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t$
$=\sqrt{e}e^{2\int_{0}^{1}\Gamma_{t}\;\mathrm{d}t}\Big{(}\int_{0}^{1}\int_{(-a,a)^{d}}|v^{n}_{t}(\psi^{n}_{t}(x))-\tilde{v}_{t}(\psi^{n}_{t}(x))|^{2}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t$ (18)
$\quad+\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus(-a,a)^{d}}|v^{n}_{t}(\psi^{n}_{t}(x))-\tilde{v}_{t}(\psi^{n}_{t}(x))|^{2}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t\Big{)}.$
The second term can be bounded by
$\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus(-a,a)^{d}}|v^{n}_{t}(\psi^{n}_{t}(x))$ $-\tilde{v}_{t}(\psi^{n}_{t}(x))|^{2}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t$
$\leq 2\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus(-a,a)^{d}}\Big{(}|v^{n}_{t}(\psi^{n}_{t}(x))|^{2}+|\tilde{v}_{t}(\psi^{n}_{t}(x))|^{2}\Big{)}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t.$ (19)
Now
$|v^{n}_{t}(\psi^{n}_{t}(x))|^{2}\leq\max_{i\in\{1,\dots,n\}}|v_{t}(\psi^{n}_{t}(x)|X_{i}^{*})|^{2}\leq\frac{|X_{i}^{*}-(1-\sigma_{\min})\psi^{n}_{t}(x)|^{2}}{(1-(1-\sigma_{\min})t)^{2}}\leq 2\frac{|X_{i}^{*}|^{2}+(1-\sigma_{\min})^{2}|\psi^{n}_{t}(x)|^{2}}{(1-(1-\sigma_{\min})t)^{2}}.$
As $\mathbb{P}^{*}$ has compact support within $(-a,a)^{d}$,
$|X_{i}^{*}|^{2}\leq d\cdot a^{2}.$
To bound $\psi_{t}^{n}(x)$ for $x\in\mathbb{R}^{d}\setminus(-a,a)^{d}$, we use that
$|\psi^{n}_{t}(x)|\leq|x|.$ (20)
To see why this is true, consider two cases. First, let $t$ be small such that $\psi^{n}_{t}(x)\notin(-a,a)^{d}.$ Then for every $X_{i}^{*},i\in\{1,\dots,n\}$ we have by construction that $v^{n}_{t}(\psi_{t}^{n}(x)|X_{i}^{*})$ is in the smallest convex cone that includes $\psi_{t}^{n}(x)$ and $(-1,1)^{d}$, where $\psi_{t}^{n}(x)$ is the vertex of the convex cone. Any linear combination of $v^{n}_{t}(\psi_{t}^{n}(x)|X_{i}^{*}),i=1,\dots,n,$ must be in this convex cone as well. Therefore $v_{t}^{n}(x)$ must be in the convex cone. Since
$\frac{d}{dt}\psi^{n}_{t}(x)=v^{n}_{t}(\psi_{t}(x)),\quad\psi^{n}_{0}(x)=x,$
this means that for small $t$, $\psi_{t}$ pushes the initial condition $x$ in the direction of $(-a,a)^{d}$ and therefore (20) is true.
Second, let $t^{\prime}$ be the smallest $t$ such that $\psi^{n}_{t^{\prime}}(x)\in(-a,a)^{d}.$ Then for all $t>t^{\prime}$ we have that $\psi^{n}_{t}(x)\in(-a,a)^{d}.$ This holds since within $(-a,a)^{d}\setminus(-1,1)^{d}$ using the same argument as in the first case, $v_{t}^{n}(x)$ is a linear combination of $v^{n}_{t}(\psi_{t}^{n}(x)|X_{i}^{*})$ and thus $\psi_{t}$ pushes $x$ in the direction of $(-1,1)^{d}.$ If $t$ is such that $\psi_{t}(x)\in(-1,1)^{d},$ there cannot be a $t^{\prime}>t$ such that $\psi_{t^{\prime}}(x)\notin(-a,a)^{d},$ since by continuity of $\psi_{t},$ there must be a $t^{\prime\prime}$ with $t<t^{\prime\prime}<t^{\prime}$ such that $\psi_{t^{\prime\prime}}(x)\in(-a,a)^{d}.$ In this case the corresponding vector field is oriented towards $(-1,1)^{d},$ which leads to $\psi_{t}$ pushing $x$ back to $(-1,1)^{d}.$ Hence, in any case for $x\notin(-a,a)^{d}$, we have that (20) is true.
Therefore we get for (19) for $\sigma_{\min}\leq 1$
$\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus(-a,a)^{d}}$ $|v^{n}_{t}(\psi^{n}_{t}(x))-\tilde{v}_{t}(\psi^{n}_{t}(x))|^{2}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t\leq 8\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus(-a,a)^{d}}\frac{da^{2}+(1-\sigma_{\min})^{2}|x|^{2}}{(1-(1-\sigma_{\min})t)^{2}}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t$
$\leq\frac{8da^{2}}{\sigma_{\min}^{2}}\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus(-a,a)^{d}}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t+\frac{(1-\sigma_{\min})^{2}}{\sigma_{\min}^{2}}\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus(-a,a)^{d}}|x|^{2}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t.$
For the first term we get
$\frac{8da^{2}}{\sigma_{\min}^{2}}\int_{\mathbb{R}^{d}\setminus(-a,a)^{d}}p_{0}(x)\;\mathrm{d}x$ $=\frac{8da^{2}}{\sigma_{\min}^{2}}\Big{(}\int_{\mathbb{R}\setminus(-a,a)}\varphi(x)\;\mathrm{d}x\Big{)}^{d},$
where $\varphi$ is the pdf of the one dimensional kernel distribution. Since $a>1$, we get for the second term
$\frac{(1-\sigma_{\min})^{2}}{\sigma_{\min}^{2}}\int_{\mathbb{R}^{d}\setminus(-a,a)^{d}}|x|^{2}p_{0}(x)\;\mathrm{d}x$ $=\frac{(1-\sigma_{\min})^{2}}{\sigma_{\min}^{2}}\sum_{i=1}^{d}\int_{\mathbb{R}^{d}\setminus(-a,a)^{d}}x_{i}^{2}p_{0}(x)\;\mathrm{d}x_{i}$
$\leq\frac{(1-\sigma_{\min})^{2}}{\sigma_{\min}^{2}}d\Big{(}\int_{\mathbb{R}\setminus(-a,a)}x^{2}\varphi(x)\;\mathrm{d}x\Big{)}^{d}.$
By assumption $\varphi$ decays faster than $e^{-x}.$ Using the upper incomplete $\Gamma$-function as defined in Lemma A.2 and Lemma A.2, we obtain for $a>2$
$\int_{\mathbb{R}\setminus(-a,a)}\varphi(x)\;\mathrm{d}x$ $\leq\int_{\mathbb{R}\setminus(-a,a)}x^{2}\varphi(x)\;\mathrm{d}x$
$\leq\int_{\mathbb{R}\setminus(-a,a)}x^{2}e^{-x}\;\mathrm{d}x$
$=2\int_{a}^{\infty}x^{2}e^{-x}\;\mathrm{d}x$
$=2\Gamma(3,a)$
$\leq 6e^{-a}a^{2}.$
Hence
$\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus(-a,a)^{d}}|v^{n}_{t}(\psi^{n}_{t}(x))$ $-\tilde{v}_{t}(\psi^{n}_{t}(x))|^{2}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t\leq c\frac{a^{2d+2}e^{-da}}{\sigma_{\min}^{2}}.\qed$
Next we want to choose $a$ in Lemma 4.1 and a neural network of finite size (depending on $n$) such that the bound of $\mathsf{W}_{1}(\mathbb{P}^{\psi^{n}(Z)},\mathbb{P}^{\hat{\psi}(Z)})$ decays at a rate of $n^{-\frac{1+\alpha}{2\alpha+d}}.$ Hence we need to approximate
$v^{n}_{t}(x)=\sum_{i=1}^{n}v_{t}(x|X_{i}^{*})\frac{p_{t}(x|X_{i}^{*})}{\sum_{i=1}^{n}p_{t}(x|X_{i}^{*})}$
and simultaneously control the Lipschitz constant of the approximation $\hat{v}$ on $(-a,a)^{d}.$ This is indeed sufficient, since using the same reasoning as in the proof of Lemma 4.1, it holds that if $x\in(-a,a)^{d}$, then $v_{t}^{n}(x)\in(-a,a)^{d}$ for every $t$. As $\hat{v}$ approximates $v^{n}$ in the supremum norm, the Lipschitz constant of $\hat{v}$ will be at least of the order of the Lipschitz constant of $v^{n}$. The same applies to the supremum norm. Hence we need bounds for both, the Lipschitz constant and the supremum norm of $v^{n}.$
Lemma 4.2.
Assume that $\operatorname{supp}(p^{*})\subset[-1,1]^{d}.$ For every $x\in(-a,a)$ we have that
$|v^{n}_{t}(x)|\leq\frac{\sqrt{d}(1+a)}{1-(1-\sigma_{\min})t}\leq\frac{\sqrt{d}(1+a)}{\sigma_{\min}}.$
Further for the Gaussian kernel and $\operatorname{supp}(p^{*})\subset[-1,1]^{d}$, for any $x\in\mathbb{R}^{d}$ and every $t\in[0,1]$
$\operatorname{Lip}(v^{n}_{t})\leq\frac{1}{\sigma_{t}}+\frac{2d}{\sigma_{t}^{3}}.$
Proof.
For the supremum norm bound, observe that
$|v^{n}_{t}(x)|$ $=\Big{|}\sum_{i=1}^{n}v_{t}(x|X_{i}^{*})\frac{p_{t}(x|X_{i}^{*})}{\sum_{i=1}^{n}p_{t}(x|X_{i}^{*})}\Big{|}$
$\leq\sum_{i=1}^{n}|v_{t}(x|X_{i}^{*})|\frac{p_{t}(x|X_{i}^{*})}{\sum_{i=1}^{n}p_{t}(x|X_{i}^{*})}$
$\leq\max_{i\in\{1,\dots,n\}}|v_{t}(x|X_{i}^{*})|$
$=\max_{i\in\{1,\dots,n\}}\frac{|(\sigma_{t}-t)X_{i}^{*}+x|}{1-(1-\sigma_{\min})t}$
$\leq\max_{i\in\{1,\dots,n\}}\frac{|X_{i}^{*}|+|x|}{1-(1-\sigma_{\min})t}$
$\leq\frac{\sqrt{d}(1+a)}{1-(1-\sigma_{\min})t}.$
For the Lipschitz bound, note that
$\nabla_{x}v_{t}^{n}(x)$ $=\nabla_{x}\sum_{i=1}^{n}\Big{(}\frac{\frac{\partial\sigma_{t}}{\partial t}}{\sigma_{t}}x-\mu_{t}(X_{i})+\frac{\partial\mu_{t}}{\partial t}(X_{i})\Big{)}\frac{p_{t}(x|X_{i})}{\sum_{j=1}^{n}p_{t}(x|X_{j})}$ (21)
$=\nabla_{x}\frac{-(1-\sigma_{\min})}{\sigma_{t}}x-\nabla_{x}\frac{-(1-\sigma_{\min})t}{\sigma_{t}}\frac{\sum_{i=1}^{n}X_{i}p_{t}(x|X_{i})}{\sum_{j=1}^{n}p_{t}(x|X_{j})}+\nabla_{x}\frac{\sum_{i=1}^{n}X_{i}p_{t}(x|X_{i})}{\sum_{j=1}^{n}p_{t}(x|X_{j})}.$
Now for the partial derivative of the $\ell$-st coordinate function with respect to $x_{k},$ $\ell,k\in\{1,\dots,d\},$ we get
$\frac{\partial}{\partial x_{k}}\frac{\sum_{i=1}^{n}X_{i,\ell}p_{t}(x|X_{i})}{\sum_{j=1}^{n}p_{t}(x|X_{j})}$
$=\frac{\big{(}\sum_{i=1}^{n}\big{(}-\frac{x_{k}-tX_{ik}}{\sigma_{t}^{2}}\big{)}X_{i,\ell}p_{t}(x|X_{i})\big{)}(\sum_{j=1}^{n}p_{t}(x|X_{j}))-\big{(}\sum_{i=1}^{n}X_{i,\ell}p_{t}(x|X_{i})\big{)}\big{(}\sum_{j=1}^{n}\big{(}-\frac{x_{k}-tX_{jk}}{\sigma_{t}^{2}}\big{)}p_{t}(x|X_{j})\big{)}}{\big{(}\sum_{j=1}^{n}p_{t}(x|X_{j})\big{)}^{2}}$
$=\frac{t}{\sigma_{t}^{2}}\Big{(}\frac{\sum_{i=1}^{n}X_{ik}X_{i\ell}p_{t}(x|X_{i})}{\sum_{j=1}^{n}p_{t}(x|X_{i})}-\frac{\big{(}\sum_{i=1}^{n}X_{ik}p_{t}(x|X_{i})\big{)}\big{(}\sum_{i=1}^{n}X_{i\ell}p_{t}(x|X_{i})\big{)}}{\big{(}\sum_{j=1}^{n}p_{t}(x|X_{i})\big{)}^{2}}\Big{)}.$
Since $\operatorname{supp}(p^{*})\subset[-1,1]^{d},$ we can bound
$\frac{\partial}{\partial x_{k}}\frac{\sum_{i=1}^{n}X_{i,\ell}p_{t}(x|X_{i})}{\sum_{j=1}^{n}p_{t}(x|X_{j})}\leq\frac{2t}{\sigma^{2}_{t}}.$
Using $t\in[0,1],$ $\sigma_{\min}\leq 1,$ we get for (21)
$\nabla_{x}v_{t}^{n}(x)\leq\frac{1}{\sigma_{t}}I_{d}+\frac{2}{\sigma_{t}^{3}}J_{d},$
where $I_{d}$ denotes the $d\times d$ identity matrix, $J_{d}$ denotes the $d\times d$ matrix consisting of ones and $\leq$ denotes entry wise inequality. Using the mean value theorem, we obtain for $x,y\in\mathbb{R}^{d}$
$|v_{t}^{n}(x)-v_{t}^{n}(y)|$ $\leq\Big{\|}\frac{1}{\sigma_{t}}I_{d}+\frac{2}{\sigma_{t}^{3}}J_{d}\Big{\|}_{2}|x-y|.$
As
$\Big{\|}\frac{1}{\sigma_{t}}I_{d}+\frac{2}{\sigma_{t}^{3}}J_{d}\Big{\|}_{2}$ $\leq\Big{\|}\frac{1}{\sigma_{t}}I_{d}\Big{\|}_{2}+\Big{\|}\frac{2}{\sigma_{t}^{3}}J_{d}\Big{\|}_{2}$
$=\frac{1}{\sigma_{t}}+\frac{2d}{\sigma_{t}^{3}},$
we get the desired bound. ∎
From [Gühring et al., 2020, Corollary 4.2] we know that for any $f\in C^{2}$ with bounded second partial derivatives we can approximate $f$ and its Lipschitz constant simultaneously using a ReLU network as defined in (16). To be more precise, for every $\varepsilon\in(0,1/2)$ there is a ReLU network $f_{\mathsf{NN}}$ such that
$\left\|f_{\mathsf{NN}}-f\right\|_{\mathcal{H}^{1}\left((0,1)^{d}\right)}\leq\varepsilon.$ (22)
The number of hidden layers $L$ of $f_{\mathsf{NN}}$ is bounded by $c\cdot\log\left(\varepsilon^{-n/(n-s)}\right)$ and the number of nonzero weights $M$ is bounded by $c\cdot\varepsilon^{-d/(n-s)}\cdot\log^{2}\left(\varepsilon^{-n/(n-s)}\right)$, where $c$ is a constant independent from $n$. Note that Gühring et al. [2020] refer to ReLU networks as defined in (16) as „standard neural networks“, since they also consider networks with skip connections.
Looking at the proof of Lemma 4.2 and using standard analysis, we can bound the second partial derivatives of $v^{n}_{t}$ by $\frac{C}{\sigma_{t}^{5}}$ for some $C>0$. Now we are ready to proof Theorem 3.6.
Proof.
From Lemma 4.1, we know that
$\mathsf{W}_{1}(\mathbb{P}^{\psi_{1}^{n}(Z)},\mathbb{P}^{\hat{\psi}_{1}(Z)})\leq\sqrt{2e}e^{\int_{0}^{1}\tilde{L}_{t}\;\mathrm{d}t}\inf_{\tilde{v}\in\mathcal{M}}\Big{(}\int_{0}^{1}\int_{(-a,a)^{d}}|v^{n}_{t}(\psi^{n}_{t}(x))-\tilde{v}_{t}(\psi^{n}_{t}(x))|^{2}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t+c\frac{a^{2d+2}e^{-da}}{\sigma_{\min}^{2}}\Big{)}^{1/2},$
where we used the choice of (10) before dividing the area in (18) in the proof of Lemma 4.1 and $\tilde{L}$ is the Lipschitz constant of $\tilde{v}$ at time $t$. We neglect an arbitrary small additional error on the right hand side to account for the use of the infimum, as the error bound corresponding to the kernel density estimator in (15) is nonzero anyway. Note that due to the restriction of the network class,
$e^{\int_{0}^{1}\tilde{L}_{t}\;\mathrm{d}t}\leq e\cdot e^{\int_{0}^{1}\frac{2d+1}{\sigma_{t}^{3}}\;\mathrm{d}t}.$
Integration yields for $n$ big enough
$\int_{0}^{1}\frac{1}{\sigma_{t}^{3}}\;\mathrm{d}t=\int_{0}^{1}\frac{1}{(1-(1-\sigma_{\min})t)^{3}}\;\mathrm{d}t\leq\frac{1}{2(1-\sigma_{\min})\sigma_{\min}^{2}}-\frac{1}{2(1-\sigma_{\min})}\leq\frac{1}{\sigma_{\min}^{2}}.$
To bound the tail, we need to assure $a$ is chosen such that
$e^{\frac{2d+1}{\sigma_{\min}^{2}}}\frac{a^{d+1}e^{-\frac{da}{2}}}{\sigma_{\min}}\leq n^{-\frac{1+\alpha}{2\alpha+d}}.$
As in the setting $d\geq 1,$ we can use that for $a\geq 15$
$a^{d+1}e^{-da/2}\leq e^{-\frac{da}{8}}.$
For $\sigma_{\min}\sim n^{-\frac{1}{2\alpha+d}}$, the above inequality is satisfied if
$a\geq\max\big{(}\frac{8}{d}\big{(}\frac{2+\alpha}{2\alpha+d}\log(n)+(2d+1)n^{\frac{2}{2\alpha+d}}\big{)},15\big{)}.$
To cover $(-a,a)^{d}$ with cubes of size $(0,1)^{d}$, we need $\lceil 2a\rceil^{d}$ little cubes.
Now we want to bound the approximation term. We can bound
$\Big{(}\int_{0}^{1}\int_{(-a,a)^{d}}|v^{n}_{t}(\psi^{n}_{t}(x))-\tilde{v}_{t}(\psi^{n}_{t}(x))|^{2}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t\Big{)}^{1/2}$ $\leq\|v^{n}_{t}-\tilde{v}_{t}\|_{\infty,(-a,a)^{d}}.$
Note that $v_{t}^{n}\in C^{\infty}$ in case of the Gaussian kernel. Hence we can apply [Gühring et al., 2020, Corollary 4.2]. The bound of the second partial derivatives as well as the supremum norm bound of $v^{n}$ influence the coefficients in [Gühring et al., 2020, Corollary 4.2], therefore we construct the network such that $n^{-\frac{5}{2\alpha+d}}v^{n}$ is approximated and choose the last weight matrix such that it scales the result up. This addition of one layer does not influence the order of the layers and non zero weights of the network. However, the approximation error gets scaled up too, hence we need to ensure a smaller error. We need to choose $\varepsilon$ in (22) such that
$e^{(2d+1)n^{\frac{2}{2\alpha+d}}}\sqrt{\varepsilon}\leq n^{-\frac{6+\alpha}{2\alpha+d}}\quad\Longleftrightarrow\quad\varepsilon\leq n^{-\frac{12+2\alpha}{2\alpha+d}}e^{-(4d+2)n^{\frac{2}{2\alpha+d}}}.$
The addition of one cutting layer to ensure the assumptions of Lemma 4.1 does not influence the order of the layers and non zero weights of the network either. Hence choosing
$L$ $\gtrsim\Big{(}\max\Big{(}\log(n)+n^{\frac{2}{2\alpha+d}},15\Big{)}\Big{)}^{d}\cdot\big{(}\log^{2}(n)+n^{\frac{4}{2\alpha+d}}\big{)},$
$M$ $\gtrsim\Big{(}\max\Big{(}\log(n)+n^{\frac{2}{2\alpha+d}},15\Big{)}\Big{)}^{d}\cdot\big{(}n^{\frac{12+2\alpha}{2\alpha+d}}e^{(4d+2)n^{\frac{2}{2\alpha+d}}}\big{)}^{d}\cdot\big{(}\log^{2}(n)+n^{\frac{4}{2\alpha+d}}\big{)},$
yields the desired rate. ∎
Proof of Theorem 3.8
Proof.
Recall the definitions $\mu_{n}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}}$ and $K_{\sigma_{\min}}\coloneqq\frac{1}{\sigma_{\min}^{d}}K(\frac{\cdot}{\sigma_{\min}})$ from Theorem 3.3. We start again by decomposing
$\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})$ $\leq\sup_{f\in\mathrm{Lip}(1)}\int f(z)(p^{*}(z)-(K_{\sigma_{\min}}\ast p^{*})(z))\;\mathrm{d}z$
$\quad+\sup_{f\in\mathrm{Lip}(1)}\int f(z)((K_{\sigma_{\min}}\ast p^{*})(z)-(K_{\sigma_{\min}}\ast\mu_{n})(z))\;\mathrm{d}z.$
For the first term, we know from the proof of Theorem 3.3, using the Lebesgue measure restricted to $V$, that there is a $C_{1}$ such that
$\sup_{f\in\mathrm{Lip}(1)}\int f(z)(p^{*}(z)-(K_{\sigma_{\min}}\ast p^{*})(z))\;\mathrm{d}z\leq C_{1}\sigma_{\min}^{1+\alpha}.$
For the second term we use the orthogonality of $A$ to obtain
$\sup_{f\in\mathrm{Lip}(1)}$ $\int f(z)((K_{\sigma_{\min}}\ast p^{*})(z)-(K_{\sigma_{\min}}\ast\mu_{n})(z))\;\mathrm{d}z$
$=\sup_{f\in\mathrm{Lip}(1)}\int\frac{f(z)}{(2\pi\sigma_{\min}^{2})^{d/2}}\Big{(}\int\exp\Big{(}-\frac{\|z-x\|^{2}}{2\sigma_{\min}^{2}}\Big{)}p^{*}(x)\;\mathrm{d}x-\frac{1}{n}\sum_{i=1}^{n}\exp\Big{(}-\frac{\|z-X_{i}^{*}\|^{2}}{2\sigma_{\min}^{2}}\Big{)}\Big{)}\;\mathrm{d}z$
$=\sup_{f\in\mathrm{Lip}(1)}\int\frac{f(z)}{(2\pi\sigma_{\min}^{2})^{d/2}}\Big{(}\int\exp\Big{(}-\frac{\|z-AA^{\top}z\|^{2}+\|AA^{\top}z-x\|^{2}}{2\sigma_{\min}^{2}}\Big{)}p^{*}(x)\;\mathrm{d}x$ (23)
$\qquad\qquad\qquad-\frac{1}{n}\sum_{i=1}^{n}\exp\Big{(}-\frac{\|z-AA^{\top}z\|^{2}+\|AA^{\top}z-X_{i}^{*}\|^{2}}{2\sigma_{\min}^{2}}\Big{)}\Big{)}\;\mathrm{d}z.$ (24)
By assumption, there is for every $x\in V$ a $y\in\mathbb{R}^{d^{\prime}}$ such that $x=Ay$. As the support of of $\mathbb{P}^{*}$ is subset of $V$, for every $X_{i}^{*}$ there is a $Y_{i}^{*}$ mapping into $\mathbb{R}^{d^{\prime}}$ such that $X_{i}^{*}=AY_{i}^{*}$ and $Y_{i}^{*}\sim p^{*}_{d^{\prime}}$. Hence we get
$\int\exp\Big{(}$ $-\frac{\|z-AA^{\top}z\|^{2}+\|AA^{\top}z-x\|^{2}}{2\sigma_{\min}^{2}}\Big{)}p^{*}(x)\;\mathrm{d}x$
$=\int\exp\Big{(}-\frac{\|z-AA^{\top}z\|^{2}+\|AA^{\top}z-Ay\|^{2}}{2\sigma_{\min}^{2}}\Big{)}p^{*}_{d^{\prime}}(y)\;\mathrm{d}y$
$=\int\exp\Big{(}-\frac{\|z-AA^{\top}z\|^{2}+\|A^{\top}z-y\|^{2}}{2\sigma_{\min}^{2}}\Big{)}p^{*}_{d^{\prime}}(y)\;\mathrm{d}y,$
where the last equality holds due to the orthonormal columns of $A$. For (24) we proceed analogously. Now we can decompose every $z\in\mathbb{R}^{d}$ uniquely into
$z=Au+Bv,\quad\text{ for some }\quad u\in\mathbb{R}^{d^{\prime}},v\in\mathbb{R}^{d-d^{\prime}},$
where $B\in\mathbb{R}^{d\times(d-d^{\prime})}$ is a matrix that maps $\mathbb{R}^{d-d^{\prime}}$ into $V^{\perp}.$ As $A^{\top}B=0$ and therefore
$A^{\top}z=A^{\top}(Au+Bv)=A^{\top}Au=u\quad\text{ and similar }\quad(I_{d}-AA^{\top})z=v,$
we get for (23) and (24)
$\sup_{f\in\mathrm{Lip}(1)}\int\frac{f(z)}{(2\pi\sigma_{\min}^{2})^{d/2}}\Big{(}$ $\int\exp\Big{(}-\frac{\|z-AA^{\top}z\|^{2}+\|A^{\top}z-y\|^{2}}{2\sigma_{\min}^{2}}\Big{)}p^{*}_{d^{\prime}}(y)\;\mathrm{d}y$
$-\frac{1}{n}\sum_{i=1}^{n}\exp\Big{(}-\frac{\|z-AA^{\top}z\|^{2}+\|A^{\top}z-Y_{i}^{*}\|^{2}}{2\sigma_{\min}^{2}}\Big{)}\Big{)}\;\mathrm{d}z$
$=\sup_{f\in\mathrm{Lip}(1)}\int\int\frac{f(Au+Bv)}{(2\pi\sigma_{\min}^{2})^{(d-d^{\prime})/2}}$ $\exp\Big{(}-\frac{\|v\|^{2}}{2\sigma_{\min}^{2}}\Big{)}\;\mathrm{d}v\Big{(}\int\frac{1}{(2\pi\sigma_{\min}^{2})^{d^{\prime}/2}}\exp\Big{(}-\frac{\|u-y\|^{2}}{2\sigma_{\min}^{2}}\Big{)}p^{*}_{d^{\prime}}(y)\;\mathrm{d}y$
$-\frac{1}{n}\sum_{i=1}^{n}\frac{1}{(2\pi\sigma_{\min}^{2})^{d^{\prime}/2}}\exp\Big{(}-\frac{\|u-Y_{i}^{*}\|^{2}}{2\sigma_{\min}^{2}}\Big{)}\Big{)}\;\mathrm{d}u.$
For fixed $v\in\mathbb{R}^{d-d^{\prime}}$ and $u,w\in\mathbb{R}^{(d-d^{\prime})}$, we have that
[…]
$\leq\int\frac{|f(Au+Bv)-f(Aw+Bv)|}{(2\pi\sigma_{\min}^{2})^{(d-d^{\prime})/2}}\exp\Big{(}-\frac{\|v\|^{2}}{2\sigma_{\min}^{2}}\Big{)}\;\mathrm{d}v$
$\leq\int\frac{|A(u-w)|}{(2\pi\sigma_{\min}^{2})^{(d-d^{\prime})/2}}\exp\Big{(}-\frac{\|v\|^{2}}{2\sigma_{\min}^{2}}\Big{)}\;\mathrm{d}v$
$=|u-w|,$
where we used the Lipschitz continuity of $f$ in the second inequality, the orthogonality of the columns of $A$ and the density of $\mathcal{N}(0,I_{d-d^{\prime}})$ in the last equality. We conclude that the integral is Lipschitz in $u$ with Lipschitz constant $1$. Hence
$\sup_{f\in\mathrm{Lip}(1)}\int\int\frac{f(Au+Bv)}{(2\pi\sigma_{\min}^{2})^{(d-d^{\prime})/2}}\exp\Big{(}-\frac{\|v\|^{2}}{2\sigma_{\min}^{2}}\Big{)}\;\mathrm{d}v\Big{(}$ $\int\frac{1}{(2\pi\sigma_{\min}^{2})^{d^{\prime}/2}}\exp\Big{(}-\frac{\|u-y\|^{2}}{2\sigma_{\min}^{2}}\Big{)}p^{*}_{d^{\prime}}(y)\;\mathrm{d}y$
$-\frac{1}{n}\sum_{i=1}^{n}\frac{1}{(2\pi\sigma_{\min}^{2})^{d^{\prime}/2}}\exp\Big{(}-\frac{\|u-Y_{i}^{*}\|^{2}}{2\sigma_{\min}^{2}}\Big{)}\Big{)}\;\mathrm{d}u$
$\leq\sup_{f\in\mathrm{Lip}(1)}\int f(u)\;\mathrm{d}v\Big{(}$ $\int\frac{1}{(2\pi\sigma_{\min}^{2})^{d^{\prime}/2}}\exp\Big{(}-\frac{\|u-y\|^{2}}{2\sigma_{\min}^{2}}\Big{)}p^{*}_{d^{\prime}}(y)\;\mathrm{d}y$
$-\frac{1}{n}\sum_{i=1}^{n}\frac{1}{(2\pi\sigma_{\min}^{2})^{d^{\prime}/2}}\exp\Big{(}-\frac{\|u-Y_{i}^{*}\|^{2}}{2\sigma_{\min}^{2}}\Big{)}\Big{)}\;\mathrm{d}u.$
Now we can proceed along the proof of the second term of Theorem 3.3 to obtain the desired bound. ∎
Proof of Theorem 3.9
Proof.
We proceed completly analogeous to Theorem 3.6 replacing $d$ with $d^{\prime}$. However, since the vectorfield we want to approximate is still $d$ dimensional, we still need $\lceil 2a\rceil^{d}$ little cubes. Hence we choose
$L$ $\gtrsim\Big{(}\max\Big{(}\log(n)+n^{\frac{2}{2\alpha+d^{\prime}}},15\Big{)}\Big{)}^{d}\cdot\big{(}\log^{2}(n)+n^{\frac{4}{2\alpha+d^{\prime}}}\big{)},$
$M$ $\gtrsim\Big{(}\max\Big{(}\log(n)+n^{\frac{2}{2\alpha+d^{\prime}}},15\Big{)}\Big{)}^{d}\cdot\big{(}n^{\frac{12+2\alpha}{2\alpha+d^{\prime}}}e^{(4d^{\prime}+2)n^{\frac{2}{2\alpha+d^{\prime}}}}\big{)}^{d}\cdot\big{(}\log^{2}(n)+n^{\frac{4}{2\alpha+d^{\prime}}}\big{)},$
to obtain the desired rate. ∎
Appendix A Appendix
Theorem A.1.
Assume $p^{*}\in B_{1,\infty}^{\alpha},\alpha\in(0,1]$ with $\|p^{*}\|_{B_{1,\infty}^{\alpha}(\mathbb{R}^{d})}\leq M$. Further assume $K$ is a nonnegative kernel $d$-dimensional kernel such that $\int K_{\sigma_{\min}}(y)y\;\mathrm{d}y=0$,
$\int\max(|u|^{2},|u|^{d+4})K^{2}(u)\;\mathrm{d}u<\infty$ $\quad\text{and}\quad\int|x|^{d+4}p^{*}(x)\;\mathrm{d}x<\infty.$
If $\sigma_{\min}\leq 1$, then there are $C_{1}$ and $C_{2}$ such that for any $\tilde{v}\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ and $v$ from (8),
$\mathbb{E}\big{[}\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})\big{]}\leq C_{1}\sigma_{\min}^{1+\alpha}+\frac{C_{2}}{\sqrt{n\sigma_{\min}^{d}}}.$
Proof.
As in Theorem 3.3,
$\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})$ $=\sup_{f\in\mathrm{Lip}(1)}\int f(x)(p^{*}(z)-(K_{\sigma_{\min}}\ast\mu_{n})(z))\;\mathrm{d}z$
$\leq\sup_{f\in\mathrm{Lip}(1)}\int f(z)(p^{*}(z)-(K_{\sigma_{\min}}\ast p^{*})(z))\;\mathrm{d}z$
$\quad+\sup_{f\in\mathrm{Lip}(1)}\int f(z)((K_{\sigma_{\min}}\ast p^{*})(z)-(K_{\sigma_{\min}}\ast\mu_{n})(z))\;\mathrm{d}z.$ (25)
For the first term, we obtain as in Theorem 3.3
$\sup_{f\in\mathrm{Lip}(1)}\int f(z)(p^{*}(z)-(K_{\sigma_{\min}}\ast p^{*})(z))\;\mathrm{d}z\leq C_{1}\sigma_{\min}^{1+\alpha}.$
Second term: To bound the expectation of (25) we first use that by [Villani, 2008, Theorem 6.15]
$\sup_{f\in\mathrm{Lip}(1)}\int f(z)((K_{\sigma_{\min}}\ast p^{*})(z)-(K_{\sigma_{\min}}\ast\mu_{n})(z))\;\mathrm{d}z\leq\int|z||(K_{\sigma_{\min}}\ast p^{*})(z)-(K_{\sigma_{\min}}\ast\mu_{n})(z)|\;\mathrm{d}z.$
Then for $\rho>\frac{d}{2}$
$\mathbb{E}[\int$ $|z||(K_{\sigma_{\min}}\ast p^{*})(z)-(K_{\sigma_{\min}}\ast\mu_{n})(z)|\;\mathrm{d}z]$
$=\int(1+|z|)^{-\rho}|z|(1+|z|)^{\rho}\mathbb{E}\Big{[}\Big{|}\frac{1}{n}\sum_{i=1}^{n}K_{\sigma_{\min}}(z-X_{i})-\mathbb{E}[K_{\sigma_{\min}}(z-X_{i})]\Big{|}\Big{]}\;\mathrm{d}z$
$\leq\frac{1}{n}\sqrt{\int(1+|z|)^{-2\rho}}\;\mathrm{d}z\sqrt{\int|z|^{2}(1+|z|)^{2\rho}\mathbb{E}\Big{[}\Big{(}\sum_{i=1}^{n}K_{\sigma_{\min}}(z-X_{i})-\mathbb{E}[K_{\sigma_{\min}}(z-X_{i})]\Big{)}^{2}\Big{]}\;\mathrm{d}z},$ (26)
where the equality holds by Fubini’s theorem and the inequality holds by the Jensen inequality together with the Cauchy-Schwarz inequality.
Next we bound the expectation. Since $\mathbb{E}[K_{\sigma_{\min}}(z-X_{i})-\mathbb{E}[K_{\sigma_{\min}}(z-X_{i})]]=0$ and $X_{1},\dots,X_{n}$ are i.i.d. we have that
$\mathbb{E}\Big{[}\Big{(}\sum_{i=1}^{n}K_{\sigma_{\min}}(z-X_{i})-\mathbb{E}[K_{\sigma_{\min}}(z-X_{i})]\Big{)}^{2}\Big{]}$ $=n\operatorname{Var}\big{(}K_{\sigma_{\min}}(z-X_{1})\big{)}$
$\leq n\mathbb{E}[(K_{\sigma_{\min}}(z-X_{1}))^{2}].$
Therefore
$\int|z|^{2}(1+|z|)^{2\rho}$ $\mathbb{E}\Big{[}\Big{(}\sum_{i=1}^{n}K_{\sigma_{\min}}(z-X_{i})-\mathbb{E}[K_{\sigma_{\min}}(z-X_{i})]\Big{)}^{2}\Big{]}\;\mathrm{d}z$
$=n\int\mathbb{E}\Big{[}|X_{1}+\sigma_{\min}y|^{2}\big{(}1+|X_{1}+\sigma_{\min}y|^{2\rho}\big{)}^{2}K^{2}(y)\Big{]}\;\mathrm{d}y$
$\lesssim\frac{n}{\sigma_{\min}^{d}}\int\mathbb{E}\big{[}|X_{1}+\sigma_{\min}y|^{2}+|X_{1}+\sigma_{\min}y|^{2\rho+2}\big{]}K^{2}(y)\;\mathrm{d}y$
$\lesssim\frac{n}{\sigma_{\min}^{d}}\int\big{(}\mathbb{E}\big{[}|X_{1}|^{2\rho+2}\big{]}+\sigma_{\min}^{2}|y|^{2}+\sigma_{\min}^{2\rho+2}|y|^{2\rho+2}\big{)}K^{2}(y)\;\mathrm{d}y$
$\lesssim\frac{n}{\sigma_{\min}^{d}}(1+\sigma^{2}_{\min}+\sigma_{\min}^{2\rho+2})\int\max(|y|,|y|^{2\rho+2})K^{2}(y)\;\mathrm{d}y.$
Using that $\sigma_{\min}\leq 1$ we can bound (26) for $\rho>\frac{d}{2}$ further by
$\mathbb{E}[\int|z||(K_{\sigma_{\min}}\ast p^{*})(z)-(K_{\sigma_{\min}}\ast\mu_{n})(z)|\;\mathrm{d}z]$ $\lesssim\frac{1}{\sqrt{n\sigma_{\min}^{d}}}\int\max(|y|,|y|^{2\rho+2})K^{2}(y)\;\mathrm{d}y.$
For $p=\frac{d}{2}+1$ the last integral is finite and we thus obtain
$\mathbb{E}[\int|z||(K_{\sigma_{\min}}\ast p^{*})(z)$ $-(K_{\sigma_{\min}}\ast\mu_{n})(z)|\;\mathrm{d}z]\leq\frac{C_{2}}{\sqrt{n\sigma_{\min}^{d}}}.$
In the end, we get for (25) that
$\mathbb{E}\big{[}\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})\big{]}\leq C_{1}\sigma_{\min}^{1+\alpha}+\frac{C_{2}\max(1,{\sigma_{\min}},\sigma_{\min}^{p+1})}{\sqrt{n\sigma_{\min}^{d}}}.\qed$
For comprehensiveness, we translate the following result from German including the proof.
Lemma A.2.
[Gabcke, 1979, Satz 4.4.3] Let $a\geq 1$ and $x>a.$ The upper incomplete gamma function
$\Gamma(a,x):=\int_{x}^{\infty}e^{-v}v^{a-1}\;\mathrm{d}v$ (27)
can be bounded by
$\Gamma(a,x)\leq ae^{-x}x^{a-1}.$
Proof.
Let $a\geq 0$. Substituting $v=1/w$ in (27) leads to
[…]
Set
$h(w):=e^{-1/w}w^{-a-1},$
the first derivative of $h$ is
$h^{\prime}(w)=e^{-1/w}w^{-a-2}\left[\frac{1}{w}-(a+1)\right].$
If $x>a+1$, then $h^{\prime}(w)>0$ for all $w\in[0,\frac{1}{x}]$. Hence $h$ is strictly increasing and
$\Gamma(a,x)\leq\frac{1}{x}h\left(\frac{1}{x}\right)=e^{-x}x^{a}\quad(a\geq 0,x>a+1).$ (28)
For $a\geq 1$ integration by parts in (27) leads to
$\Gamma(a,x)=e^{-x}x^{a-1}+(a-1)\Gamma(a-1,x).$
Using (28) with $a-1$ instead of $a,$ we get
$\Gamma(a,x)\leq e^{-x}x^{a-1}+(a-1)e^{-x}x^{a-1}=ae^{-x}x^{a-1}\quad(a\geq 1,x>a).\qed$