### v2

# On the minimax optimality of Flow Matching through the connection to kernel density estimation

On the minimax optimality of Flow Matching through the connection to kernel density estimation
Lea Kunkel and Mathias Trabs [^1] We thank the participants of the Oberwolfach Mini-Workshop 2508b ”Statistical Challenges for Deep Generative Models” for inspiring discussions and helpful comments on this project.
(Ruhr University Bochum and Karlsruhe Institute of Technology)
Flow Matching has recently gained attention in generative modeling as a simple and flexible alternative to diffusion models. While existing statistical guarantees adapt tools from the analysis of diffusion models, we take a different perspective by connecting Flow Matching to kernel density estimation. We first verify that the kernel density estimator matches the optimal rate of convergence in Wasserstein distance up to logarithmic factors, improving existing bounds for the Gaussian kernel. Based on this result, we prove that for sufficiently large networks, Flow Matching achieves the optimal rate up to logarithmic factors. If the target distribution lies on a lower-dimensional manifold, we show that the kernel density estimator profits from the smaller intrinsic dimension on a small tube around the manifold. The faster rate also applies to Flow Matching, providing a theoretical foundation for its empirical success in high-dimensional settings.
Keywords: Generative models, continuous normalizing flows, rate of convergence, kernel density estimator, Wasserstein distance, distribution estimation
MSC 2020: 62E17, 62G07, 68T07
1 Introduction
The goal of generative models is to learn a map $\psi$ that transforms a random variable $Z$ drawn from a simple and known latent distribution $\mathbb{U}$ into a new variable $\psi(Z)$ whose distribution closely approximates the unknown target distribution $\mathbb{P}^{*}$. Hence, generative models are naturally linked to the classical non-parametric distribution estimation problem, enhanced by the additional requirement of enabling fast and efficient sampling from the learned distribution.
Flow Matching, a generative model introduced by Lipman et al. (2023), has recently attained significant interest due to its considerably more straightforward construction in comparison to diffusion models (building up on Sohl-Dickstein et al. (2015)), which have been regarded as the state-of-the-art generative method. Instead of learning $\psi$ directly, like GANs (Goodfellow et al., 2014), Flow Matching relies on learning a vector field $v$ that describes the direction of movement to transfer the probability mass from $\mathbb{U}$ to $\mathbb{P}^{*}$. This leads to the following ODE in time $t\in[0,1]$,
$\frac{d}{dt}\psi_{t}(x)=v_{t}(\psi_{t}(x)),\quad\psi_{0}(x)=x.$
The solution to this ODE yields a full path of probability densities via the push-forward distributions
$p_{t}=[\psi_{t}]_{\#}p_{0},$
where $p_{0}$ is the density of $\mathbb{U}$. The vector field $v$ should be chosen such that $p_{1}$ is close to the density $p^{*}$ of $\mathbb{P}^{*}$.
Originally, normalizing flows were used to transform a latent distribution into an approximation of an unknown distribution by an invertible map constructed according to the change of variable theorem. This method was defined by Tabak and Vanden-Eijnden (2010) and Tabak and Turner (2013), but popularized by Rezende and Mohamed (2015) and Dinh et al. (2015). For an overview see Kobyzev et al. (2020). Continuous normalizing flows (CNF), introduced by Chen et al. (2018), use a neural ordinary differential equation (ODE) to construct this mapping via a maximum likelihood method. However, the training requires simulations of the ODE, which results in high computational cost, see Grathwohl et al. (2019). The Flow Matching approach of Lipman et al. (2023) circumvents this by using a regression objective to train a CNF, which avoids simulations during training,
$\Psi(\tilde{v})\coloneqq\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}\big[\left|\tilde{v}_{t}(X_{t})-v_{t}\left(X_{t}|Y\right)\right|^{2}\big],$ (1)
where $p_{t}(\cdot|Y)$ is a time dependent conditional density path associated to the conditional vector field $v_{t}(\cdot|Y)$ and both $p_{t}(\cdot|Y)$ and $v_{t}(\cdot|Y)$ are known by construction. This regression objective is in certain cases similar to the score matching objective (Hyvärinen, 2005; Vincent, 2011), which is used to train diffusion models based on stochastic differential equations, introduced by Song et al. (2021).
Flow Matching algorithms have been successfully used in many different applications that benefit from efficient sampling, such as text-to-speech (Guo et al., 2024) and text-to-image (Yang et al., 2025; Esser et al., 2024) settings, the production of novel molecular or protein structures (Dunn and Koes, 2024; Bose et al., 2024) or the construction of surrogate models in high energy physics (Bieringer et al., 2024). It has also been adjusted theoretically to different settings. Atanackovic et al. (2024) adapted Flow Matching to the case of interacting particles, Gat et al. (2024) explored the discrete setting. Chen and Lipman (2024) generalized the Euclidean setting to the Riemannian setting, allowing for more general geometries. Kerrigan et al. (2024) extended Flow Matching to function spaces. However, the statistical properties have only recently been studied by Fukumizu et al. (2025) in the Wasserstein $2$ distance. They do not use the exact setting of Lipman et al. (2023), but rather introduce stopping times that depend on the number of samples, enabling the transfer of methods known from the statistical analysis of diffusion models.
In this work, we demonstrate that Flow Matching in the setting of Lipman et al. (2023) is closely related to the classical kernel density estimation (KDE). This connection allows us to analyze Flow Matching from a new perspective. First, we show that the motivation of Flow Matching also holds for its empirical counterparts where we replace the unknown target density $p^{*}$ in (1) with the empirical measure of the observations. For sufficiently large network classes, the resulting generative algorithm coincides exactly with a kernel density estimator, where the kernel is given by the density of the latent distribution. We show convergence rates for the kernel density estimator in Wasserstein $1$ distance in a setting that allows for the Gaussian kernel which is the most popular choice in Flow Matching applications. Our rates coincide with the minimax optimal rates from Niles-Weed and Berthet (2022) up to logarithmic factors. We then transfer our analysis to Flow Matching using neural network classes of parameterized vector fields. This particularly shows that even in case of perfect approximation, which corresponds to heavy overparameterization in practice, the Flow Matching estimator does not deteriorate to the empirical measure, but to an estimator that is minimax optimal. Put another way, Flow Matching can be, up to constants, at least as good as a KDE even without the presence of network-specific structures.
Unlike Fukumizu et al. (2025), we do not build on the similarities to diffusion models. Separating the error of the kernel density estimator allows us to analyse empirical risk minimization without the need for Chernoff-type bounds. This avoids the problems pointed out by Yakovlev and Puchkin (2025) in the analysis of diffusion models. Overall, the analysis of Flow Matching is more delicate, since Girsanov’s theorem does not apply and thus the strategy of Chen et al. (2023c) cannot be used. Therefore, our bound depends exponentially on the Lipschitz constant of the vector field. This is one of the reasons for the difficulties in the proof of Fukumizu et al. (2025) [^2] The bounds of the integral in the exponent of the first equation on p. 14 (in the ArXiV version) are not correct. The separation of the outer integral does not imply the same separation of the integral in the exponent. The correct factor in the integral is $\exp(2\int_{T_{0}}^{t_{j}}L_{u}du)$, where $L_{u}$ is the Lipschitz constant of the approximated vector field, which depends on $T_{0}$ and is thus not bounded by a universal constant.. The main theoretical novelty of this work compared to standard approaches in the statistical literature on diffusions and Flow Matching is to use the KDE as a reference model for the empirical risk minimizer. Therefore, we do not recover the classical trade-off in network size. Since in general, assuming that the Lipschitz constant of the vector field is bounded with a fixed constant requires an assumption on the unknown distribution (Kunkel, 2025), our approach requires significantly less assumptions on the unknown distribution. Our analysis allows us to show convergence rates in Wasserstein $1$ distance in the case where overparameterized neural networks are used for the vector field.
While these rates are minimax optimal up to a logarithmic constant, they suffer from the curse of dimensionality. In case $\mathbb{P}^{*}$ is supported on a lower-dimensional manifold, we show that, on a small tube around the manifold, the mass of the generated distribution concentrates according to the unknown distribution much faster, depending on the intrinsic dimension and thus circumventing the curse of dimensionality. This result again applies to both, the KDE and the Flow Matching estimator. For the former, our results are the first that do not require a structural adaption of the KDE to the manifold itself. Therefore, they contribute to the statistical understanding of the KDE under the manifold hypothesis previously studied by Divol (2022); Berenfeld and Hoffmann (2021).
Related work
Lipman et al. (2023) use a fixed latent distribution. Similar approaches for flows between two possibly unknown distributions $\mathbb{P}$ and $\mathbb{Q}$ are studied by Tong et al. (2024), Liu et al. (2022) and Albergo and Vanden-Eijnden (2023). Tong et al. (2024) also generalize the mentioned methods. Gao et al. (2024) prove a suboptimal rate of convergence for certain forms of Flow Matching not exactly conforming to the construction on Lipman et al. (2023) in the Wasserstein $2$ distance. Benton et al. (2024) analyzed Flow Matching excluding the approximation error by imposing assumptions on the covariance that lead to global Lipschitz bounds on the vector field. They also focus on different constructions than Lipman et al. (2023). Gong et al. (2025) study the properties of ReLU networks to approximate a vector field corresponding to higher order trajectories.
As Lipman et al. (2023) point out, Flow Matching is closely related to diffusion models. For an overview of generative diffusion models, see Cao et al. (2024). Indeed, even in cases that are not constructed to be consistent with diffusion models, the approximation of a score function has similar properties to the approximation of the Flow Matching vector field. Fukumizu et al. (2025) build up on this connection. The statistical properties of score matching are an area of ongoing research, see for example Chen et al. (2023a, b, c); Oko et al. (2023); Tang and Yang (2024); Azangulov et al. (2024); Zhang et al. (2024); Yakovlev and Puchkin (2025). Marzouk et al. (2023) study the statistical properties of CNFs trained by likelihood maximization.
The kernel density or Parzen-Rosenblatt estimator (Rosenblatt, 1956; Parzen, 1962) is the classical method for estimating a smooth density. For an overview in the univariate case, see Tsybakov (2009), Chapter 1.2 or Devroye and Lugosi (2001) and in the multivariate case see Scott (1992). Typically, the kernel density estimator is analyzed using the mean squared error or the $L_{1}$ distance. Despite its longstanding use, kernel density estimation is still the subject of ongoing research, for example in density estimation on unknown manifolds (Berry and Sauer, 2017; Berenfeld and Hoffmann, 2021; Divol, 2022; Wu and Wu, 2022). In particular, Divol (2022) uses the Wasserstein metric to evaluate the performance of a kernel density estimator, however he needs to assume properties of the kernel that are not satisfied by standard choices such as the Gaussian kernel.
Outline
In Section˜2 we introduce the problem setting mathematically and recall all the definitions needed to define the conditional Flow Matching objective from Lipman et al. (2023). We introduce the empirical counterparts, validate the empirical conditional Flow Matching objective, and portray the connection to the kernel density estimator. In Section˜3 we first separate the error corresponding to the kernel density estimator and the vector field approximation. We then proceed to bound the former error in Section˜3.1. Incorporating theory of network approximation, we bound the latter error and obtain a rate of convergence for the entire Flow Matching procedure in Section˜3.2. Subsequently, we show in Section˜3.3 that the rate can be improved if the unknown distribution is supported on a lower-dimensional manifold. Finally, we illustrate our results in Section˜4. All proofs are deferred to Section˜5.
1.1 Some notations
We write $X\sim p$ as a synonym for $X\sim\mathbb{P}$ in case $\mathbb{P}$ has a density $p$ with respect to the Lebesgue measure and $\mathbb{P}$ is the distribution of the random variable $X$. By $\lesssim$ we denote $a\leq c\cdot b,$ where $a,b\in\mathbb{R}$ and $c$ is a constant independent of $n$. For $d\in\mathbb{N}$, we use the Euclidean norm $|\cdot|$ on $\mathbb{R}^{d}$ and use $\|\cdot\|$ for associated operator norms. For a set $A\subset\mathbb{R}^{d}$, we denote the tubular neighborhood of the set $A$ with radius $r>0$ with $B(r,A)$. The supremum norm and the $L_{1}$-norm of a function $f\colon\Omega\rightarrow\mathbb{R}^{d^{\prime\prime}}$ with $d^{\prime},d^{\prime\prime}\in\mathbb{R}$ and $\Omega\subset\mathbb{R}^{d^{\prime}}$ is defined as
$\|f\|_{\infty,\Omega}\coloneqq\sup_{x\in\Omega}|f(x)|\quad\text{and}\quad\|f\|_{1,\Omega}\coloneqq\int_{\Omega}|f(x)|\;\mathrm{d}x.$
In case of $\Omega=\mathbb{R}^{d^{\prime}},$ we write $\|f\|_{\infty,\mathbb{R}^{d^{\prime}}}=\|f\|_{\infty}$ and $\|f\|_{1,\mathbb{R}^{d^{\prime}}}=\|f\|_{1}.$
For $d^{\prime\prime}=1,\alpha\in(0,1]$ and $\Lambda>0$, the Besov ball $B_{1,\infty}^{\alpha}(\Lambda,\infty)$ is defined as
$B_{1,\infty}^{\alpha}(\Lambda,\Omega)\coloneqq\big\{f\in L_{1}(\Omega)\colon|f|_{B_{1,\infty}^{\alpha}(\Omega)}\coloneqq\sup_{t>0}t^{-\alpha}\omega_{1}(f,t)_{1}\leq\Lambda\big\}$ (2)
where
$\omega_{1}(f,t)_{1}:=\sup_{0<|h|\leq t}\int|f(x)-f(x+h)|\;\mathrm{d}x,\quad t>0,x\in\Omega.$
The difference is set to zero if $x+h\notin\Omega.$ Replacing $|f|_{B_{1,\infty}^{\alpha}(\Omega)}<\Lambda$ with $|f|_{B_{1,\infty}^{\alpha}(\Omega)}<\infty$ in (2), we obtain the Besov space $B_{1,\infty}^{\alpha}(\Omega)$. On $B_{1,\infty}^{\alpha}(\Omega)$ we define the norm $\|f\|_{B_{1,\infty}^{\alpha}(\Omega)}=\|f\|_{L_{1}(\Omega)}+|f|_{B_{1,\infty}^{\alpha}(\Omega)}$. Compared to more commonly used smoothness classes, such as Hölder classes, the Besov class defined above only imposes integrated and not pointwise smoothness. This allows for local irregularities like jumps.
Finally, let $C^{k}(\Omega)$ denote the set of functions $f$ such that the coordinate functions are each $k$-times continuously differentiable and the supremum norm of the derivatives are finite. For $k\in\mathbb{N}_{0}^{d}$ and $|k|=\sum_{i=1}^{d}k_{i}$ set $D^{k}=\frac{\partial^{|k|}}{\partial x_{1}^{k_{1}}\ldots\partial x_{d}^{k_{d}}},$ where $\frac{\partial}{\partial x_{i}}$ denotes the weak partial derivative.
2 Empirical Flow Matching
Assume we observe an i.i.d. sample $X^{*}_{1},\dots,X^{*}_{n}$ from an unknown distribution $\mathbb{P}^{*}$ on a compact set $\mathcal{X}\subset[-1,1]^{d}.$ Further assume that $\mathbb{P}^{*}$ has a density $p^{*}$ with respect to the Lebesgue measure. The goal of generative modeling is to mimic the distribution $\mathbb{P}^{*}.$
For a time dependent vector field $v\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ we consider the flow $\psi\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ given as the solution to the ODE
$\frac{d}{dt}\psi_{t}(x)=v_{t}(\psi_{t}(x)),\quad\psi_{0}(x)=x.$ (3)
For a fixed latent distribution with Lebesgue density $p_{0}$, the vector field $v$ generates a probability density path $p\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{>0}$ with $\int p_{t}(x)\mathrm{d}x=1$ for all $t$ via the push-forward distributions
$p_{t}=[\psi_{t}]_{\#}p_{0},\quad\text{i.e. }\psi_{t}(Z)\sim p_{t}\text{ for }Z\sim p_{0}.$
The ODE (3) corresponds to the Lagrangian description (in terms of particle trajectories) of the conservation of mass formula (Villani, 2008, p.14). The change of variables formula links it to the Eulerian description: A necessary and sufficient condition for $v_{t}$ to generate $p_{t}$ is
$\frac{d}{dt}p_{t}+\operatorname{div}(p_{t}v_{t})=0,$ (4)
see (Villani, 2008, p. 14) .
Approximating a given vector field $v$ that generates a certain density path $p_{t}$ using a function $\tilde{v}$ leads to the Flow Matching objective
$\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1]\\ X_{t}\sim p_{t}\end{subarray}}\big[|v_{t}(X_{t})-\tilde{v}_{t}(X_{t})|^{2}\big]$ (5)
by Lipman et al. (2023).
They have consider a probability path of the form
$p_{t}(x)=\int p_{t}(x|y)p^{*}(y)\;\mathrm{d}y,$ (6)
where $p_{t}(\cdot|y)\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a conditional probability path generated by some vector field $v_{t}(\cdot|y)\colon\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ for $y\in\mathbb{R}^{d}$. The vector field generating (6) is then given by
$v_{t}(x)=\int v_{t}(x|y)\frac{p_{t}(x|y)p^{*}(y)}{p_{t}(x)}dy.$ (7)
In this setting Lipman et al. (2023) show that the Flow Matching objective (5) is equivalent to the conditional Flow Matching objective
$\Psi(\tilde{v})\coloneqq\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}\big[\left|\tilde{v}_{t}(X_{t})-v_{t}\left(X_{t}|Y\right)\right|^{2}\big].$ (8)
As $p^{*}$ is unknown, in practice, the expectation in $Y\sim p^{*}$ is replaced by the empirical counterpart based on the observations $X_{1}^{*},\dots,X_{n}^{*}$. This leads to the empirical counterparts of (6) and (7) given by
$p^{n}_{t}(x)$ $=\frac{1}{n}\sum_{i=1}^{n}p_{t}(x|X_{i}^{*})\qquad\text{and}\qquad v^{n}_{t}(x)=\sum_{i=1}^{n}v_{t}(x|X_{i}^{*})\frac{p_{t}(x|X_{i}^{*})}{\sum_{j=1}^{n}p_{t}(x|X_{j}^{*})}.$ (9)
With this modification we recover the sufficient condition for $v^{n}_{t}$ to generate $p^{n}_{t}$ analogously to (Lipman et al., 2023, Theorem 1).
Lemma 2.1.
If $v_{t}(\cdot|X_{i}^{*})$ generates $p_{t}(\cdot|X_{i}^{*})$ for all $i=1,\dots,n,$ then $v^{n}_{t}$ generates $p^{n}_{t}.$
The motivation for the conditional Flow Matching objective (8) is the equivalence to the unconditioned Flow Matching objective with respect to the optimizing arguments (Lipman et al., 2023, Theorem 2). Using the empirical counterparts, this still holds true.
Theorem 2.2.
Using $p^{n}_{t}$ and $v^{n}_{t}$ from (9), and a class of parameterized vector fields $\mathcal{N},$ which is constructed such that the minimal arguments exists, we have that
$\underset{\hat{v}\in\mathcal{N}}{\mathrm{argmin}}\int_{0}^{1}\mathbb{E}_{X_{t}\sim p^{n}_{t}}\big[|\hat{v}_{t}(X_{t})-v^{n}_{t}(X_{t})|^{2}\big]\;\mathrm{d}t=\underset{\hat{v}\in\mathcal{N}}{\mathrm{argmin}}\int_{0}^{1}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\tilde{X}_{t}\sim p_{t}(\cdot|X_{i}^{*})}\big[|\hat{v}_{t}(\tilde{X}_{t})-v_{t}(\tilde{X}_{t}|X_{i}^{*})|^{2}\big]\;\mathrm{d}t.$
Hence the empirical conditional Flow Matching objective
$\tilde{\Psi}(\tilde{v})\coloneqq\int_{0}^{1}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\tilde{X}_{t}\sim p_{t}(\cdot|X_{i}^{*})}\big[|\tilde{v}_{t}(\tilde{X}_{t})-v_{t}(\tilde{X}_{t}|X_{i}^{*})|^{2}\big]\;\mathrm{d}t$ (10)
is justified theoretically. Compared to Gao et al. (2024), we use $v_{t}(\cdot|X_{i}^{*})$ directly as proposed by Lipman et al. (2023) and stick to the entire time interval. Note that for simplicity we do not sample $t.$
From minimizing (10) in $\tilde{v}\in\mathcal{N},$ we obtain an optimal argument $\hat{v}$. Solving the ODE (3) using $\hat{v}$, we obtain a flow $\hat{\psi}_{t}$, i.e. $\hat{\psi}$ is given by
[…]
We use this flow to push forward the known, latent distribution to time $t=1$. In accordance with the goal of generative modeling, the distribution of this pushforward should mimic $\mathbb{P}^{*}.$
In order to apply Flow Matching, we have to construct a class of conditional probability paths. Let $Z\sim\mathbb{U}$ and let $K$ denote the density of $\mathbb{U}$. Consider the functions $\sigma\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{>0}$ and $\mu\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}.$ Set for $t\in[0,1]$ and a given $X_{i}^{*},i\in\{1,..,n\}$,
$\psi_{t}(Z|X_{i}^{*})\coloneqq\sigma_{t}(X_{i}^{*})Z+\mu_{t}(X_{i}^{*}).$ (12)
The density of $\psi_{t}(Z|X_{i}^{*})$ is by the transformation formula
$p_{t}(x|X_{i}^{*})=\frac{1}{\sigma^{d}_{t}(X_{i}^{*})}K\Big(\frac{x-\mu_{t}(X_{i}^{*})}{\sigma_{t}(X_{i}^{*})}\Big).$
We call $p_{t}(\cdot|X_{i}^{*})$ the conditional kernel probability path. Setting
$\frac{d}{dt}\psi_{t}(x|X_{i}^{*})=v_{t}(\psi_{t}(x|X_{i}^{*})|X_{i}^{*}),$ (13)
we recover the same result like in (Lipman et al., 2023, Theorem 3):
Lemma 2.3.
Let $p_{t}\left(x|X_{i}^{*}\right)$ be a conditional kernel probability path, and $\psi_{t}(\cdot|X_{i}^{*})$ its corresponding flow as in (12). Then, the unique vector field that defines $\psi_{t}(\cdot|X_{i}^{*})$ via (13) has the form:
$v_{t}\left(x|X_{i}^{*}\right)=\frac{\frac{\partial\sigma_{t}}{\partial t}\left(X_{i}^{*}\right)}{\sigma_{t}\left(X_{i}^{*}\right)}\big(x-\mu_{t}\left(X_{i}^{*}\right)\big)+\frac{\partial\mu_{t}}{\partial t}\left(X_{i}^{*}\right).$
Set $\sigma_{\min}>0.$ To flow from $p^{n}_{0}(x)=\frac{1}{n}\sum_{i=1}^{n}K(x)=K(x)$ to $p^{n}_{1}(x)=\frac{1}{n\sigma^{d}_{\min}}\sum_{i=1}^{n}K(\frac{x-X_{i}^{*}}{\sigma_{\min}}),$ we can choose any differentiable functions $\sigma_{t}$ and $\mu_{t}$ such that
$\mu_{0}=0,\quad\mu_{1}=1\quad\text{and}\quad\sigma_{0}=1,\quad\sigma_{1}=\sigma_{\min}.$
At time $t=1$ the distribution $\mathbb{P}^{\psi_{1}^{n}(Z)}$ then coincides with the kernel density estimator
$p^{n}_{1}(x)=\frac{1}{n\sigma^{d}_{\min}}\sum_{i=1}^{n}K\Big(\frac{x-X_{i}^{*}}{\sigma_{\min}}\Big),$ (14)
where the kernel is given by the latent distribution $\mathbb{U}$. Choosing $\mathbb{U}=\mathcal{N}_{d}(0,1)$, i.e., we consider the $d$-dimensional Gaussian kernel $K(x)=(2\pi)^{-d/2}\exp(-|x|^{2}/2)$, yields the proposed flow from Lipman et al. (2023, Section 4). Moreover, considering general kernels is in line with methods by Tong et al. (2024), Liu et al. (2022) and Albergo and Vanden-Eijnden (2023), which transform an unknown distribution to another.
3 Rate of convergence
The aim of this work is to evaluate how well Flow Matching performs depending on the number of observations $n.$ To this end, we have to control the distance between $\mathbb{P}^{*}$ and $\mathbb{P}^{\hat{\psi}_{1}(Z)}$ with flow $\hat{\psi}$ from (11). As as evaluation metric, we use the Wasserstein $1$ metric.
On the normed vector space $(\mathbb{R}^{d},|\cdot|)$ the Wasserstein $1$ distance between two probability distributions $\mathbb{P}$ and $\mathbb{Q}$ on $\mathbb{R}^{d}$ is defined as
$\mathsf{W}_{1}(\mathbb{P},\mathbb{Q})$ $\coloneqq\Big(\inf_{\pi\in\Pi(\mathbb{P},\mathbb{Q})}\int_{\mathcal{X}}|x-y|\;\mathrm{d}\pi(x,y)\Big),$
where $\Pi$ is the set of all distributions whose marginal distributions correspond to $\mathbb{P}$ and $\mathbb{Q}$ respectively. By (Villani, 2008, Theorem 5.10(i)), the following duality holds:
$\mathsf{W}_{1}(\mathbb{P},\mathbb{Q})=\sup_{W\in\operatorname{Lip}(1)}\mathbb{E}_{\begin{subarray}{c}X\sim\mathbb{P}\\ Y\sim\mathbb{Q}\end{subarray}}[W(X)-W(Z)].$ (15)
On the space of probability measures with finite first moments, the Wasserstein $1$ distance metrizes weak convergence. A comprehensive overview of the advantages of this distance over other measures that metrize weak convergence can be found in (Villani, 2008, p. 98 f.). Due to its compatibility with dimension reduction settings such as the manifold hypothesis, the Wasserstein $1$ distance is a popular choice for the evaluation of generative models, see e.g. Schreuder et al. (2021); Liang (2021); Stéphanovitch et al. (2024); Kunkel and Trabs (2025).
Another popular measure of dissimiliarity is the total variation distance,
$\operatorname{TV}(\mathbb{P},\mathbb{Q})\coloneqq\sup_{A\subset\mathcal{A}}|\mathbb{P}(A)-\mathbb{Q}(A)|,$
where $\mathcal{A}$ is the Borel-$\sigma$ algebra. This metric has frequently been used to evaluate diffusion models (Oko et al., 2023; Yakovlev and Puchkin, 2025). Example˜3.1 illustrates why the total variation distance is not suitable to evaluate Flow Matching models.
Example 3.1.
Let $d=1,Z\sim\mathcal{N}(0,1),\varepsilon>0$ and define the two vector fields $v^{(1)},v^{(2)}\colon[0,1]\times\mathbb{R}\rightarrow\mathbb{R}$
$v^{(1)}_{t}(x)\coloneqq 0,\quad v^{(2)}_{t}(x)\coloneqq\varepsilon\sin\Big(\frac{x}{\varepsilon}\Big),\quad\text{for all }t\in[0,1].$
Let $\psi^{(1)},\psi^{(2)}$ be the corresponding solutions to the ODE (3). Then
$\|v^{(1)}-v^{(2)}\|_{\infty}\overset{\varepsilon\rightarrow 0}{\rightarrow}0,\quad\mathsf{W}_{1}(\mathbb{P}^{\psi_{1}^{(1)}(Z)},\mathbb{P}^{\psi_{2}^{(1)}(Z)})\leq\varepsilon\overset{\varepsilon\rightarrow 0}{\rightarrow}0\quad\text{and}\quad\underset{\varepsilon\rightarrow 0}{\liminf}\operatorname{TV}(\mathbb{P}^{\psi_{1}^{(1)}(Z)},\mathbb{P}^{\psi_{2}^{(1)}(Z)})\geq c,$
for a constant $c>0$ .
The proof of Example˜3.1 shows that for convergence in total variation distance, one needs control over the gradient of $\psi$ with regard to $x$. Considering the change of variables theorem and the fact that the total variation distance operates on a density level, this is unsurprising. However, the gradient is not controlled by the Flow Matching objective. Assuming gradient control over the vector field in space can lead to different results, see Su et al. (2025).
The following theorem provides a first comparison of the performance of $\mathbb{P}^{\hat{\psi}_{1}(Z)}$ and the empirical flow $\mathbb{P}^{\psi^{n}_{1}(Z)}$ in the Wasserstein $1$ distance.
Theorem 3.2.
Assume that all functions in $\mathcal{N}$ are Lipschitz continuous for fixed $t$ with Lipschitz constant $\Gamma_{t}.$ Then for any $\tilde{v}\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d},$ with $\tilde{v}\in\mathcal{N}$ and $v^{n}$ from (9)
$\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{\psi}_{1}(Z)})\leq\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})+\sqrt{2e}e^{\int_{0}^{1}\Gamma_{t}\;\mathrm{d}t}\|v^{n}-\tilde{v}\|_{\infty}$. (16)
Using the construction from (12) to (14), $\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})$ is the estimation error of a kernel density estimator. The second term $\|v^{n}-\hat{v}\|_{\infty}$ depends on the set $\mathcal{N}$ and its ability to approximate $v^{n}$. For a sufficiently rich class $\mathcal{M}$ this second term will be negligible in our analysis. A corresponding error decomposition can be shown for the Wasserstein $2$ distance, too.
The proof of Theorem˜3.2 uses Grönwall’s Lemma, which leads to the factor $\exp(\int_{0}^{1}\Gamma_{t}\;\mathrm{d}t)$ in the second term in (16). This is standard in the analysis of ODEs, in context of flow-based generative models see Fukumizu et al. (2025) and Albergo and Vanden-Eijnden (2023). Compared to the SDE setting of diffusion models, there is no such result like Girsanov’s theorem, which is typically used to study these models, following Chen et al. (2023c), and circumvents the exponential dependence on the Lipschitz constant. This greatly complicates the analysis of Flow Matching. Given this dependence, a Lipschitz regularization in the Flow Matching objective seems to be theoretically beneficial. Imposing the assumption that the vector field defined in (7) is Lipschitz continuous implies an extensive restriction on the class of feasible $\mathbb{P}^{*}$, see Kunkel (2025).
Remark 3.3.
A bound for arbitrary Lipschitz $\tilde{v}_{t}$ cannot be better than $\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})$. If $v^{n}_{t}$ is contained in $\mathcal{N},$ then the Picard-Lindelöf theorem yields $\hat{\psi}_{t}=\psi^{n}_{t}$ and hence $\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{\psi}_{1}})=\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}}).$
Due to Theorem˜3.2, we can bound both error terms individually. We start with the error corresponding to the kernel density estimator in Section˜3.1. Subsequently, we bound the second error term for approximation classes $\mathcal{M}$ consisting of feedforward neural networks in Section˜3.2 to obtain our final result.
3.1 Rate of convergence of the KDE in Wasserstein distance
To bound $\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi_{1}^{n}(Z)})$ in Theorem˜3.2, we have to analyze the corresponding kernel density estimator in the Wasserstein $1$ distance. For compact kernels, existing results for the $L_{1}$-error of the KDE can be exploited. However, this can only lead to the suboptimal convergence rate $n^{-\frac{\alpha}{2\alpha+d}}$. Indeed, the Lipschitz regularity of the test function in (15) can be exploited to obtain a faster rate. Additionally, to allow for latent distributions with unbounded support, especially the Gaussian distribution used by Lipman et al. (2023), we need to investigate kernels with unbounded support.
Theorem 3.4.
Assume $d\geq 2,p^{*}\in B_{1,\infty}^{\alpha}\left(\Lambda,\mathbb{R}^{d}\right),\alpha\in(0,1],\Lambda>0$ and $\operatorname{supp}\left(p^{*}\right)$ bounded. Assume $K$, used in the Flow Matching setting via (14), is a nonnegative $d$-dimensional kernel such that
$\int yK(y)\;\mathrm{d}y=0,\quad\text{and}\quad\int|y|^{1+\alpha}K(y)\;\mathrm{d}y<\infty.$
Further assume $K\in C^{\frac{d+2}{2}}$ with $\|D^{k}K\|_{1}\leq C$ for $k\in\mathbb{N}_{0}^{d}$ such that $|k|\leq\frac{d+2}{2}$ and some $C>0$. Then there are $C_{1}$ and $C_{2}$ such that for $v^{n}$ from (9),
$\mathbb{E}\big[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})\big]\leq C_{1}\sigma_{\min}^{1+\alpha}+\frac{C_{2}}{\sqrt{\sigma_{\min}^{d-2}}}\frac{\log n}{\sqrt{n}}.$
For $\sigma_{\min}\sim(n/\log^{2}n)^{-\frac{1}{2\alpha+d}}$ we obtain $\mathbb{E}\big[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})\big]=\mathcal{O}((n/\log^{2}n)^{-\frac{1+\alpha}{2\alpha+d}}).$
The proof separates the bias induced by smoothing and the stochastic error after smoothing. Using the dual form (15) and the relationship between differentiation and convolution, the bound of the bias exploits both, the $\alpha$-smoothness of $\mathbb{P}^{*}$ and the Lipschitz-smoothness of test functions in (15). The the stochastic error is bounded by considering $(K_{\sigma_{\min}}\ast f)$ as the smoothed test-function and then taking advantage of a result by Schreuder (2020). Noting that the proof relies on the dual form of the Wasserstein $1$ distance, a bound for the Wasserstein $2$ distance would require a completly different approach.
The convergence rate $(n/\log^{2}n)^{-\frac{1+\alpha}{2\alpha+d}}$ coincides up to the logarithmic factor with the lower bound by Niles-Weed and Berthet (2022, Theorem 3) and thus the above rate is minimax optimal up to the logarithm. Divol (2022) also considers rates of convergence for kernel density estimators. In case of $d\neq 2,$ he obtains the optimal rate for distributions bounded away from zero. However, the kernels he considers must be smooth radial functions with support bounded in $(0,1)^{d}$, (Divol, 2022, Condition A). Hence, his result does not apply to the Gaussian kernel.
Remark 3.5.
  1. 1.
    The assumption $\|D^{k}K\|_{1}\leq C$ for $k\in\mathbb{N}_{0}^{d}$ and $C>0$ is satisfied by the Gaussian kernel.
  2. 2.
    For kernels that to not satisfy the differentiability condition in Theorem˜3.4, we show in Theorem˜A.1 a rate of convergence of $n^{-\frac{1+\alpha}{2\alpha+2+d}}.$ For nonzero $\alpha$ and $d$ big enough, this rate decays faster than $n^{-\frac{1}{d}},$ the rate of convergence of the empirical measure (Dudley, 1969; Boissard and Gouic, 2014). This result replaces the assumption that $\operatorname{supp}(p^{*})$ is bounded by a moment assumption on $p^{*}.$
  3. 3.
    The second term in Theorem˜3.4 is an upper bound for the expected value of $\mathsf{W}_{1}(K_{\sigma_{\min}}\ast\mathbb{P}_{n},K_{\sigma_{\min}}\ast\mathbb{P}^{*}).$ Results on the convergence of the smooth empirical measure have been discussed in case of the Gaussian kernel. Goldfeld et al. (2020, Proposition 1) obtain a rate in $\mathcal{O}\big(\frac{1}{\sqrt{n\sigma_{\min}^{d}}}\big)$ for subgaussian distributions $\mathbb{P}^{*}$, which coincides with the rate obtained in Theorem˜A.1. Weed (2018) looks at distributions on [1,1][-1,1] and obtains a rate in $\mathcal{O}\big(\frac{1}{\sqrt{n}\sigma_{\min}^{d}}\big).$ Hence the results above improves these results for distributions in $B^{\alpha}_{1,\infty}$ on bounded support.
3.2 Flow Matching with parameterized vector field
If the empirical vector field $v^{n}$ is contained in $\mathcal{N}$, then the second error term in Theorem˜3.2 vanishes and the Flow Matching procedure inherits the optimal rate of convergence from Theorem˜3.4. In practice, (10) is optimized over a class of neural networks to find a good approximate of $v^{n}$. We thus have to take this approximation error into account.
To analyze the ability of a network to approximate $v^{n},$ it is necessary to know the properties of $v^{n}$ for all $t\in[0,1].$ Hence we are going to specify $\sigma$, $\mu$ and a latent distribution for the subsequent analysis.
Assumption 3.6.
  1. 1.
    We consider the following choices of $\sigma\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{>0}$ and $\mu\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}:$
    $\sigma_{t}(X_{i}^{*})=1-(1-\sigma_{\min})t\quad\text{and}\quad\mu_{t}(X_{i}^{*})=tX_{i}^{*}.$
  2. 2.
    We choose $\mathbb{U}=\mathcal{N}(0,I_{d})$ for the latent distribution.
The choice of $\sigma_{t}$ and $\mu_{t}$ in ˜3.6 coincides with the setting of Lipman et al. (2023) in Example II. This is not a unique choice, for other choices used in the literature in similar settings, see (Tong et al., 2024, Table 1). To avoid losing information, if $\mathbb{U}=\mathcal{N}(0,I_{d})$, then the choice of $\mu_{t}$ should satisfy $|\mu_{t}(X_{i}^{*})|\leq|X_{i}^{*}|$. $\sigma_{t}$ should map to $[0,1]$ to avoid unnecessarily increasing the Lipschitz constant $\Gamma_{t}$ of the vector field. Generalizing the linear choice of $\sigma_{t}$ to
$\sigma_{t}(X_{i}^{*})=(1-(1-\sigma_{\min}^{\frac{1}{k}}{})^{k}$
for a $k>0$ has only a negligible effect. The curvature cancels out in the construction of $v_{t}(\cdot|X_{i}^{*}),$ see Lemma˜2.3.
In order to obtain quantitative results for the neural network approximation, we further need to specify the neural networks. We use fully connected feedforward ReLU networks, which enjoy nice approximation properties, see for example Yarotsky (2017); Gühring et al. (2020); Kohler and Langer (2021); Schmidt-Hieber (2020); Suzuki (2019).
Let us fix some notation: define the Rectified Linear Unit (ReLU) activation function, $\phi\colon\mathbb{R}\rightarrow\mathbb{R},\phi(x)=\max(0,x)$. For a vector $v=\left(v_{1},\ldots,v_{p}\right)\in\mathbb{R}^{p}$, we define the shifted ReLU activation function $\phi_{v}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{p}$ as
$\phi_{v}(x)=\left(\phi\left(x_{1}-v_{1}\right),\ldots,\phi\left(x_{p}-v_{p}\right)\right),\quad x=\left(x_{1},\ldots,x_{p}\right)\in\mathbb{R}^{p}.$
For a number $L\in\mathbb{N}$ of hidden layers and a vector $\mathcal{A}=\left(p_{0},p_{1},\ldots,p_{L+1}\right)\in$ $\mathbb{N}^{L+2}$, a neural network of depth $L+1$ and architecture $\mathcal{A}$ is a function of the form
$f:\mathbb{R}^{p_{0}}\rightarrow\mathbb{R}^{p_{L+1}},\quad f(x)=W_{L}\circ\phi_{v_{L}}\circ W_{L-1}\circ\phi_{v_{L-1}}\circ\cdots\circ W_{1}\circ\phi_{v_{1}}\circ W_{0}\circ x,$ (17)
where $W_{i}\in\mathbb{R}^{p_{i+1}\times p_{i}}$ are weight matrices and $v_{i}\in\mathbb{R}^{p_{i}}$ are shift vectors.
Now we want to combine the approximation properties of ReLU networks with our above analysis to evaluate the performance of the Flow Matching mechanism from (11) where $\mathcal{N}=\mathsf{NN}(L,S,\Gamma)$ is a set of ReLU networks $v\colon[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d},(t,x)\mapsto v_{t}(x)$ with a fixed maximal number of layers $L$, at most $S$ nonzero weights and Lipschitz constant of $v_{t}$ of at most $\Gamma$ for any $t$. We obtain the following rate:
Theorem 3.7.
Grant Assumption 3.6 and assume $p^{*}\in B^{\alpha}_{1,\infty}(\Lambda,[-1,1]^{d})$ for $\alpha\in(0,1],\Lambda>0$. Set $\sigma_{\min}=n^{-\frac{1}{2\alpha+d}}.$ Then we have for $n$ big enough
$\mathbb{E}\big[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{\psi}_{1}})\big]\lesssim n^{-\frac{1+\alpha}{2\alpha+d}}\log^{2}(n),$
where $\hat{\psi}$ is given by (11) with $\mathcal{N}=\mathsf{NN}(L_{n},S_{n},\Gamma_{n})$ where $L_{n}\geq O(n^{\ell})$, $S_{n}\geq O(\exp(n^{\ell}))$ and $\Gamma_{n}=O(n^{\frac{3}{2\alpha+d}})$ for some $\ell>0$.
The proof of Theorem˜3.7 combines Theorem˜3.2 with Theorem˜3.4, the tail behavior of the Gaussian distribution and an approximation result of Gühring et al. (2020). In a nutshell, we approximate the vector field $v^{n}$ on a growing but bounded set with an accuracy that compensates the exponential dependence on the Lipschitz constant. Given a Wasserstein $2$ version of Theorem˜3.4, the same approach could be used the extend the result to $\mathbb{E}[\mathsf{W}_{2}(\mathbb{P}^{*},\mathbb{P}^{\hat{\psi}_{1}(Z)})].$
Remark 3.8.
  1. 1.
    The Lipschitz constraint in Theorem˜3.7 is typically not enforced in practice. However, we only require a bound on the Lipschitz constant of order $\Gamma_{n}=\frac{2d+1}{\sigma_{t}^{3}}+\frac{1}{2}$ which is only a mild restriction for $\sigma_{\min}\to 0$ as $n\to\infty$.
  2. 2.
    The network size in Theorem˜3.7 is not bounded from above. The empirical reference measure $\mathbb{P}^{\psi_{n}(Z)}$ in Theorem˜3.2 circumvents the need for concentration inequalities, which in turn avoids complexity bounds. Thus there is no trade-off in approximation power and network size.
  3. 3.
    Combining Theorem˜3.4 with Theorem˜3.2 and the proof strategy used in Theorem˜3.7 leads to a rate of convergence of $n^{-\frac{1+\alpha}{2\alpha+2+d}}$ for $p^{*}$ with unbounded support that satisfy a moment condition.
  4. 4.
    Compared to the literature on diffusion models and the first results on Flow Matching, we do not use early stopping times, but rather set $\sigma_{\min}>0$ according to $n.$ In terms of variance, both procedures are interchangeable: calculating $\sigma_{t_{*}}$ in (Fukumizu et al., 2025, Theorem 9) leads to exactly the same variance. However, early stopping leads to a bias induced by the mean function $\mu$, which is not $1$ at a time $t<1.$ This could be fixed using an adapted mean function, which is $1$ at the time of the early stopping, but this leads to much more complicated structures of $\mu.$
Theorem˜3.7 shows that Flow Matching achieves minimax optimal rates (up to logarithmic factors) for certain classes of unknown densities, giving some justification of their the empirical success. There are several limitations and potential avenues for further research in the results of this work.
First of all the networks used for Theorem˜3.7 are very large and do not correspond to the networks used in practice. While theoretical worst-case error bounds will always lead to networks that are not exactly small, the work of Yakovlev and Puchkin (2025) gives hope for improvement. Up to now it is an open question, whether a smaller network would come at the cost of a non-optimal rate. In addition, the size of the network and the overall result could benefit from the theoretical use of more sophisticated networks, like the U-net (Ronneberger et al., 2015) construction in Lipman et al. (2023).
We have focused on the statistical analysis of Flow Matching, neglecting the optimization process as a source of additional error as well as the numerical error for solving the ODE (11). Including these errors would be more in line with the real world. The same applies for including the sampling of $t\sim\mathbb{U}$ in the Flow Matching objective.
3.3 Density estimation on a manifold
The rate in Theorem˜3.7 depends on the dimension $d$ and therefore exhibits the classical curse of dimensionality. In this subsection, we analyze a case, in which the curse of dimensionality can be circumvented and the rate depends on some intrinsic dimension $d^{\prime}<d$.
To this end, we consider a single-chart $d^{\prime}$-dimensional manifold $\mathcal{M}$, i.e. there exists a smooth, invertible map $g\colon\mathbb{R}^{d^{\prime}}\rightarrow\mathbb{R}^{d}$ such that
$\mathcal{M}=g(\mathbb{R}^{d^{\prime}}).$
This setting is commonly assumed in generative modeling, specifically in GANs, see for example Schreuder et al. (2021). For our analysis, we impose the following assumption on the smoothness of $g$.
Assumption 3.9.
Let $g\in C^{2}$ and $\|g\|_{\infty}\leq M,\|Dg(y)\|\leq M_{1},\|D(Dg(y))\|\leq M_{2}$. The Jacobian $Dg$ is such that for every $y\in\mathbb{R}^{d^{\prime}}$, $0<m<\lambda_{\min}(Dg(y))<\lambda_{\max}(Dg(y))<M_{3}$, where $\lambda_{\min}$ is the smallest and $\lambda_{\max}$ is the largest eigenvalue and $m,M_{3}>0$ are constants. The inverse $g^{-1}$ of $g$ is twice continuously differentiable.
We further assume that $\mathcal{M}$ has a positive reach $\tau>0$. The notion of the reach of a manifold is due to Federer (1959). First, we define the distance function to $\mathcal{M}$ for $x\in\mathbb{R}^{d}$ as $\operatorname{dist}(x,M)=\inf_{z\in\mathcal{M}}|x-z|$. Then, the medial axis of $\mathcal{M}$ is given by
$\operatorname{Med}(\mathcal{M})=\{x\in\mathbb{R}^{d}\mid\exists\ y,z\in\mathcal{M},y\neq z,|y-x|=|z-x|=\operatorname{dist}(x,\mathcal{M})\}.$
The reach of $\mathcal{M}$ is then defined as
$\tau=\inf_{x\in\mathcal{M}}\operatorname{dist}(x,\operatorname{Med}(\mathcal{M})).$
For an illustration of the concept of the reach see Figure˜4 in Section˜5.3. For $r<\tau$ and $r<c_{0}=\frac{2m^{2}\tau}{2M_{2}\tau+3mM_{1}}$, let $\pi\colon\mathbb{R}^{d}\rightarrow\mathcal{M}\cup\{0\}$ be a function such that $\pi\big|_{B(r,\mathcal{M})}$ is the orthogonal projection onto $\mathcal{M}$ and $\pi(B(r,\mathcal{M})^{c})=\{0\}$. In this subsection, we assume that $\mathbb{P}^{*}$ is supported on $\mathcal{M}$ and that there is a $\alpha$-Besov smooth distribution $\mathbb{P}^{*}_{d^{\prime}}$ on $\mathbb{R}^{d^{\prime}}$ such that $g(Y)\sim\mathbb{P}^{*}$ when $Y\sim\mathbb{P}^{*}_{d^{\prime}}$. The next theorem bounds the estimation error of the KDE in a proximity of the manifold.
Theorem 3.10.
Assume $\mathbb{P}^{*}_{d^{\prime}}$ has a Lebesgue density $p^{*}_{d^{\prime}}\in B^{\alpha}_{1,\infty}(\Lambda,[-1,1]^{d^{\prime}})$ for $\alpha\in(0,1],\Lambda>0$. Further assume that the kernel $K$ satisfies the assumptions of Theorem˜3.4. Then
$\mathbb{E}\big[\mathsf{W}_{1}(\mathbb{P}^{*},\pi_{\#}\mathbb{P}^{\psi^{n}_{1}(Z)})\big]\leq C_{1}\sigma_{\min}^{1+\alpha}+\frac{C_{2}}{\sqrt{\sigma_{\min}^{d^{\prime}-2}}}\frac{\log n}{\sqrt{n}}.$
For $\sigma_{\min}\sim(n/\log^{2}(n))^{-\frac{1}{2\alpha+d^{\prime}}}$ we obtain $\mathbb{E}\big[\mathsf{W}_{1}(\mathbb{P}^{*},\pi_{\#}\mathbb{P}^{\psi^{n}_{1}}(Z))\big]=\mathcal{O}\big((n/\log^{2}(n))^{-\frac{1+\alpha}{2\alpha+d^{\prime}}}\big)$.
The proof of Theorem˜3.10 is inspired and uses techniques from Aamari et al. (2019) and Genovese et al. (2014). The proof starts with a decomposition of the projection into tangential vectors and a remainder. The remainder is bounded using a Taylor series and the reach assumption. The rest of the proof follows along the lines of Theorem˜3.4, but is carefully adapted to the manifold and projection setting. While the rigorous arguments are very technical, in a nutshell, the tangential noise is retraced to $\mathbb{R}^{d^{\prime}}$ and the space dependence of this retraction is shown to be small. Just as Theorem˜3.4, the proof strongly relies on the dual form of the Wasserstein $1$ metric, thus extensions to Wasserstein $2$ are not straightforward.
The above result is in line with the similar results for kernel density estimators for distributions on manifolds, c.f. Berenfeld and Hoffmann (2021); Divol (2022). Note however that in these articles, the kernel is normalized with respect to the lower dimension, i.e. they consider $h^{-d^{\prime}}K(\cdot/h)$ for bandwidth $h>0$, while our generative setting enforces us to prove Theorem˜3.10 for kernels which are normalized with respect to the ambient space dimension, i.e. $h^{-d}K(\cdot/h)$ (and $h=\sigma_{\min}$ in our notation). Furthermore, in Divol (2022, Theorem 3.1), the kernel estimator is evaluated on the manifold. Since the Flow Matching estimator results in generated samples and we evaluate the distribution of these samples, we cannot measure the Wasserstein distance along the manifold without the tube-projection, since the density of a distribution restricted to a null-set is not unique. This is different from the density estimation settings, where the density is either constructed directly on the manifold (Divol, 2022) or evaluated pointwise (Berenfeld and Hoffmann, 2021). Section˜4 will give another motivation for this evaluation measure.
Theorem 3.11.
Grant Assumption 3.6 and assume the setting of Theorem˜3.10. Set $\sigma_{\min}=n^{-\frac{1}{2\alpha+d^{\prime}}}.$ Then there are sequences $L_{n},S_{n},\Gamma_{n}\in\mathbb{N}$ such that for $n$ big enough
$\mathbb{E}\big[\mathsf{W}_{1}(\mathbb{P}^{*},\pi_{\#}\mathbb{P}^{\hat{\psi}_{1}(Z)})\big]\lesssim n^{-\frac{1+\alpha}{2\alpha+d^{\prime}}}\log(n)^{2},$
where $\hat{\psi}$ is given by (11) with $\mathcal{N}=\mathsf{NN}(L_{n},M_{n},\Gamma_{n})$.
Theorem˜3.11 shows that the Flow Matching estimator can approximate $\mathbb{P}^{*}$ with a faster rate on a small tube around the manifold, exploiting the smoothness of $\mathbb{P}^{*}$ on the manifold. It uses exactly the same argument as Theorem˜3.7. For diffusion models, similar behavior has been observed by Tang and Yang (2024); Azangulov et al. (2024).
4 Numerical illustration
While the empirical success is widely known, the main argument of this work is that in an overparameterized setting the Flow Matching algorithm is at least as good as the kernel density estimator. In a full dimensional setting, we still recover a minimax optimal rate of convergence. Particularly, the Flow Matching algorithm does not mimic the empirical measure, but a smoothed version. This enables leveraging smoothness, even in an overparameterized setting. In this section, we want to illustrate this comparison using a constructed example.
To this end, we consider the following setting: the distribution $\mathbb{P}^{*}$ is the uniform distribution on the graph of the sine curve restricted to [3,3][-3,3]. Thus, $\mathbb{P}^{*}$ lies on a $1$-dimensional manifold in $\mathbb{R}^{2}$. We assume that we have $200$ samples from $\mathbb{P}^{*}$. In this setting, we train a Flow Matching model with three hidden SeLU layers of width $512$ and $\mathbb{U}=\mathcal{N}(0,I_{2})$, for a definition of the SeLU activation function see Klambauer et al. (2017). Our implementation builds up on the code provided by Tong et al. (2024), we use the Adam optimizer (Kingma and Ba, 2014) with the standard parameters $(\operatorname{lr}=0.001$, $\beta_{1}=0.9$ and $\beta_{2}=0.999)$ and Poli et al. (2025) to solve the neural ODE with the solver dopri5 and set $\mathrm{atol}=\mathrm{rtol}=10^{-5}$. In order to analyse the model without avoidable additional smoothing, we do not use minibatches. We evaluated the estimator after 10 000,20 000,30 00010\ 000,20\ 000,30\ 000 and 40 00040\ 000 updates to avoid committing to a fixed number of training iterations. As a comparison, we use a KDE with a Gaussian kernel. For both methods, we generate $512$ samples. To interpret the KDE as a generative model, we draw an index uniformly, sample from $\mathcal{N}(0,I_{2})$ and then add the noise to the sample of the index. An example of this setting and generated samples from the generative models for $\sigma_{\min}\in\{0.5,0.1,0.05,0.01\}$ can be found in Figure˜1. We repeated the following experiments for networks with width of $32$ and $8$, the results can be found in Figure˜5 in Appendix˜A.
(a) FM, $\sigma_{\min}=0.5$
(b) FM, $\sigma_{\min}=0.1$
(c) FM, $\sigma_{\min}=0.05$
(d) FM, $\sigma_{\min}=0.01$
(e) KDE, $\sigma_{\min}=0.5$
(f) KDE, $\sigma_{\min}=0.1$
(g) KDE, $\sigma_{\min}=0.05$
(h) KDE, $\sigma_{\min}=0.01$
Figure 1: Exemplarily result of the Flow Matching (FM) estimator and the KDE. The generated samples are the blue dots, the training sample is represented by black dots on the red sine curve. Top row shows generated samples of the FM estimator after $40000$ runs.
By construction, we know that the KDE does not detect the manifold, but spreads noise in two dimensions. The question is whether the Flow Matching estimator is capable of detecting the manifold, without imposing structural assumptions on the manifold. To measure this, we calculate the mean distance of the generated samples to the manifold. We repeated the same experiment independently drawing new samples $10$ times. The results are gathered in Figure˜2(a). For all choices of $\sigma_{\min}$ except the largest, the KDE achieves the smallest mean distance, which strongly opposes the claim that the Flow Matching estimator detects the manifold better that the KDE.
Naturally, for very small $\sigma_{\min}$, the KDE is very close to the empirical distribution. In proofs, i.e. the proof of Theorem˜3.4 only a proper choice of $\sigma_{\min}$ leads an the optimal rate. The Flow Matching estimator does not behave the same, the model seems to smooth intrinsically. Albeit we ran the model for a very large number of training iterations, there is no mechanism to prevent the convergence towards a local minimum. Therefore this effect could be due to the optimization error. More desirable is an intrinsic smoothing. In order to measure this effect, we calculated the maximal distance between two projections on the sine curves in the $10$ runs. In Figure˜2(b) it is visible that for smaller $\sigma_{\min}$, the Flow Matching estimator approximately covers the manifold better than then KDE, whereas this cannot be seen for larger $\sigma_{\min}$. While for the uniform distribution that is considered here, gaps in the covering are a satisfactory indication for the quality of estimation, this measure fails for other distributions. In hindsight, this motivates the projected measure evaluated in Theorem˜3.10.
(a) Mean distance of the generated samples to the manifold. Note that the shape of the sine curve leads to a smaller expected distance than $\sigma_{\min}$, this effect is smaller for smaller $\sigma_{\min}$.
(b) Largest gap on the manifold between the projections of the generated samples.
Figure 2: Comparison between the Flow Matching estimator and the KDE in terms of distance to the manifold and the largest gap of the projection estimator on the manifold.
Our results indicate that in order to theoretically find situations in which the Flow Matching algorithm outperforms the KDE in very general settings without structural assumptions on the unknown distribution, classical distances between probability measures might not the right choice. While creating an image that is different from the training images appears as creative and is favorable, the performance of the corresponding distribution in the Wasserstein $1$ distance is not necessarily good. In context of GANs, an inclusion of the distance between the generated distribution and the empirical measure has recently been investigated by Vardanyan et al. (2024).
5 Proofs
Before we begin with the proofs, let us introduce some additional notation: The spectral norm of $A$ is denoted by $\|A\|_{2}.$ We define the ceiling of $x\in\mathbb{R}$ as $\lceil x\rceil:=\min\{k\in\mathbb{Z}|k\geq x\}.$
For $\alpha\in(0,1]$, we define the Hölder space $\mathcal{H}^{\alpha}$ by
$\mathcal{H}^{\alpha}(\Omega)\coloneqq\big\{f\colon\Omega\rightarrow\mathbb{R}^{d}\colon\|f\|_{\mathcal{H}^{\alpha}(\Omega)}<\infty\big\},\quad\text{ where }\quad\|f\|_{\mathcal{H}^{\alpha}(\Omega)}\coloneqq\sup_{x,y\in\Omega}\frac{|f(x)-f(y)|}{|x-y|^{\alpha}}.$ (18)
For $\alpha=1,$ we obtain the space of Lipschitz functions. We abbreviate $\operatorname{Lip}(f)=\|f\|_{\mathcal{H}^{1}(\Omega)}.$ Again, in case of $\Omega=\mathbb{R}^{d},$ we omit $\Omega.$
5.1 Proofs of Section˜2
Proof of Lemma˜2.1.
A necessary and sufficient condition for $v_{t}$ to generate $p_{t}$ is given in (4). We thus verify
$\frac{d}{dt}p^{n}_{t}(x)$ $=\frac{1}{n}\sum_{i=1}^{n}\frac{d}{dt}p_{t}(x|X_{i}^{*})=-\frac{1}{n}\sum_{i=1}^{n}\operatorname{div}(p_{t}(x|X_{i}^{*})v_{t}(x|X_{i}^{*}))$
$=-\operatorname{div}\Big(\frac{1}{n}\sum_{i=1}^{n}p_{t}(x|X_{i}^{*})\sum_{i=1}^{n}\frac{p_{t}(x|X_{i}^{*})v_{t}(x|X_{i}^{*})}{\sum_{i=1}^{n}p_{t}(x|X_{i}^{*})}\Big)=-\operatorname{div}(p^{n}_{t}(x)v^{n}_{t}(x)).\qed$
Proof of Theorem˜2.2.
For fixed $t\in[0,1]$ we have
$|\hat{v}_{t}(x)-v^{n}_{t}(x)|^{2}$ $=|\hat{v}_{t}(x)|^{2}-2\langle\hat{v}_{t}(x),v^{n}_{t}(x)\rangle+|v^{n}_{t}(x)|^{2},$
$|\hat{v}_{t}(x)-v_{t}(x|X_{i}^{*})|^{2}$ $=|\hat{v}_{t}(x)|^{2}-2\langle\hat{v}_{t}(x),v_{t}(x|X_{i}^{*})\rangle+|v_{t}(x|X_{i}^{*})|^{2}.$
The last term does not influence the minimal argument in $\hat{v}$. For the first two we have
$\mathbb{E}_{X_{t}\sim p^{n}_{t}}[|\hat{v}_{t}(X_{t})|^{2}]=\int|\hat{v}_{t}(x)|^{2}p^{n}_{t}(x)\;\mathrm{d}x=\frac{1}{n}\sum_{i=1}^{n}\int|\hat{v}_{t}(x)|^{2}p_{t}(x|X_{i}^{*})\;\mathrm{d}x=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\tilde{X}_{t}\sim p_{t}(\cdot|X_{i}^{*})}[|\hat{v}_{t}(\tilde{X}_{t})|^{2}],$
and
$\mathbb{E}_{X_{t}\sim p^{n}_{t}}[\langle\hat{v}_{t}(X_{t}),v^{n}_{t}(X_{t})\rangle]$ $=\int\langle\hat{v}_{t}(x),v^{n}_{t}(x)\rangle p^{n}_{t}(x)\;\mathrm{d}x$
$=\int\Big\langle\hat{v}_{t}(x),\sum_{i=1}^{n}v_{t}(x|X_{i}^{*})\frac{p_{t}(x|X_{i}^{*})}{\sum_{i=1}^{n}p_{t}(x|X_{i}^{*})}\Big\rangle\frac{1}{n}\sum_{i=1}^{n}p_{t}(x|X_{i}^{*})\;\mathrm{d}x$
$=\frac{1}{n}\sum_{i=1}^{n}\int\langle\hat{v}_{t}(x),v_{t}(x|X_{i}^{*})\rangle p_{t}(x|X_{i}^{*})\;\mathrm{d}x$
$=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\tilde{X}_{t}\sim p_{t}(\cdot|X_{i}^{*})}[\langle\hat{v}_{t}(\tilde{X}_{t}),v_{t}(\tilde{X}_{t}|X_{i}^{*})\rangle].\qed$
Proof of Lemma˜2.3.
As $\sigma_{t}(X_{i}^{*})>0$ for all $t,X_{i}^{*},$ we have that
$\psi_{t}^{-1}(x|X_{i}^{*})=\frac{x-\mu_{t}(X_{i}^{*})}{\sigma_{t}(X_{i}^{*})}.$
Hence
$\frac{\partial\psi_{t}}{\partial t}(z|X_{i}^{*})\Big|_{z=\psi_{t}^{-1}(x|X_{i}^{*})}=v_{t}(x|X_{i}^{*}),$
and
$\frac{\partial\psi_{t}}{\partial t}(x|X_{i}^{*})=\frac{\partial\sigma_{t}}{\partial t}(X_{i}^{*})x+\frac{\partial\mu_{t}}{\partial t}(X_{i}^{*}).$
Thus, we get
$v_{t}\left(x|X_{i}^{*}\right)=\frac{\frac{\partial\sigma_{t}}{\partial t}\left(X_{i}^{*}\right)}{\sigma_{t}\left(X_{i}^{*}\right)}\big(x-\mu_{t}\left(X_{i}^{*}\right)\big)+\frac{\partial\mu_{t}}{\partial t}\left(X_{i}^{*}\right).\qed$
5.2 Proofs and extended results of Section˜3
Proof of Example˜3.1.
For the Wasserstein $1$ distance, we can bound
$\mathsf{W}_{1}(\mathbb{P}^{\psi_{1}^{(1)}(Z)},\mathbb{P}^{\psi_{2}^{(1)}(Z)})$ $\leq\mathbb{E}_{Z\sim\mathcal{N}(0,1)}[|\psi_{1}^{(1)}(Z)-\psi_{1}^{(2)}(Z)|]$
$=\mathbb{E}_{Z\sim\mathcal{N}(0,1)}[|Z-\psi_{1}^{(2)}(Z)|]$
$=\mathbb{E}_{Z\sim\mathcal{N}(0,1)}\Big[\Big|\int_{0}^{1}v_{t}^{(2)}(\psi^{(2)}_{t}(Z))\;\mathrm{d}t\Big|\Big]$
$=\mathbb{E}_{Z\sim\mathcal{N}(0,1)}[|\varepsilon\sin\Big(\frac{\psi^{(2)}_{t}(Z)}{\varepsilon}\Big)|]$
$\leq\varepsilon.$
For the total variation distance, we first note that by the change of variables theorem, we have for every element of a sequence of diffeomorophisms $(f_{k})_{k\in\mathbb{N}}$
$f_{k}(Z)\sim\frac{\varphi(f_{k}^{-1}(\cdot))}{|f_{k}^{\prime}(f_{k}^{-1}(\cdot))|},$
where $\varphi$ is the density of $\mathcal{N}(0,1)$ and where we identified the density with the distribution. By definition of $v^{(1)}$, we have that $\psi^{(1)}_{t}(Z)\sim\varphi$. In case $f_{k}^{\prime}(x)>0$ for all $x\in\mathbb{R}$, we obtain for the total variation distance
$\operatorname{TV}(\mathcal{N}(0,1),\mathbb{P}^{f_{k}(Z)})$ $=\frac{1}{2}\int\Big|\varphi(x)-\frac{\varphi(f_{k}^{-1}(\cdot))}{|f^{\prime}(f_{k}^{-1}(\cdot))|}\Big|\;\mathrm{d}x$
$=\frac{1}{2}\int|\varphi(f_{k}(x))f_{k}^{\prime}(x)-\varphi(x)|\;\mathrm{d}x$
$=\frac{1}{2}\int|\varphi(x)(f_{k}^{\prime}(x)-1)+f_{k}^{\prime}(x)(\varphi(f_{k}(x))-\varphi(x))|\;\mathrm{d}x$
$\geq\frac{1}{2}\int|\varphi(x)(f_{k}^{\prime}(x)-1)|\;\mathrm{d}x-\frac{1}{2}\int|f_{k}^{\prime}(x)(\varphi(f_{k}(x))-\varphi(x))|\;\mathrm{d}x.$ (19)
In case $\sup_{k\in\mathbb{N}}\|f_{k}^{\prime}\|_{\infty}\leq C<\infty$ and $\|f_{k}-\operatorname{Id}\|_{\infty}\rightarrow 0$, we have that for $k$ large enough
$|f_{k}(x)-x|\leq 1\quad\Longrightarrow\quad|f_{k}(x)|\geq|x|-1,$
which implies
$|f_{k}^{\prime}(x)(\varphi(f_{k}(x))-\varphi(x))|\leq C\Big(\frac{1}{\sqrt{2\pi}}e^{-\frac{(|x|-1)^{2}}{2}}+\varphi(x)\Big).$
Since the right hand side of the above equation is integrable, we can use dominated convergence and the continuity of $\varphi$ to conclude that the second term in (19) converges towards 0 for $k\rightarrow\infty$. However, for the first term, we also need that $f^{\prime}\rightarrow(\operatorname{Id})^{\prime}=1$. In the Flow Matching setting, this corresponds to the derivative of $\psi^{(2)}_{1}$ with respect to $x$ converging to $1$, which is not controlled via the Flow Matching objective.
In the given example, we can solve the corresponding ODEs and obtain
$\frac{\partial\psi^{(1)}_{t}(x)}{\partial t}=0,\quad$ $\psi^{(1)}_{0}(x)=x,$ $\Longrightarrow\quad$ $\psi^{(1)}_{t}(x)=x$
$\frac{\partial\psi^{(2)}_{t}(x)}{\partial t}=\varepsilon\sin\Big(\frac{\psi^{(2)}_{t}(x)}{\varepsilon}\Big),\quad$ $\psi^{(2)}_{0}(x)=x,$ $\Longrightarrow\quad$ $\psi^{(2)}_{t}(x)=2\varepsilon\arctan\Big(e^{t}\tan\Big(\frac{x}{2\varepsilon}\Big)\Big)+2\pi\varepsilon\Big\lfloor\frac{x}{2\pi\varepsilon}+\frac{1}{2}\Big\rfloor,$
where $\psi^{(2)}_{t}$ is defined for almost all $x\in\mathbb{R}$. Note that $\big\lfloor\frac{x}{2\pi\varepsilon}+\frac{1}{2}\big\rfloor$ is chosen such that we obtain a continuous, differentiable and bijective function on $\mathbb{R}$. The solution of $\psi^{(2)}$ can be checked calculating the derivative with respect to $t$ and using $\sin(2\theta)=\frac{2\tan(\theta)}{1+\tan^{2}(\theta)}$. The derivative with respect to $x$ of $\psi_{t}^{(2)}$ is given by
$\frac{\partial\psi^{(2)}_{t}}{\partial x}=\frac{e^{t}\sec^{2}\big(\frac{x}{2\varepsilon}\big)}{e^{2t}\tan^{2}\big(\frac{x}{2\varepsilon}\big)+1}=\frac{e^{t}}{\big(e^{2t}-1\big)\sin^{2}\big(\frac{x}{2\varepsilon}\big)+1}>0.$
Since $t\in[0,1]$, the derivative is bounded. Therefore, we can use the lower bound from (19). By construction $\psi_{1}^{(2)}(x)\rightarrow x$ for $\varepsilon\rightarrow 0$, thus the second term in (19) converges towards 0. For the second term
$\int\Big|$ $\varphi(x)\Big(\frac{e}{\big(e^{2}-1\big)\sin^{2}\big(\frac{x}{2\varepsilon}\big)+1}-1\Big)\Big|\;\mathrm{d}x$
$\geq\frac{1}{\sqrt{2\pi e}}\int_{-1}^{1}\Big|\Big(\frac{e}{\big(e^{2}-1\big)\sin^{2}\big(\frac{x}{2\varepsilon}\big)+1}-1\Big)\Big|\;\mathrm{d}x$
$\geq\frac{1}{2\sqrt{2\pi e}}\int_{-1}^{1}\mathds{1}\Big\{\frac{e}{\big(e^{2}-1\big)\sin^{2}\big(\frac{x}{2\varepsilon}\big)+1}\leq\frac{1}{2}\Big\}\;\mathrm{d}x$
$=\frac{1}{2\sqrt{2\pi e}}\int_{-1}^{1}\mathds{1}\Big\{x\in\bigcup_{n\in\mathbb{Z}}\Big[2\varepsilon\Big(\pi n+\arcsin\Big(\sqrt{\frac{2e-1}{e^{2}-1}}\Big)\Big),2\varepsilon\Big(\pi(n+1)-\arcsin\Big(\sqrt{\frac{2e-1}{e^{2}-1}}\Big)\Big)\Big]\Big\}\;\mathrm{d}x.$
Denoting $\gamma\coloneqq\arcsin\Big(\sqrt{\frac{2e-1}{e^{2}-1}}\Big)$, the length of the intervals in the bound is $2\varepsilon((n+1)\pi-\gamma)-2\varepsilon(n\pi+\gamma)=2\varepsilon(\pi-2\gamma).$ Next we count how many intervals are lie completely in [1,1][-1,1]. We have that
$2\varepsilon(\pi n+\gamma)\geq-1\Longleftrightarrow n\geq-\frac{1}{2\varepsilon\pi}-\frac{\gamma}{\pi},\quad 2\varepsilon(\pi(n+1)-\gamma)\leq 1\Longleftrightarrow n\leq\frac{1}{2\varepsilon\pi}+\frac{\gamma}{\pi}-1.$
Therefore, we have
$\Big\lfloor\frac{1}{2\varepsilon\pi}+\frac{\gamma}{\pi}-1\Big\rfloor-\Big\lceil-\frac{1}{2\varepsilon\pi}-\frac{\gamma}{\pi}\Big\rceil$
intervals that lie completely in [1,1][-1,1]. We bound this number via
$\Big\lfloor\frac{1}{2\varepsilon\pi}+\frac{\gamma}{\pi}-1\Big\rfloor-\Big\lceil-\frac{1}{2\varepsilon\pi}-\frac{\gamma}{\pi}\Big\rceil\geq\frac{1}{2\varepsilon\pi}+\frac{\gamma}{\pi}-2-(l-\frac{1}{2\varepsilon\pi}-\frac{\gamma}{\pi}+1)=\frac{1}{\varepsilon\pi}+\frac{\gamma}{\pi}-3.$
Thus, we can bound
$\int\Big|\varphi(x)\Big(\frac{e}{\big(e^{2}-1\big)\sin^{2}\big(\frac{x}{2\varepsilon}\big)+1}-1\Big)\Big|\;\mathrm{d}x$ $\geq\frac{1}{2\sqrt{2\pi e}}2\varepsilon(\pi-2\gamma)\big(\frac{1}{\varepsilon\pi}+\frac{\gamma}{\pi}-3\big)$
$=\frac{1}{2\sqrt{2\pi e}}\Big(\frac{2}{\pi}(\pi-2\gamma)+2\varepsilon\gamma\frac{\pi-2\gamma}{\pi}-6\varepsilon(\pi-2\gamma)\Big)$
$\overset{\varepsilon\rightarrow 0}{\longrightarrow}\frac{\pi-2\gamma}{\pi\sqrt{2\pi e}}.$
Setting $0<c<\frac{\pi-2\gamma}{\pi\sqrt{2\pi e}}$ yields the result. ∎
Proof of Theorem˜3.2.
As $\mathsf{W}_{1}$ and $\mathsf{W}_{2}$ satisfy the triangle inequality (see Villani (2008, p. 94)),
$\mathsf{W}_{q}(\mathbb{P}^{*},\mathbb{P}^{\hat{\psi}_{1}(Z)})\leq\mathsf{W}_{q}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})+\mathsf{W}_{q}(\mathbb{P}^{\psi^{n}_{1}(Z)},\mathbb{P}^{\hat{\psi}_{1}(Z)}).$
Using (Villani, 2008, Remark 6.5) we get
$\mathsf{W}_{1}(\mathbb{P}^{\psi^{n}_{1}(Z)},\mathbb{P}^{\hat{\psi}_{1}(Z)})$ $\leq\mathsf{W}_{2}(\mathbb{P}^{\psi^{n}_{1}(Z)},\mathbb{P}^{\hat{\psi}_{1}(Z)})$
$\leq\sqrt{\mathbb{E}[|\psi^{n}_{1}(Z)-\hat{\psi}_{1}(Z)|^{2}]}$
$=\Big(\int|\psi^{n}_{1}(x)-\hat{\psi}_{1}(x)|^{2}p_{0}(x)\;\mathrm{d}x\Big)^{1/2}.$
Like in (Albergo and Vanden-Eijnden, 2023, Proposition 3) set
$Q_{t}=\int|\psi^{n}_{t}(x)-\hat{\psi}_{t}(x)|^{2}p_{0}(x)\;\mathrm{d}x.$
Since $p^{*},\psi,$ $\hat{\psi}$ and $v_{t}^{n}$ are integrable with respect to the Lebesgue measure, we can use the dominated convergence theorem and interchange integration and differentiation to obtain
$\frac{d}{dt}Q_{t}$ $=2\int(\psi^{n}_{t}(x)-\hat{\psi}_{t}(x))^{\top}(v^{n}_{t}(\psi^{n}_{t}(x))-\hat{v}_{t}(\hat{\psi}_{t}(x)))p_{0}(x)\;\mathrm{d}x$
$=2\int(\psi^{n}_{t}(x)-\hat{\psi}_{t}(x))^{\top}(v^{n}_{t}(\psi^{n}_{t}(x))-\hat{v}_{t}(\psi^{n}_{t}(x)))p_{0}(x)\;\mathrm{d}x$
$\quad+2\int(\psi^{n}_{t}(x)-\hat{\psi}_{t}(x))^{\top}(\hat{v}_{t}(\psi^{n}_{t}(x))-\hat{v}_{t}(\hat{\psi}_{t}(x)))p_{0}(x)\;\mathrm{d}x.$
As $0\leq|\psi^{n}_{t}(x)-\hat{\psi}_{t}(x)-(v^{n}_{t}(\psi^{n}_{t}(x))-\hat{v}_{t}(\psi^{n}_{t}(x)))|^{2},$ we get
$2(\psi^{n}_{t}(x)-\hat{\psi}_{t}(x))^{\top}(v^{n}_{t}(\psi^{n}_{t}(x))-\hat{v}_{t}(\psi^{n}_{t}(x)))\leq|\psi^{n}_{t}(x)-\hat{\psi}_{t}(x)|^{2}+|v^{n}_{t}(\psi^{n}_{t}(x))-\hat{v}_{t}(\psi^{n}_{t}(x))|^{2},$
and since for fixed $t$ the vector field $\hat{v}_{t}$ is $\Gamma_{t}$ Lipschitz continuous
$2(\psi^{n}_{t}(x)-\hat{\psi}_{t}(x))^{\top}(\hat{v}_{t}(\psi^{n}_{t}(x))-\hat{v}_{t}(\hat{\psi}_{t}(x)))\leq 2\Gamma_{t}|\psi^{n}_{t}(x)-\hat{\psi}_{t}(x)|^{2}.$
Therefore
$\frac{d}{dt}Q_{t}\leq(1+2\Gamma_{t})Q_{t}+2\int|v^{n}_{t}(\psi^{n}_{t}(x))-\hat{v}_{t}(\psi^{n}_{t}(x))|^{2}p_{0}(x)\;\mathrm{d}x.$
Now we use the formulation of Grönwall’s lemma in Walter (1970) and use $Q_{0}=0$ (this holds since $\psi_{0}(x)=\hat{\psi}_{0}(x)=x$). This leads to
$Q_{1}\leq e^{\int_{0}^{1}1+2\Gamma_{t}\;\mathrm{d}t}2\int_{0}^{1}\int|v^{n}_{t}(\psi^{n}_{t}(x))-\hat{v}_{t}(\psi^{n}_{t}(x))|^{2}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t.$ (20)
By Theorem˜2.2 the vector field $\hat{v}$ is chosen such that it minimizes $\int_{0}^{1}\mathbb{E}_{X_{t}\sim p_{t}}\big[|\hat{v}_{t}(X_{t})-v^{n}_{t}(X_{t})|^{2}\big]\;\mathrm{d}t$ and $X_{t}=\psi_{t}(Z),$ thus we get for every $\tilde{v}\in\mathcal{N}$
$Q_{1}$ $\leq 2e^{\int_{0}^{1}1+2\Gamma_{t}\;\mathrm{d}t}\int_{0}^{1}\int|v^{n}_{t}(\psi^{n}_{t}(x))-\tilde{v}_{t}(\psi^{n}_{t}(x))|^{2}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t$
$\leq 2e^{\int_{0}^{1}1+2\Gamma_{t}\;\mathrm{d}t}\int_{0}^{1}\|v^{n}_{t}-\tilde{v}_{t}\|^{2}_{\infty}p_{0}(x)\;\mathrm{d}t,$
$\leq 2e^{\int_{0}^{1}1+2\Gamma_{t}\;\mathrm{d}t}\|v^{n}-\tilde{v}\|^{2}_{\infty}.$
Taking the square root leads to the result. ∎
Proof of Theorem˜3.4.
Set $\mu_{n}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}}.$ We denote $K_{\sigma_{\min}}\coloneqq\frac{1}{\sigma_{\min}^{d}}K(\frac{\cdot}{\sigma_{\min}}).$ By the triangle inequality of the supremum we get
$\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})$ $=\sup_{f\in\mathrm{Lip}(1)}\int f(z)(p^{*}(z)-(K_{\sigma_{\min}}\ast\mu_{n})(z))\;\mathrm{d}z$
$\leq\sup_{f\in\mathrm{Lip}(1)}\int f(z)(p^{*}(z)-(K_{\sigma_{\min}}\ast p^{*})(z))\;\mathrm{d}z$
$\quad+\sup_{f\in\mathrm{Lip}(1)}\int f(z)((K_{\sigma_{\min}}\ast p^{*})(z)-(K_{\sigma_{\min}}\ast\mu_{n})(z))\;\mathrm{d}z.$
The first term is the bias of the kernel density estimator and the second term refers to the stochastic error.
First term: Using Fubini’s theorem and $\int K_{\sigma_{\min}}(y)\;\mathrm{d}y=1$, we rewrite the first term as
$\sup_{f\in\mathrm{Lip}(1)}\int f(z)(p^{*}(z)$ $-(K_{\sigma_{\min}}\ast p^{*})(z))\;\mathrm{d}z$
$=\sup_{f\in\mathrm{Lip}(1)}\int f(z)\Big(\int K_{\sigma_{\min}}(y)p^{*}(z)\;\mathrm{d}y-\int K_{\sigma_{\min}}(y)p^{*}(z-y)\;\mathrm{d}y\Big)\;\mathrm{d}z$
$=\sup_{f\in\mathrm{Lip}(1)}\int K_{\sigma_{\min}}(y)\Big(\int f(z)p^{*}(z)\;\mathrm{d}z-\int f(z)p^{*}(z-y)\;\mathrm{d}z\Big)\;\mathrm{d}y$
$=\sup_{f\in\mathrm{Lip}(1)}\int K_{\sigma_{\min}}(y)\big((f\ast p^{*}(-\cdot))(0)-(f\ast p^{*}(-\cdot))(y)\big)\;\mathrm{d}y.$
Then, since $|p^{*}|_{B^{\alpha}_{1,\infty}}\leq M,$ we have that
$|\nabla(f\ast p^{*}(-\cdot))(x)-\nabla(f\ast p^{*}(-\cdot))(y)|$ $=\Big|\int\nabla f(z)p^{*}(x-z)\;\mathrm{d}z-\int\nabla f(z)p^{*}(y-z)\;\mathrm{d}z\Big|$
$\leq\sqrt{d}M|x-y|^{\alpha},$
where $\nabla$ denotes the vector of weak partial derivatives and we use that $|\nabla f|\leq\sqrt{d}$ for $f\in\operatorname{Lip}(1)$.
Therefore by the generalized mean value theorem, for every $y$ there exists a $\xi_{y}$ and a $C_{1}^{\prime}>0$ such that
$\sup_{f\in\mathrm{Lip}(1)}\int$ $K_{\sigma_{\min}}(y)\Big(\int f(z)p^{*}(z)\;\mathrm{d}z-\int f(z)p^{*}(z-y)\;\mathrm{d}z\Big)\;\mathrm{d}y$
$=\sup_{f\in\mathrm{Lip}(1)}\int K_{\sigma_{\min}}(y)(\nabla(f\ast p^{*}(-\cdot))(\xi_{y}))^{\top}y\;\mathrm{d}y$
$=\sup_{f\in\mathrm{Lip}(1)}\int K_{\sigma_{\min}}(y)(\nabla(f\ast p^{*}(-\cdot))(\xi_{y})-\nabla(f\ast p^{*}(-\cdot))(0))^{\top}y\;\mathrm{d}y$
$\quad+\int K_{\sigma_{\min}}(y)(\nabla(f\ast p^{*}(-\cdot))(0))^{\top}y\;\mathrm{d}y$
$\leq\sup_{f\in\mathrm{Lip}(1)}\int|K_{\sigma_{\min}}(y)||\nabla(f\ast p^{*}(-\cdot))(\xi_{y})-\nabla(f\ast p^{*}(-\cdot))(0)||y|\;\mathrm{d}y$
$\quad+\int K_{\sigma_{\min}}(y)(f\ast p^{*}(-\cdot))(0))^{\top}y\;\mathrm{d}y$
$\leq C_{1}^{\prime}\int|K_{\sigma_{\min}}(y)||y|^{1+\alpha}\;\mathrm{d}y+\int K_{\sigma_{\min}}(y)(\nabla(f\ast p^{*}(-\cdot))(0))^{\top}y\;\mathrm{d}y.$
For the first term we have that
$\int K_{\sigma_{\min}}(y)|y|^{1+\alpha}\;\mathrm{d}y$ $=\frac{1}{\sigma_{\min}^{d}}\int K(\frac{y}{\sigma_{\min}})|y|^{1+\alpha}\;\mathrm{d}y$
$=\sigma_{\min}^{1+\alpha}\int K(u)|u|^{1+\alpha}\;\mathrm{d}u.$
For the last term we use the assumption $\int K_{\sigma_{\min}}(y)y\;\mathrm{d}y=0.$
Hence we get
$\sup_{f\in\mathrm{Lip}(1)}\int f(z)(p^{*}(z)-(K_{\sigma_{\min}}\ast p^{*})(z))\;\mathrm{d}z\leq\sigma_{\min}^{1+\alpha}C_{1}^{\prime}\int K(u)|u|^{1+\alpha}\;\mathrm{d}u.$
By assumption the last term is finite. Setting $C_{1}\coloneqq C_{1}^{\prime}\int K(u)|u|^{1+\alpha}\;\mathrm{d}u$ yields the first term in the final bound.
Second term: We rewrite the second term as
$\sup_{f\in\mathrm{Lip}(1)}\int f(z)((K_{\sigma_{\min}}\ast p^{*})(z)$ $-(K_{\sigma_{\min}}\ast\mu_{n})(z))\;\mathrm{d}z$
$=\sup_{f\in\mathrm{Lip}(1)}\int\int f(z)K_{\sigma_{\min}}(z-x)(p^{*}(x)-\mu_{n}(x))\;\mathrm{d}x\;\mathrm{d}z$
$=\sup_{f\in\mathrm{Lip}(1)}\int(K_{\sigma_{\min}}(-\cdot)\ast f)(x)p^{*}(x)\;\mathrm{d}x-\int(K_{\sigma_{\min}}(-\cdot)\ast f)(x)\mu_{n}(x)\;\mathrm{d}x$
$=\sup_{\begin{subarray}{c}g=K_{\sigma_{\min}}(-\cdot)\ast f\\ f\in\mathrm{Lip}(1)\end{subarray}}\int g(x)(p^{*}(x)-\mu_{n}(x))\;\mathrm{d}x.$
For any multi index $k\in\mathbb{N}^{d}_{0}$ with $|k|\geq 1$, choose an index $j$ with $k_{j}\geq 1$ and set $k=\ell+e_{j}$ where $|k|=|\ell|+1$. Then using Young’s inequality and the properties of convolution, we obtain
$\|D^{k}(K_{\sigma_{\min}}\ast f)\|_{\infty}\leq\|D^{e_{j}}f\|_{\infty}\|D^{\ell}K_{\sigma_{\min}(-\cdot)}\|_{1}\leq 1\cdot\|D^{\ell}K_{\sigma_{\min}}\|_{1}.$
By assumption $\|D^{\ell}K\|_{1}\leq C$ and hence $\|D^{\ell}K_{\sigma_{\min}}\|_{1}\leq C\sigma_{\min}^{-|k|+1}.$ Thus
$\|D^{k}(K_{\sigma_{\min}}\ast f)\|_{\infty}\lesssim\sigma_{\min}^{-|k|+1}.$ (21)
Now we can use Schreuder (2020, Theorem 4). For even $d,$ set the smoothness parameter to $\frac{d}{2}.$ Then there exists a constant $C^{\prime}_{2}>0$ such that
$\sup_{\begin{subarray}{c}g=K_{\sigma_{\min}}\ast f\\ f\in\mathrm{Lip}(1)\end{subarray}}\int g(x)(p^{*}(x)-\mu_{n}(x))\;\mathrm{d}x\leq\frac{C^{\prime}_{2}}{\sigma_{\min}^{\frac{d}{2}-1}}n^{-1/2}\log(n).$
For odd $d>1,$ we know that
$\|D^{\frac{d-1}{2}}(f\ast K_{\sigma_{\min}})\|_{\infty}\leq\|D^{1}f\|_{\infty}\|D^{\frac{d-3}{2}}K_{\sigma_{\min}}\|_{1}\leq C^{\prime}_{2}\sigma_{\min}^{-\frac{d-3}{2}},$
and
$\frac{D^{\frac{d-1}{2}}(f\ast K_{\sigma_{\min}})(x)-D^{\frac{d-1}{2}}(f\ast K_{\sigma_{\min}})(y)}{|x-y|^{1/2}}\leq 2C^{\prime}_{2}\sigma_{\min}^{-\frac{d-3}{2}}\leq 2C^{\prime}_{2}\sigma_{\min}^{-\frac{d}{2}+1}.$
Then we can use Schreuder (2020, Theorem 4) with $\alpha=\frac{d}{2}$ again. Therefore, there is a $C^{\prime\prime}_{2}>0$ such that
$\sup_{\begin{subarray}{c}g=K_{\sigma_{\min}}\ast f\\ f\in\mathrm{Lip}(1)\end{subarray}}\int g(x)(p^{*}(x)-\mu_{n}(x))\;\mathrm{d}x\leq\frac{C^{\prime\prime}_{2}}{\sigma_{\min}^{\frac{d}{2}-1}}n^{-1/2}\log(n).$
Combining both terms we conclude that there are constants $C_{1}$ and $C_{2}$ with
$\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})\leq C_{1}\sigma_{\min}^{1+\alpha}+\frac{C_{2}}{\sigma_{\min}^{\frac{d}{2}-1}}n^{-1/2}\log(n).\qed$
Proof of Theorem˜3.7
The proof of Theorem˜3.7 requires some additional results. The proof of the subsidiary lemmata can be found in Section˜5.3. In order to control the error $e^{\int_{0}^{1}\Gamma_{t}\;\mathrm{d}t}\|v^{n}-\hat{v}\|_{\infty}$ (where we omitted constants) we need to approximate the function $v^{n}$ and control the Lipschitz constant of the approximation $\hat{v}_{t}$ for fixed $t.$ However, we cannot expect to approximate a non-trivial function on $\mathbb{R}^{d}$ with respect to the supremum norm. Additionally, $v^{n}_{t}$ as constructed by Lipman et al. (2023) is only locally Lipschitz in $x$ for fixed $t$ and the bound of the Lipschitz constant grows for $\sigma_{\min}\rightarrow 0$ (see Lemma˜5.2). Hence we cannot expect to construct a network $\hat{v}$ that approximates $v^{n}$ and its Lipschitz constant on $\mathbb{R}^{d}$ with a small absolute error for every $x\in\mathbb{R}.$
In the given setting, a rapid decay of the kernel function results in less significant approximation errors outside of the support of $p^{*}.$ We first show that in case of rapidly decreasing kernels the error $\mathsf{W}_{1}(\mathbb{P}^{\psi(Z)},\mathbb{P}^{\hat{\psi}(Z)})$ can be decomposed into two approximation errors, one depending on the approximation of $v$ on a compact domain and one corresponding to the area outside with little probability mass.
Lemma 5.1.
Assume that $\operatorname{supp}(p^{*})\subset(-1,1)^{d}$ and $\sigma_{\min}\leq 1.$ Additionally, assume that $K(x)=\prod_{i=1}^{d}\varphi(x_{i}),$ where $\varphi$ is a one dimensional symmetric density that decays faster than $x\mapsto e^{-\lambda x}$ for $|x|>2$ and some $\lambda>0$. Then for every $a>1$ there is a network $\tilde{v}$, which is cutoff outside $(-a,a)^{d}$ such that
$\mathsf{W}_{1}(\mathbb{P}^{\psi_{1}^{n}(Z)},\mathbb{P}^{\hat{\psi}_{1}(Z)})\leq\sqrt{2e}e^{\int_{0}^{1}\Gamma_{t}\;\mathrm{d}t}\Big(\int_{0}^{1}\int_{(-a,a)^{d}}|v^{n}_{t}(\psi^{n}_{t}(x))-\tilde{v}_{t}(\psi^{n}_{t}(x))|^{2}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t+c\frac{a^{2d+2}e^{-da}}{\sigma_{\min}^{2}}\Big)^{1/2}.$
Next we want to choose $a$ in Lemma˜5.1 and a neural network of finite size (depending on $n$) such that the bound of $\mathsf{W}_{1}(\mathbb{P}^{\psi^{n}(Z)},\mathbb{P}^{\hat{\psi}(Z)})$ decays at a rate of $n^{-\frac{1+\alpha}{2\alpha+d}}.$ Hence we need to approximate
$v^{n}_{t}(x)=\sum_{i=1}^{n}v_{t}(x|X_{i}^{*})\frac{p_{t}(x|X_{i}^{*})}{\sum_{i=1}^{n}p_{t}(x|X_{i}^{*})}$
and simultaneously control the Lipschitz constant of the approximation $\hat{v}$ on $(-a,a)^{d}.$ This is indeed sufficient, since using the same reasoning as in the proof of Lemma˜5.1, it holds that if $x\in(-a,a)^{d}$, then $v_{t}^{n}(x)\in(-a,a)^{d}$ for every $t$. As $\hat{v}$ approximates $v^{n}$ in the supremum norm, the Lipschitz constant of $\hat{v}$ will be at least of the order of the Lipschitz constant of $v^{n}$. The same applies to the supremum norm. Hence we need bounds for both, the Lipschitz constant and the supremum norm of $v^{n}.$
Lemma 5.2.
Assume that $\operatorname{supp}(p^{*})\subset[-1,1]^{d}.$ For every $x\in(-a,a)$ we have that
$|v^{n}_{t}(x)|\leq\frac{\sqrt{d}(1+a)}{1-(1-\sigma_{\min})t}\leq\frac{\sqrt{d}(1+a)}{\sigma_{\min}}.$
Further for the Gaussian kernel and $\operatorname{supp}(p^{*})\subset[-1,1]^{d}$, for any $x\in\mathbb{R}^{d}$ and every $t\in[0,1]$
$\operatorname{Lip}(v^{n}_{t})\leq\frac{1}{\sigma_{t}}+\frac{2d}{\sigma_{t}^{3}}.$
From (Gühring et al., 2020, Corollary 4.2) we know that for any $f\in C^{2}$ with bounded second partial derivatives we can approximate $f$ and its Lipschitz constant simultaneously using a ReLU network as defined in (17). To be more precise, for every $\varepsilon\in(0,1/2)$ there is a ReLU network $f_{\mathsf{NN}}$ such that
$\left\|f_{\mathsf{NN}}-f\right\|_{\mathcal{H}^{1}\left((0,1)^{d}\right)}\leq\varepsilon.$ (22)
The number of hidden layers $L$ of $f_{\mathsf{NN}}$ is bounded by $c\cdot\log\left(\varepsilon^{-2/(2-1)}\right)$ and the number of nonzero weights $S$ is bounded by $c\cdot\varepsilon^{-d/(2-1)}\cdot\log^{2}\left(\varepsilon^{-2/(2-1)}\right)$, where $c$ is a constant independent from $n$. Note that Gühring et al. (2020) refer to ReLU networks as defined in (17) as „standard neural networks“, since they also consider networks with skip connections.
Looking at the proof of Lemma˜5.2 and using standard analysis, we can bound the second partial derivatives of $v^{n}_{t}$ by $\frac{C}{\sigma_{t}^{5}}$ for some $C>0$. Now we are ready to proof Theorem˜3.7.
Proof of Theorem˜3.7.
From Lemma˜5.1, we know that
$\mathsf{W}_{1}(\mathbb{P}^{\psi_{1}^{n}(Z)},\mathbb{P}^{\hat{\psi}_{1}(Z)})\leq\sqrt{2e}e^{\int_{0}^{1}\tilde{L}_{t}\;\mathrm{d}t}\inf_{\tilde{v}\in\mathcal{N}}\Big(\int_{0}^{1}\int_{(-a,a)^{d}}|v^{n}_{t}(\psi^{n}_{t}(x))-\tilde{v}_{t}(\psi^{n}_{t}(x))|^{2}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t+c\frac{a^{2d+2}e^{-da}}{\sigma_{\min}^{2}}\Big)^{1/2},$
where we used the choice of (11) before dividing the area in (35) in the proof of Lemma˜5.1 and $\tilde{L}$ is the Lipschitz constant of $\tilde{v}$ at time $t$. We neglect an arbitrary small additional error on the right hand side to account for the use of the infimum, as the error bound corresponding to the kernel density estimator in (16) is nonzero anyway. Note that due to the restriction of the network class,
$e^{\int_{0}^{1}\tilde{L}_{t}\;\mathrm{d}t}\leq e\cdot e^{\int_{0}^{1}\frac{2d+1}{\sigma_{t}^{3}}\;\mathrm{d}t}.$
Integration yields for $n$ big enough
$\int_{0}^{1}\frac{1}{\sigma_{t}^{3}}\;\mathrm{d}t=\int_{0}^{1}\frac{1}{(1-(1-\sigma_{\min})t)^{3}}\;\mathrm{d}t\leq\frac{1}{2(1-\sigma_{\min})\sigma_{\min}^{2}}-\frac{1}{2(1-\sigma_{\min})}\leq\frac{1}{\sigma_{\min}^{2}}.$
To bound the tail, we need to assure $a$ is chosen such that
$e^{\frac{2d+1}{\sigma_{\min}^{2}}}\frac{a^{d+1}e^{-\frac{da}{2}}}{\sigma_{\min}}\leq n^{-\frac{1+\alpha}{2\alpha+d}}.$
As in the setting $d\geq 1,$ we can use that for $a\geq 15$
$a^{d+1}e^{-da/2}\leq e^{-\frac{da}{8}}.$
For $\sigma_{\min}\sim n^{-\frac{1}{2\alpha+d}}$, the above inequality is satisfied if
$a\geq\max\big(\frac{8}{d}\big(\frac{2+\alpha}{2\alpha+d}\log(n)+(2d+1)n^{\frac{2}{2\alpha+d}}\big),15\big).$
To cover $(-a,a)^{d}$ with cubes of size $(0,1)^{d}$, we need $\lceil 2a\rceil^{d}$ little cubes.
Now we want to bound the approximation term. We can bound
$\Big(\int_{0}^{1}\int_{(-a,a)^{d}}|v^{n}_{t}(\psi^{n}_{t}(x))-\tilde{v}_{t}(\psi^{n}_{t}(x))|^{2}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t\Big)^{1/2}$ $\leq\|v^{n}_{t}-\tilde{v}_{t}\|_{\infty,(-a,a)^{d}}.$
Note that $v_{t}^{n}\in C^{\infty}$ in case of the Gaussian kernel. Hence we can apply (Gühring et al., 2020, Corollary 4.2). The bound of the second partial derivatives as well as the supremum norm bound of $v^{n}$ influence the coefficients in (Gühring et al., 2020, Corollary 4.2), therefore we construct the network such that $n^{-\frac{5}{2\alpha+d}}v^{n}$ is approximated and choose the last weight matrix such that it scales the result up. This addition of one layer does not influence the order of the layers and non zero weights of the network. However, the approximation error gets scaled up too, hence we need to ensure a smaller error. We need to choose $\varepsilon$ in (22) such that
$e^{(2d+1)n^{\frac{2}{2\alpha+d}}}\sqrt{\varepsilon}\leq n^{-\frac{6+\alpha}{2\alpha+d}}\quad\Longleftrightarrow\quad\varepsilon\leq n^{-\frac{12+2\alpha}{2\alpha+d}}e^{-(4d+2)n^{\frac{2}{2\alpha+d}}}.$
The addition of one cutting layer to ensure the assumptions of Lemma˜5.1 does not influence the order of the layers and non zero weights of the network either. Hence choosing
$L_{n}$ $\gtrsim\Big(\max\Big(\log(n)+n^{\frac{2}{2\alpha+d}},15\Big)\Big)^{d}\cdot\big(\log^{2}(n)+n^{\frac{4}{2\alpha+d}}\big),$
$S_{n}$ $\gtrsim\Big(\max\Big(\log(n)+n^{\frac{2}{2\alpha+d}},15\Big)\Big)^{d}\cdot\big(n^{\frac{12+2\alpha}{2\alpha+d}}e^{(4d+2)n^{\frac{2}{2\alpha+d}}}\big)^{d}\cdot\big(\log^{2}(n)+n^{\frac{4}{2\alpha+d}}\big),$
yields the desired rate. ∎
Proof of Theorem˜3.10.
For this proof, we need some auxiliary lemmata, whose proofs are postponed to Section˜5.3. First, we need to bound the Lipschitz constant of $\pi$ on $U_{r}$.
Lemma 5.3.
The mapping $\pi\Big|_{B(r,\mathcal{M})}$ is Lipschitz continuous with Lipschitz constant $\frac{1}{1-\frac{r}{\tau}}$.
Then we decompose as in Theorem˜3.4
$\mathsf{W}_{1}(\mathbb{P}^{*},\pi_{\#}\mathbb{P}^{\psi^{n}_{1}}(Z))\leq\mathsf{W}_{1}(\mathbb{P}^{*},\pi_{\#}\mathbb{P}^{\psi_{1}}(Z))+\mathsf{W}_{1}(\pi_{\#}\mathbb{P}^{\psi_{1}}(Z),\pi_{\#}\mathbb{P}^{\psi^{n}_{1}}(Z)).$ (23)
First term: We first note that
$\mathsf{W}_{1}(\mathbb{P}^{*},\pi_{\#}\mathbb{P}^{\psi_{1}}(Z))$ $=\sup_{\begin{subarray}{c}f\in\operatorname{Lip}(1)\\ f(0)=0\end{subarray}}\mathbb{E}_{\begin{subarray}{c}X\sim\mathbb{P}^{*}\\ U\sim\mathcal{N}(0,I_{d})\end{subarray}}\big[f(X)-f(\pi(X+\sigma_{\min}U))\big]$
$=\sup_{\begin{subarray}{c}f\in\operatorname{Lip}(1)\\ f(0)=0\end{subarray}}\mathbb{E}_{\begin{subarray}{c}Y\sim\mathbb{P}^{*}_{d^{\prime}}\\ U\sim\mathcal{N}(0,I_{d})\end{subarray}}\big[f(g(Y))-f(\pi(g(Y)+\sigma_{\min}U))\big]$
$=\sup_{\begin{subarray}{c}f\circ g=\varphi\\ f\in\operatorname{Lip}(1)\\ f(0)=0\end{subarray}}\mathbb{E}\big[\varphi(Y)-\varphi(\rho_{\sigma_{\min}}(Y,U))\big],$ (24)
where we define the $\mathbb{R}^{d^{\prime}}$ counterpart of the projection as
$\rho_{\sigma_{\min}}(y,u)\coloneqq g^{-1}(\pi(g(y)+\sigma_{\min}u)).$
In the following, we derive a second order Taylor expansion of $\rho$, that linearizes along the manifold. This step of the proof is sketched in Figure˜3.
Figure 3: Setting for the .
To this end, we start by bounding the projection of a disturbed point on the manifold by the magnitude of the distribution and a vector in the tangent space of the point. Let $x\in\mathcal{M}$. Denote the tangent space of the point $x$ as $T_{x}\mathcal{M}$ and the normal space as $N_{x}\mathcal{M}$. Define the $r$ surrounding on $\mathcal{M}$ as
$\mathcal{M}_{x,r}\coloneqq\{x+\omega+h_{x}(\omega)\colon\omega\in T_{x}\mathcal{M},|\omega|\leq r\},$
where $h_{x}\colon T_{x}\mathcal{M}\rightarrow N_{x}\mathcal{M}$ is the function that maps a vector in the tangent space to the vector in the normal space such that $x+\omega+h_{x}(\omega)\in\mathcal{M}$. Note that this is a slightly larger set than $\mathcal{M}\cap B(r,\{x\})$. Since $x\in\mathcal{M}$, we have that $h(0)=0$. Since $T_{x}\mathcal{M}$ is a tangent space, $Dh(0)=0$. Additionally $\|D^{2}h(\omega)\|\lesssim\frac{1}{\tau}$, see Aamari et al. (2019, Proposition 2.3) referring to Niyogi et al. (2008, Proposition 6.1). A second order Taylor expansion, like in Tang and Yang (2024); Divol (2022) yields
$|h(\omega)|\lesssim\frac{1}{\tau}|\omega|^{2}.$ (25)
For $z\in B(r,\{x\})$, let $z=x+v$, with $|v|\leq r<1$. There is a unique decomposition $v=v_{T}+v_{N}$ with $v_{T}\in T_{x}\mathcal{M}$ and $v_{N}\in N_{x}\mathcal{M}$. Then $\pi(z)=x+\omega^{*}+h(\omega^{*})$, where $\omega^{*}$ is the minimizer of
$|z-(x+\omega+h(\omega))|^{2}=|v-\omega-h(\omega)|^{2}=|v_{T}-\omega|^{2}+|v_{N}-h(\omega)|^{2}$ (26)
and $\pi(z)\in\mathcal{M}_{x,r}$. Since the objective (26) is convex and since $\mathcal{M}_{x,r}$ is compact, the unique minimizer satisfies
$\omega^{*}-v_{T}=-Dh(\omega^{*})^{\top}(h(\omega^{*})-v_{N}).$
Taking norms yields
$|\omega^{*}-v_{T}|\leq|Dh(\omega^{*})||h(\omega^{*})-v_{N}|\leq\|Dh(\omega^{*})\|(|h(\omega^{*})|+|v_{N}|).$ (27)
Due to the Pythagorean theorem, we can bound $|v_{N}|\leq|v|$ and $|v_{T}|\leq|v|$. Since $\omega^{*}$ is the minimizer of (26), we know that
$|\omega^{*}-v_{T}|^{2}\leq|\omega^{*}-v_{T}|^{2}+|v_{N}-h(\omega^{*})|^{2}\leq|v_{N}-h(v_{T})|^{2}\leq 2(|v|^{2}+|h(v_{T})|^{2}).$
Using (25), we obtain using $|v|<1$
$|\omega^{*}|\leq|v_{T}|+|\omega^{*}-v_{T}|\lesssim|v|+\sqrt{|v|^{2}+\frac{1}{\tau}|v|^{4}}\lesssim\frac{1}{\min(\sqrt{\tau},1)}|v|.$
Inserting this bound into (27) and using that by Taylor expansion $\|Dh(\omega^{*})\|\lesssim\frac{1}{\tau}|\omega^{*}|$ leads to
$|\omega^{*}-v_{T}|\lesssim\frac{1}{\min(\tau,\tau^{3/2})}|v|^{2}.$
Thus
$\pi(x+v)=x+\omega^{*}+h(\omega^{*})$
$\Longleftrightarrow\quad$ $\pi(x+v)-(x+v_{T})=\omega^{*}-v_{T}+h(\omega^{*})$
$\Longrightarrow\quad$ $|\pi(x+v)-(x+v_{T})|\leq|\omega^{*}-v_{T}|+|h(\omega^{*})|$
$\Longrightarrow\quad$ $|\pi(x+v)-(x+v_{T})|\lesssim|v|^{2}.$
Therefore
$\pi(x+v)=x+v_{T}+R(x,v),\quad\text{where}\quad|R(x,v)|\lesssim|v|^{2}.$
Applying this to the projection of $g(y)+\sigma_{\min}u$ leads to
$\pi(g(y)+\sigma_{\min}u)=g(y)+\sigma_{\min}u_{T}+R(g(y),\sigma_{\min}u),$
where $\|R(g(y),\sigma_{\min}u)\|\lesssim\sigma_{\min}^{2}|u|^{2}$. Thus
$\pi(g(y)+\sigma_{\min}u)-g(y)=\sigma_{\min}u_{T}+R(g(y),\sigma_{\min}u)=:\eta.$
A Taylor expansion of $g^{-1}$ around $g(y)$ yields
$g^{-1}(\pi(g(y)+\sigma_{\min}u))=g^{-1}(g(y)+\eta)=y+Dg^{-1}(g(y))\eta+O(|\eta|^{2}),$
where $O(\|\eta\|^{2})$ represents a vector whose norm is of order $\|\eta\|^{2}$. Inserting the definition of $\eta$ leads to
$\rho_{\sigma_{\min}}(y,u)=g^{-1}(\pi(g(y)+\sigma_{\min}u))$ $=y+Dg^{-1}(g(y))(\sigma_{\min}u_{T}+R(g(y),\sigma_{\min}u)))$
$\quad+O(|\sigma_{\min}u_{T}+R(g(y),\sigma_{\min}u)|^{2})$
$=y+\sigma_{\min}A(y)u+R^{\prime}(g(y),\sigma_{\min}u),$
where $|R^{\prime}(g(y),\sigma_{\min}u)|\lesssim\sigma_{\min}^{2}|u|^{2}$ and $A(y)\coloneqq Dg^{-1}(g(y))\circ P_{T_{g(y)}\mathcal{M}}$ with $P_{T_{g(y)}\mathcal{M}}$ being the orthogonal projection onto the tangent space of $\mathcal{M}$ at the point $g(y)$.
This leads to, using that $\pi$ sets everything to 0 outside for $|u|>r$ and setting $\rho\coloneqq\rho_{1}$
$\int\int\varphi(\rho_{\sigma_{\min}}(y,u))p^{*}_{d^{\prime}}(y)K(u)\;\mathrm{d}y\;\mathrm{d}u$ $=\int_{B(0,r)}\int\varphi(\rho_{1}(y,u))p^{*}_{d^{\prime}}(y)K_{\sigma_{\min}}(u)\;\mathrm{d}y\;\mathrm{d}u$
$\quad+\int_{B(0,r)^{c}}\int\varphi(\rho_{1}(y,u))p^{*}_{d^{\prime}}(y)K_{\sigma_{\min}}(u)\;\mathrm{d}y\;\mathrm{d}u.$
For the first term, we obtain
$\int_{B(0,r)}\int\varphi(\rho_{1}(y,u))p^{*}_{d^{\prime}}(y)K_{\sigma_{\min}}(u)\;\mathrm{d}y\;\mathrm{d}u$ $\lesssim\int_{B(0,r)}\int\varphi(y+A(y)u)p^{*}_{d^{\prime}}(y)K_{\sigma_{\min}}(u)\;\mathrm{d}y\;\mathrm{d}u$
$\quad+\int_{B(0,r/\sigma_{\min})}\int|R^{\prime}(g(y),\sigma_{\min}u)|p^{*}_{d^{\prime}}(y)K(u)\;\mathrm{d}y\;\mathrm{d}u$
$\lesssim\int_{B(0,r)}\int\varphi(y+A(y)u)p^{*}_{d^{\prime}}(y)K_{\sigma_{\min}}(u)\;\mathrm{d}y\;\mathrm{d}u+\sigma_{\min}^{2}.$ (28)
For the integral over $B(0,r)^{c}$, we change the order of integration. For $y$ fixed, we define two sets
$A_{0}(y)\coloneqq\{u\in B(0,r)^{c}\mid g(y)+u\notin B(r,\mathcal{M})\},\quad A_{\text{far}}\coloneqq\{u\in B(0,r)^{c}\mid g(y)+u\in B(r,\mathcal{M})\}.$
$A_{0}(y)$ consists of all $u$ such that $\pi(g(y)+u)=0$. $A_{\text{far}}$ consists of all $u$ such that $g(y)+u$ is close to the manifold, which implies that $|u|\geq 2\tau-r$ by the reach condition and the construction of the projection. Inserting this integral split and (28) into (24) leads to
$\sup_{\begin{subarray}{c}f\circ g=\varphi\\ f\in\operatorname{Lip}(1)\\ f(0)=0\end{subarray}}\mathbb{E}\big[\varphi(Y)-\varphi(\rho_{\sigma_{\min}}(Y,U))\big]$ $\lesssim\sup_{\begin{subarray}{c}f\circ g=\varphi\\ f\in\operatorname{Lip}(1)\\ f(0=0)\end{subarray}}\int_{B(0,r)}\int(\varphi(y)-\varphi(y+A(y)u))p^{*}_{d^{\prime}}(y)K_{\sigma_{\min}}(u)\;\mathrm{d}y\;\mathrm{d}u$ (29)
$\quad+\sigma_{\min}^{2}+\int\int_{A_{0}(y)}(\varphi(y)-\varphi(g^{-1}(0)))p^{*}_{d^{\prime}}(y)K_{\sigma_{\min}}(u)\;\mathrm{d}u\;\mathrm{d}y$ (30)
$\quad+\int\int_{A_{\text{far}}}(\varphi(y)-\varphi(g^{-1}(\pi(g(y)+u))))p^{*}_{d^{\prime}}(y)K_{\sigma_{\min}}(u)\;\mathrm{d}u\;\mathrm{d}y.$ (31)
For $\sigma_{\min}$ small enough, which will translate into $n$ large enough, the (30) is bounded by $\sigma_{\min}^{2}$, since the tails of the Gaussian decay exponential.
For (31), we first note that under ˜3.9 the Jacobian of the inverse $g^{-1}$ of $g$ is bounded. We are going to calculate this bound as a byproduct of the proof of Lemma˜5.4. Therefore there is a $\xi_{y,u}$ with $|\xi_{y,u}|\leq r$ such that
$|\varphi(g^{-1}(g(y)))-\varphi(g^{-1}(\pi(g(y)+u)))|\lesssim|g(y)-\pi(g(y)+u)|=|g(y)-g(y)-u-\xi_{y,u}|\leq|u|+r.$
Then for $\sigma_{\min}$ small enough
$\int\int_{A_{\text{far}}}(\varphi(y)-\varphi(g^{-1}(\pi(g(y)+u))))p^{*}_{d^{\prime}}(y)K_{\sigma_{\min}}(u)\;\mathrm{d}u\;\mathrm{d}y\leq\int\int_{B(0,r)^{c}}(|u|+r)p^{*}_{d^{\prime}}(y)K_{\sigma_{\min}}(u)\;\mathrm{d}u\;\mathrm{d}y\lesssim\sigma_{\min}^{2}.$
To bound (29), we first define
$\ell(u)\coloneqq\int\varphi(y+A(y)u)p^{*}_{d^{\prime}}(y)\;\mathrm{d}y.$
Then, we can rewrite
$\int_{B(0,r)}\int(\varphi(y)-\varphi(y+A(y)u))p^{*}_{d^{\prime}}(y)K_{\sigma_{\min}}(u)\;\mathrm{d}y\;\mathrm{d}u=\int_{B(0,r)}K_{\sigma_{\min}}(u)(\ell(0)-\ell(u))\;\mathrm{d}u.$ (32)
Analogous to Theorem˜3.4, we can use the multivariate mean value theorem to conclude that for every $u$ there is a $\xi_{u}$ with $|\xi_{u}|\leq|u|$ such that
$\int_{B(0,r)}K_{\sigma_{\min}}(u)(\ell(0)-\ell(u))\;\mathrm{d}u=\int_{B(0,r)}K_{\sigma_{\min}}(u)\nabla\ell(\xi_{u})^{\top}u\;\mathrm{d}u=\int_{B(0,r)}K_{\sigma_{\min}}(u)(\nabla\ell(\xi_{u})-\nabla\ell(0))^{\top}u\;\mathrm{d}u,$
where the last inequality follows from the symmetry of $K_{\sigma_{\min}}$. Given that the mapping $F_{u}(y)\coloneqq y+A(y)u$ is a bijection for all $u\in B(0,r)$, and denote the inverse by $S_{u}(z)$. If that is the case, then
$\nabla\ell(u)=\int A(y)^{\top}(\nabla\varphi)(y+A(y)u)p^{*}_{d^{\prime}}(y)\;\mathrm{d}y=\int A(S_{u}(z))^{\top}(\nabla\varphi)(z)|\operatorname{det}(DS_{u}(z))|p^{*}_{d^{\prime}}(S_{u}(z))\;\mathrm{d}z.$ (33)
The following lemma shows that this is justified as well as several properties of the mapping $F_{u}$ and its inverse, that will be used in the following. Intuitively, the properties follow from the fact that $F_{u}$ is a small perturbation of the identity.
Lemma 5.4.
For any $u\in\mathbb{R}^{d}$ such that $|u|\leq r$, the mapping $F_{u}\colon\mathbb{R}^{d^{\prime}}\rightarrow\mathbb{R}^{d^{\prime}},\ F_{u}(y)\coloneqq y+A(y)u$ is a diffeomorphism. Further
$\mathrm{(I)}\quad$ $\sup_{y\in\mathbb{R}^{d^{\prime}}}\|A(y)\|<1,$
$\mathrm{(II)}\quad$ $\sup_{y\in\mathbb{R}^{d^{\prime}}}\|DA(y)\|\leq c_{0}<\frac{1}{r},$
$\mathrm{(III)}\quad$ $F_{u}\text{ is injective},$
$\mathrm{(IV)}\quad$ $S_{u}(z)=z-A(S_{u}(z))u,$
$\mathrm{(V)}\quad$ $\sup_{y\in\mathbb{R}^{d^{\prime}}}\|DS_{u}(y)\|\lesssim\frac{1}{1-c_{0}r},$
$\mathrm{(VI)}\quad$ $\sup_{y\in\mathbb{R}^{d^{\prime}}}||\operatorname{det}(DS_{u}(y))|-1|=O(|u|).$
Using Lemma˜5.4 and the fact that $\varphi$ is Lipschitz, we obtain
$|\nabla\ell(\xi_{u})-\nabla\ell(0)|$ $=\Big|\int(\nabla\varphi(z))^{\top}\Big(A(S_{\xi_{u}}(z))p^{*}_{d^{\prime}}(S_{\xi_{u}}(z))|\operatorname{det}(DS_{\xi_{u}}(z))|-A(z)p^{*}_{d^{\prime}}(z)\Big)\;\mathrm{d}z\Big|$
$\leq\int|\nabla\varphi(z)|\|A(S_{\xi_{u}}(z))-A(z)\|p^{*}_{d^{\prime}}(S_{\xi_{u}}(z))|\operatorname{det}(DS_{\xi_{u}}(z))|\;\mathrm{d}z$
$\quad+\int|\nabla\varphi(z)|\|A(z)\|p^{*}_{d^{\prime}}(S_{\xi_{u}}(z))|\operatorname{det}(DS_{\xi_{u}})-1|\;\mathrm{d}z$
$\quad+\int|\nabla\varphi(z)|\|A(z)\||p^{*}_{d^{\prime}}(S_{\xi_{u}}(z))-p^{*}_{d^{\prime}}(z)|\;\mathrm{d}z$
$\lesssim|u|+|u|^{\alpha}.$
Plugging everything into (32), we obtain
$\int_{B(0,r)}K_{\sigma_{\min}}(u)(\ell(0)-\ell(u))\;\mathrm{d}u\leq\int_{B(0,r)}K_{\sigma_{\min}}(u)(|u|+|u|^{\alpha})|u|\;\mathrm{d}u\lesssim\sigma_{\min}^{2}+\sigma_{\min}^{1+\alpha}.$
Since $\sigma_{\min}\leq 1$, we have shown
$\mathsf{W}_{1}(\mathbb{P}^{*},\pi_{\#}\mathbb{P}^{\psi_{1}}(Z))\lesssim\sigma_{\min}^{1+\alpha}.$
Second term: First, we define the set
$\mathcal{F}\coloneqq\Big\{h\colon\mathbb{R}^{d}\rightarrow\mathbb{R}\mid h=f\circ\varphi,f\in\operatorname{Lip}(1),f(0)=0,\varphi\in\operatorname{Lip}(\frac{\tau}{\tau-r}),\exists x\in\mathcal{M}\text{ s.t. }\varphi(x)=x\Big\}.$
For the second term in (23), we can use the Lipschitz bound from Lemma˜5.3 to bound
$\mathsf{W}_{1}(\pi_{\#}\mathbb{P}^{\psi_{1}}(Z),\pi_{\#}\mathbb{P}^{\psi^{n}_{1}}(Z))$ $=\sup_{\begin{subarray}{c}f\in\operatorname{Lip}(1)\\ f(0)=0\end{subarray}}\int f(\pi(x))\;\mathrm{d}(K_{\sigma_{\min}}\ast\mathbb{P}^{*}-K_{\sigma_{\min}}\ast\mathbb{P}_{n})(x)$
$=\sup_{\begin{subarray}{c}f\in\operatorname{Lip}(1)\\ f(0)=0\end{subarray}}\int_{B(\mathcal{M},r)}f(\pi(x))\;\mathrm{d}(K_{\sigma_{\min}}\ast\mathbb{P}^{*}-K_{\sigma_{\min}}\ast\mathbb{P}_{n})(x)$
$\leq\sup_{h\in\mathcal{F}}\int_{B(\mathcal{M},r)}h(x)\;\mathrm{d}(K_{\sigma_{\min}}\ast\mathbb{P}^{*}-K_{\sigma_{\min}}\ast\mathbb{P}_{n})(x)$
$\leq\sup_{h\in\mathcal{F}}\int h(x)\;\mathrm{d}(K_{\sigma_{\min}}\ast\mathbb{P}^{*}-K_{\sigma_{\min}}\ast\mathbb{P}_{n})(x)$ (34)
$\quad+\sup_{h\in\mathcal{F}}\int_{B(\mathcal{M},r)^{c}}h(x)\;\mathrm{d}(K_{\sigma_{\min}}\ast\mathbb{P}^{*}-K_{\sigma_{\min}}\ast\mathbb{P}_{n})(x).$
Since we assumed that $g$ is bounded and $\mathcal{F}$ consists of concatenations of functions that grow at most linearly, we conclude due to the exponential decay of the Gaussian distribution that for $\sigma_{\min}$ small enough with respect to $r$,
$\sup_{h\in\mathcal{F}}\int_{B(\mathcal{M},r)^{c}}h(x)\;\mathrm{d}(K_{\sigma_{\min}}\ast\mathbb{P}^{*}-K_{\sigma_{\min}}\ast\mathbb{P}_{n})(x)\lesssim\sigma_{\min}^{2}.$
For (34), we obtain
$\sup_{h\in\mathcal{F}}\int h(x)\;\mathrm{d}(K_{\sigma_{\min}}\ast\mathbb{P}^{*}-K_{\sigma_{\min}}\ast\mathbb{P}_{n})(x)$ $\lesssim\mathsf{W}_{1}(\mathbb{P}^{\psi_{1}}(Z),\mathbb{P}^{\psi^{n}_{1}}(Z))$
$=\sup_{f\in\operatorname{Lip}(1)}\int f(x)\;\mathrm{d}((K_{\sigma_{\min}}\ast\mathbb{P}^{*})-(K_{\sigma_{\min}}\ast\mathbb{P}_{n}))(x)$
$=\sup_{f\in\operatorname{Lip}(1)}\int(K_{\sigma_{\min}}(-\cdot)\ast f)(x)\;\mathrm{d}(\mathbb{P}^{*}-\mathbb{P}_{n})(x)$
$=\sup_{f\in\operatorname{Lip}(1)}\int(K_{\sigma_{\min}}(-\cdot)\ast f)(g(y))\;\mathrm{d}(\mathbb{P}^{*}_{d^{\prime}}-\mathbb{P}_{n,d^{\prime}})(y)$
$=\sup_{\begin{subarray}{c}h=(K_{\sigma_{\min}}\ast f)(g)\\ f\in\operatorname{Lip}(1)\end{subarray}}\int h(y)(p^{*}_{d^{\prime}}(y)-\mu_{n,d^{\prime}}(y))\;\mathrm{d}y,$
where $\mathbb{P}_{n,d^{\prime}}$ is the empirical measure resulting from $Y_{1},...,Y_{n}$ such that $g(Y_{i})=X_{i}$ for every $i\in\{1,...,n\}$ and $\mu_{n,d^{\prime}}=\frac{1}{n}\sum_{i=1}^{n}\delta_{Y_{i}}$. From (21), we know that for any multi-index $k\in\mathbb{N}_{0}^{d}$ with $|k|\geq 1$,
$\|D^{k}(K_{\sigma_{\min}}\ast f)\|_{\infty}\lesssim\sigma_{\min}^{-|k|+1}.$
Next, we use Faà di Bruno’s formula and the boundedness of all derivatives of $g$ to conclude
$\|D^{k}(K_{\sigma_{\min}}\ast f)(g)\|_{\infty}\lesssim\sigma_{\min}^{-|k|+1},$
where the constant depends on $k,d$ and the properties of $g$. Now we can proceed as in Theorem˜3.4 and obtain
$\mathsf{W}_{1}(\pi_{\#}\mathbb{P}^{\psi_{1}}(Z),\pi_{\#}\mathbb{P}^{\psi^{n}_{1}}(Z))\lesssim\frac{1}{\sqrt{\sigma_{\min}^{d^{\prime}-2}}}n^{-\frac{1}{2}}.$
Proof of Theorem˜3.11.
We proceed completely analogous to Theorem˜3.7 replacing $d$ with $d^{\prime}$. However, since the vector field we want to approximate is still $d$ dimensional, we still need $\lceil 2a\rceil^{d}$ little cubes. Hence we choose
$L_{n}$ $\gtrsim\Big(\max\Big(\log(n)+n^{\frac{2}{2\alpha+d^{\prime}}},15\Big)\Big)^{d}\cdot\big(\log^{2}(n)+n^{\frac{4}{2\alpha+d^{\prime}}}\big),$
$S_{n}$ $\gtrsim\Big(\max\Big(\log(n)+n^{\frac{2}{2\alpha+d^{\prime}}},15\Big)\Big)^{d}\cdot\big(n^{\frac{12+2\alpha}{2\alpha+d^{\prime}}}e^{(4d^{\prime}+2)n^{\frac{2}{2\alpha+d^{\prime}}}}\big)^{d}\cdot\big(\log^{2}(n)+n^{\frac{4}{2\alpha+d^{\prime}}}\big),$
to obtain the desired rate. ∎
5.3 Auxiliary results
Proof of Lemma˜5.1.
First, recall the proof from Theorem˜3.2. For any $a>0$ we can bound the term from (20)
$Q_{1}$ $\leq\sqrt{e}e^{2\int_{0}^{1}\Gamma_{t}\;\mathrm{d}t}\int_{0}^{1}\int|v^{n}_{t}(\psi^{n}_{t}(x))-\tilde{v}_{t}(\psi^{n}_{t}(x))|^{2}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t$
$=\sqrt{e}e^{2\int_{0}^{1}\Gamma_{t}\;\mathrm{d}t}\Big(\int_{0}^{1}\int_{(-a,a)^{d}}|v^{n}_{t}(\psi^{n}_{t}(x))-\tilde{v}_{t}(\psi^{n}_{t}(x))|^{2}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t$ (35)
$\quad+\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus(-a,a)^{d}}|v^{n}_{t}(\psi^{n}_{t}(x))-\tilde{v}_{t}(\psi^{n}_{t}(x))|^{2}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t\Big).$
The second term can be bounded by
$\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus(-a,a)^{d}}|v^{n}_{t}(\psi^{n}_{t}(x))$ $-\tilde{v}_{t}(\psi^{n}_{t}(x))|^{2}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t$
$\leq 2\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus(-a,a)^{d}}\Big(|v^{n}_{t}(\psi^{n}_{t}(x))|^{2}+|\tilde{v}_{t}(\psi^{n}_{t}(x))|^{2}\Big)p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t.$ (36)
Now
$|v^{n}_{t}(\psi^{n}_{t}(x))|^{2}\leq\max_{i\in\{1,\dots,n\}}|v_{t}(\psi^{n}_{t}(x)|X_{i}^{*})|^{2}\leq\frac{|X_{i}^{*}-(1-\sigma_{\min})\psi^{n}_{t}(x)|^{2}}{(1-(1-\sigma_{\min})t)^{2}}\leq 2\frac{|X_{i}^{*}|^{2}+(1-\sigma_{\min})^{2}|\psi^{n}_{t}(x)|^{2}}{(1-(1-\sigma_{\min})t)^{2}}.$
As $\mathbb{P}^{*}$ has compact support within $(-a,a)^{d}$,
$|X_{i}^{*}|^{2}\leq d\cdot a^{2}.$
To bound $\psi_{t}^{n}(x)$ for $x\in\mathbb{R}^{d}\setminus(-a,a)^{d}$, we use that
$|\psi^{n}_{t}(x)|\leq|x|.$ (37)
To see why this is true, consider two cases. First, let $t$ be small such that $\psi^{n}_{t}(x)\notin(-a,a)^{d}.$ Then for every $X_{i}^{*},i\in\{1,\dots,n\}$ we have by construction that $v^{n}_{t}(\psi_{t}^{n}(x)|X_{i}^{*})$ is in the smallest convex cone that includes $\psi_{t}^{n}(x)$ and $(-1,1)^{d}$, where $\psi_{t}^{n}(x)$ is the vertex of the convex cone. Any linear combination of $v^{n}_{t}(\psi_{t}^{n}(x)|X_{i}^{*}),i=1,\dots,n,$ must be in this convex cone as well. Therefore $v_{t}^{n}(x)$ must be in the convex cone. Since
$\frac{d}{dt}\psi^{n}_{t}(x)=v^{n}_{t}(\psi_{t}(x)),\quad\psi^{n}_{0}(x)=x,$
this means that for small $t$, $\psi_{t}$ pushes the initial condition $x$ in the direction of $(-a,a)^{d}$ and therefore (37) is true.
Second, let $t^{\prime}$ be the smallest $t$ such that $\psi^{n}_{t^{\prime}}(x)\in(-a,a)^{d}.$ Then for all $t>t^{\prime}$ we have that $\psi^{n}_{t}(x)\in(-a,a)^{d}.$ This holds since within $(-a,a)^{d}\setminus(-1,1)^{d}$ using the same argument as in the first case, $v_{t}^{n}(x)$ is a linear combination of $v^{n}_{t}(\psi_{t}^{n}(x)|X_{i}^{*})$ and thus $\psi_{t}$ pushes $x$ in the direction of $(-1,1)^{d}.$ If $t$ is such that $\psi_{t}(x)\in(-1,1)^{d},$ there cannot be a $t^{\prime}>t$ such that $\psi_{t^{\prime}}(x)\notin(-a,a)^{d},$ since by continuity of $\psi_{t},$ there must be a $t^{\prime\prime}$ with $t<t^{\prime\prime}<t^{\prime}$ such that $\psi_{t^{\prime\prime}}(x)\in(-a,a)^{d}.$ In this case the corresponding vector field is oriented towards $(-1,1)^{d},$ which leads to $\psi_{t}$ pushing $x$ back to $(-1,1)^{d}.$ Hence, in any case for $x\notin(-a,a)^{d}$, we have that (37) is true.
Therefore we get for (36) for $\sigma_{\min}\leq 1$
$\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus(-a,a)^{d}}$ $|v^{n}_{t}(\psi^{n}_{t}(x))-\tilde{v}_{t}(\psi^{n}_{t}(x))|^{2}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t\leq 8\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus(-a,a)^{d}}\frac{da^{2}+(1-\sigma_{\min})^{2}|x|^{2}}{(1-(1-\sigma_{\min})t)^{2}}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t$
$\leq\frac{8da^{2}}{\sigma_{\min}^{2}}\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus(-a,a)^{d}}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t+\frac{(1-\sigma_{\min})^{2}}{\sigma_{\min}^{2}}\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus(-a,a)^{d}}|x|^{2}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t.$
For the first term we get
$\frac{8da^{2}}{\sigma_{\min}^{2}}\int_{\mathbb{R}^{d}\setminus(-a,a)^{d}}p_{0}(x)\;\mathrm{d}x$ $=\frac{8da^{2}}{\sigma_{\min}^{2}}\Big(\int_{\mathbb{R}\setminus(-a,a)}\varphi(x)\;\mathrm{d}x\Big)^{d},$
where $\varphi$ is the pdf of the one dimensional kernel distribution. Since $a>1$, we get for the second term
$\frac{(1-\sigma_{\min})^{2}}{\sigma_{\min}^{2}}\int_{\mathbb{R}^{d}\setminus(-a,a)^{d}}|x|^{2}p_{0}(x)\;\mathrm{d}x$ $=\frac{(1-\sigma_{\min})^{2}}{\sigma_{\min}^{2}}\sum_{i=1}^{d}\int_{\mathbb{R}^{d}\setminus(-a,a)^{d}}x_{i}^{2}p_{0}(x)\;\mathrm{d}x_{i}$
$\leq\frac{(1-\sigma_{\min})^{2}}{\sigma_{\min}^{2}}d\Big(\int_{\mathbb{R}\setminus(-a,a)}x^{2}\varphi(x)\;\mathrm{d}x\Big)^{d}.$
By assumption $\varphi$ decays faster than $e^{-x}.$ Using the upper incomplete $\Gamma$-function as defined in Lemma˜A.2, we obtain for $a>2$
$\int_{\mathbb{R}\setminus(-a,a)}\varphi(x)\;\mathrm{d}x$ $\leq\int_{\mathbb{R}\setminus(-a,a)}x^{2}\varphi(x)\;\mathrm{d}x$
$\leq\int_{\mathbb{R}\setminus(-a,a)}x^{2}e^{-x}\;\mathrm{d}x$
$=2\int_{a}^{\infty}x^{2}e^{-x}\;\mathrm{d}x$
$=2\Gamma(3,a)$
$\leq 6e^{-a}a^{2}.$
Hence
$\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus(-a,a)^{d}}|v^{n}_{t}(\psi^{n}_{t}(x))$ $-\tilde{v}_{t}(\psi^{n}_{t}(x))|^{2}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t\leq c\frac{a^{2d+2}e^{-da}}{\sigma_{\min}^{2}}.\qed$
Proof of Lemma˜5.2.
For the supremum norm bound, observe that
$|v^{n}_{t}(x)|$ $=\Big|\sum_{i=1}^{n}v_{t}(x|X_{i}^{*})\frac{p_{t}(x|X_{i}^{*})}{\sum_{i=1}^{n}p_{t}(x|X_{i}^{*})}\Big|$
$\leq\sum_{i=1}^{n}|v_{t}(x|X_{i}^{*})|\frac{p_{t}(x|X_{i}^{*})}{\sum_{i=1}^{n}p_{t}(x|X_{i}^{*})}$
$\leq\max_{i\in\{1,\dots,n\}}|v_{t}(x|X_{i}^{*})|$
$=\max_{i\in\{1,\dots,n\}}\frac{|(\sigma_{t}-t)X_{i}^{*}+x|}{1-(1-\sigma_{\min})t}$
$\leq\max_{i\in\{1,\dots,n\}}\frac{|X_{i}^{*}|+|x|}{1-(1-\sigma_{\min})t}$
$\leq\frac{\sqrt{d}(1+a)}{1-(1-\sigma_{\min})t}.$
For the Lipschitz bound, note that
$\nabla_{x}v_{t}^{n}(x)$ $=\nabla_{x}\sum_{i=1}^{n}\Big(\frac{\frac{\partial\sigma_{t}}{\partial t}}{\sigma_{t}}x-\mu_{t}(X_{i})+\frac{\partial\mu_{t}}{\partial t}(X_{i})\Big)\frac{p_{t}(x|X_{i})}{\sum_{j=1}^{n}p_{t}(x|X_{j})}$ (38)
$=\nabla_{x}\frac{-(1-\sigma_{\min})}{\sigma_{t}}x-\nabla_{x}\frac{-(1-\sigma_{\min})t}{\sigma_{t}}\frac{\sum_{i=1}^{n}X_{i}p_{t}(x|X_{i})}{\sum_{j=1}^{n}p_{t}(x|X_{j})}+\nabla_{x}\frac{\sum_{i=1}^{n}X_{i}p_{t}(x|X_{i})}{\sum_{j=1}^{n}p_{t}(x|X_{j})}.$
Now for the partial derivative of the $\ell$-st coordinate function with respect to $x_{k},$ $\ell,k\in\{1,\dots,d\},$ we get
$\frac{\partial}{\partial x_{k}}\frac{\sum_{i=1}^{n}X_{i,\ell}p_{t}(x|X_{i})}{\sum_{j=1}^{n}p_{t}(x|X_{j})}$
$=\frac{\big(\sum_{i=1}^{n}\big(-\frac{x_{k}-tX_{ik}}{\sigma_{t}^{2}}\big)X_{i,\ell}p_{t}(x|X_{i})\big)(\sum_{j=1}^{n}p_{t}(x|X_{j}))-\big(\sum_{i=1}^{n}X_{i,\ell}p_{t}(x|X_{i})\big)\big(\sum_{j=1}^{n}\big(-\frac{x_{k}-tX_{jk}}{\sigma_{t}^{2}}\big)p_{t}(x|X_{j})\big)}{\big(\sum_{j=1}^{n}p_{t}(x|X_{j})\big)^{2}}$
$=\frac{t}{\sigma_{t}^{2}}\Bigg(\frac{\sum_{i=1}^{n}X_{ik}X_{i\ell}p_{t}(x|X_{i})}{\sum_{j=1}^{n}p_{t}(x|X_{i})}-\frac{\big(\sum_{i=1}^{n}X_{ik}p_{t}(x|X_{i})\big)\big(\sum_{i=1}^{n}X_{i\ell}p_{t}(x|X_{i})\big)}{\big(\sum_{j=1}^{n}p_{t}(x|X_{i})\big)^{2}}\Bigg).$
Since $\operatorname{supp}(p^{*})\subset[-1,1]^{d},$ we can bound
$\frac{\partial}{\partial x_{k}}\frac{\sum_{i=1}^{n}X_{i,\ell}p_{t}(x|X_{i})}{\sum_{j=1}^{n}p_{t}(x|X_{j})}\leq\frac{2t}{\sigma^{2}_{t}}.$
Using $t\in[0,1],$ $\sigma_{\min}\leq 1,$ we get for (38)
$\nabla_{x}v_{t}^{n}(x)\leq\frac{1}{\sigma_{t}}I_{d}+\frac{2}{\sigma_{t}^{3}}J_{d},$
where $I_{d}$ denotes the $d\times d$ identity matrix, $J_{d}$ denotes the $d\times d$ matrix consisting of ones and $\leq$ denotes entry wise inequality. Using the mean value theorem, we obtain for $x,y\in\mathbb{R}^{d}$
$|v_{t}^{n}(x)-v_{t}^{n}(y)|$ $\leq\Big\|\frac{1}{\sigma_{t}}I_{d}+\frac{2}{\sigma_{t}^{3}}J_{d}\Big\|_{2}|x-y|.$
As
$\Big\|\frac{1}{\sigma_{t}}I_{d}+\frac{2}{\sigma_{t}^{3}}J_{d}\Big\|_{2}$ $\leq\Big\|\frac{1}{\sigma_{t}}I_{d}\Big\|_{2}+\Big\|\frac{2}{\sigma_{t}^{3}}J_{d}\Big\|_{2}$
$=\frac{1}{\sigma_{t}}+\frac{2d}{\sigma_{t}^{3}},$
we get the desired bound. ∎
Proof of Lemma˜5.3.
Due to the triangle inequality, the largest increase in Euclidean distance is achieved when two points move in exact opposite direction. Since $\mathcal{M}$ has reach $\tau$ and $r<\tau$, the extreme case of this happens when $x$ and $y$ satisfy $|x-y|=2(\tau-r)$ and $|\pi(x)-\pi(y)|=2\tau$. This situation is sketched in Figure˜4. Note that the manifold does not need to "make a turn", the same reasoning applies to the bottleneck situation, for which we refer to Aamari et al. (2019, Figure 3).
Figure 4: Largest proportional displacement, black line is part of manifold $\mathcal{M}$, gray solid line is boundary of the reach environment, gray dotted lines are the boundaries of the $r$-tube around $\mathcal{M}$. Black dots are $x$ and $y$ respectively, the arrow symbolizes the maximal displacement.
Thus for all $x,y\in U_{r}$
$|\pi(x)-\pi(y)|\leq\frac{1}{1-\frac{r}{\tau}}|x-y|.\qed$
Proof of Lemma˜5.4.
  1. (I)
    Differentiating $g^{-1}(g(y))=y$ leads to $Dg^{-1}(g(y))Dg(y)=I_{d^{\prime}}$, which implies that $Dg^{-1}(g(y))$ is the pseudo-leftinverse of $Dg(y)$. Using the construction of the pseudoinverse via the singular value decomposition, we conclude that
    $\|Dg^{-1}(g(y))\|=\frac{1}{\lambda_{\min}(Dg(y))}\leq\frac{1}{m}<1.$
    Hence
    $\|A(y)\|\leq\|Dg^{-1}(g(y))\|\|P_{T_{g(y)}\mathcal{M}}\|\leq\frac{1}{m}.$
  2. (II)
    For any $h\in\mathbb{R}^{d^{\prime}}$, differentiating $Dg^{-1}(g(y))Dg(y)=I_{d^{\prime}}$ in direction $h$ yields
    $D(Dg^{-1}(g(y)))[h]Dg(y)+Dg^{-1}(g(y))D(Dg(y))[h]=0,$ (39)
    where $D(Dg(y))[h]=\frac{d}{dt}Dg(y+th)\big|_{t=0}$ and $D(Dg^{-1}(g(y)))[h]=\frac{d}{dt}Dg^{-1}(g(y+th))\big|_{t=0}$.
    Fix $y\in\mathbb{R}^{d^{\prime}}$ and set $x=g(y)$. Differentiating $g(g^{-1}(x))=x$ for yields
    $Dg(g^{-1}(x))Dg^{-1}(x)=I_{T_{x}\mathcal{M}}\quad\Longleftrightarrow\quad Dg(y)Dg^{-1}(g(y))=I_{T_{g(y)}\mathcal{M}},$
    where $I_{T_{x}\mathcal{M}}$ is the identity map on the tangent space. Using this right inverse of $Dg(y)$ and recalling that $P_{T_{g(y)}\mathcal{M}}$ is the orthogonal projection onto the tangent space, we obtain for (39),
    $D(Dg^{-1}(g(y)))[h]\Big|_{T_{g(y)}\mathcal{M}}=-Dg^{-1}(g(y))D(Dg(y))[h]Dg^{-1}(g(y)).$
    Taking the supremum over $h$ such that $|h|=1$ leads to
    $\|D(Dg^{-1}(g(y)))\|\leq\|Dg^{-1}(g(y))\|^{2}\|D(Dg(y))\|\leq\frac{M_{2}}{m^{2}}$
    Combining Stewart and Sun (1990, Chapter II, Theorem 4.5, Corollary 4.6) with Boissonnat et al. (2019, Corollary 3) and the fact that for $x\in[0,\frac{\pi}{2}]$ $\sin(x)\leq 3\sin\big(\frac{x}{2}\big)$ yields
    $\|P_{T_{g(y+th)}\mathcal{M}}-P_{T_{g(y)}\mathcal{M}}\|\leq\frac{3}{2\tau}|g(y+th)-g(y)|.$
    Furthermore,
    $\|D(P_{T_{g(y)}\mathcal{M}})[h]\|=\lim_{t\rightarrow 0}\frac{\|P_{T_{g(y+th)}\mathcal{M}}-P_{T_{g(y)}\mathcal{M}}\|}{|t|}\leq\lim_{t\rightarrow 0}\frac{3}{2\tau}\frac{|g(y+th)-g(y)|}{|t|}=\frac{3}{2\tau}|Dg(y)[h]|$
    Next, we differentiate for $h\in\mathbb{R}^{d^{\prime}}$
    $DA(y)[h]=D(Dg^{-1}(g(y)))[h]P_{T_{g(y)}\mathcal{M}}+Dg^{-1}(g(y))D(P_{T_{g(y)}\mathcal{M}})[h].$
    Recall that the operator norm is defined as $\sup_{|h|=1}\|DA(y)[h]\|=\sup_{|h|=1}\sup_{|v|=1}|DA(y)[h]v|$. Then, inserting all constants from ˜3.9
    $\|DA(y)\|\leq\|D(Dg^{-1}(g(y)))\|\|P_{T_{g(y)}\mathcal{M}}\|+\|Dg^{-1}(g(y))\|\|D(P_{T_{g(y)}\mathcal{M}})\|\leq\frac{M_{2}}{m^{2}}+\frac{3M_{1}}{2m\tau}=c_{0}<\frac{1}{r}.$
  3. (III)
    Assume that $F_{u}(y_{1})=F_{u}(y_{2})$. Then
    $0=F_{u}(y_{1})-F_{u}(y_{2})=y_{1}-y_{2}+(A(y_{1})-A(y_{2}))u.$
    If $y_{1}\neq y_{2}$
    $|y_{1}-y_{2}|=|(A(y_{1})-A(y_{2}))u|\leq|A(y_{1})-A(y_{2})||u|\leq c_{0}r|y_{1}-y_{2}|<|y_{1}-y_{2}|.$
    The above inequality is false, which implies $y_{1}=y_{2}$.
  4. (IV)
    Rearranging yields
    $F_{u}(S_{u}(z))=z\quad\Longleftrightarrow\quad S_{u}(z)+A(S_{u}(z))u=z\quad\Longleftrightarrow\quad S_{u}(z)=z-A(S_{u}(z)).$
  5. (V)
    In order to use the inverse function theorem, we first need to ensure that $DF_{u}$ has a nonzero determinant. For $h\in\mathbb{R}^{d^{\prime}}$, define
    $DF_{u}(y)[h]=h+(DA(y)[h])u.$
    Set the linear operator $G(y,u)\colon\mathbb{R}^{d^{\prime}}\rightarrow\mathbb{R}^{d^{\prime}}$ with $G(y,u)[h]=(DA(y)[h])u$, then
    $DF_{u}(y)=I_{d^{\prime}}+G(y,u),\quad\text{and}\quad\|G(y,u)\|\leq\|DA(y)\||u|\leq c_{0}r<1.$
    For the $i$-th largest singular value $s_{i}$, using Weyl’s singular value perturbation bound, for an English reference see e.g. Stewart (1990),
    $|s_{i}(DF_{u}(y))-s_{i}(I_{d^{\prime}})|\leq\|G(y,u)\|\leq\|DA(y)\||u|\leq c_{0}r<1.$
    Since the absolute value of the determinant is the product of the singular values, we obtain
    $(1-c_{0}r)^{d^{\prime}}\leq|\operatorname{det}(DF_{u}(y))|\leq(1+c_{0}r)^{d^{\prime}}.$
    Moreover, $t\mapsto\operatorname{det}\left(I_{d^{\prime}}+tG(y,u)\right)$ is continuous, unequal to 0 and equals 1 at $t=0$. Hence, $\operatorname{det}\left(DF_{u}(y)\right)>0$. Since $\|DA(y)\||u|\leq c_{0}r<1$, a Neumann series yields
    $\|(DF_{u}(y))^{-1}\|=\|(I_{d^{\prime}}-(-G(y,u)))^{-1}\|\leq(1-\|G(y,u)\|)^{-1}\leq\frac{1}{1-c_{0}r}$
    Since $F_{u}$ is injective, we have by the inverse function theorem that
    $DS_{u}(y)=(DF_{u}(S_{u}(y)))^{-1},\quad\text{ thus}\quad\|DS_{u}(y)\|\leq\frac{1}{1-c_{0}r}.$
  6. (VI)
    First, we differentiate $S_{u}(z)=z-A(S_{u}(z))u$ to obtain for $h\in\mathbb{R}^{d^{\prime}}$
    $DS_{u}(z)[h]=h-DA(S_{u}(z))[DS_{u}(z)[h]]u.$ (40)
    Defining the operator $B(z,u)\coloneqq(DA(S_{u}(z)))[\cdot]u$ for $v\in\mathbb{R}^{d^{\prime}}$ and rearranging yields on a matrix level
    $(I_{d^{\prime}}+B(z,u))DS_{u}(z)=I_{d^{\prime}}\quad\Longleftrightarrow\quad DS_{u}(z)=(I_{d^{\prime}}+B(z,u))^{-1},$
    where the invertibility follows from
    $\|B(z,u)\|=\sup_{|v|=1}\|(DA(S_{u}(z))[v])u\|\leq c_{0}|u|<1.$
    The inequality follows from property (II). Thus
    $\operatorname{det}(DS_{u}(z))=(\operatorname{det}(I+B(z,u)))^{-1}.$
    By the same singular value perturbation bound that was used in (V), we obtain
    $|s_{i}(I+B(z,u))-s_{i}(I)|\leq\|B(z,u)\|\leq c_{0}|u|.$
    This implies
    $(1-c_{0}|u|)^{d^{\prime}}\leq|\operatorname{det}(I+B(z,u))|\leq(1+c_{0}|u|)^{d^{\prime}},$
    which yields
    $(1+c_{0}|u|)^{-d^{\prime}}-1\leq|\operatorname{det}(DS_{u}(z))|-1\leq(1-c_{0}|u|)^{-d^{\prime}}-1.$
    We obtain
    $||\operatorname{det}(S_{u}(z))|-1|\leq\max\big(1-(1-c_{0}|u|)^{-d^{\prime}},(1+c_{0}|u|)^{-d^{\prime}}-1\big).$
    Two Taylor expansions yield for $\xi\in[0,c_{0}|u|]\subset[0,c_{0}r]$
    $|(1+\xi)^{-d^{\prime}}-1|$ $\leq\sup_{\zeta\in[0,c_{0}r]}d^{\prime}(1+\zeta)^{-d^{\prime}-1}\xi\leq d^{\prime}\xi,$
    $|(1-\xi)^{-d^{\prime}}-1|$ $\leq\sup_{\zeta\in[0,c_{0}r]}d^{\prime}(1-\zeta)^{-d^{\prime}-1}\xi\leq d^{\prime}(1-c_{0}r)^{d^{\prime}-1}\xi,$
    which implies
    $||\operatorname{det}(S_{u}(z))|-1|\leq d^{\prime}c_{0}(1-c_{0}r)^{d^{\prime}-1}|u|.$
Since $\sup_{y}\|DA(y)\||u|<1$, $F_{u}$ is surjective. Additionally, we showed that $S_{u}$ exists and from (40) and the properties of $g$ we can conclude that the derivative of $S_{u}$ is continuous, we obtain that $F_{u}$ is a diffeomorphism. ∎
Appendix A Appendix
Theorem A.1.
Assume $p^{*}\in B_{1,\infty}^{\alpha}(\Lambda,\mathbb{R}^{d}),\alpha\in(0,1],\Lambda>0$. Further assume $K$ is a non negative, $d$-dimensional kernel such that $\int K_{\sigma_{\min}}(y)y\;\mathrm{d}y=0$,
$\int\max(|u|^{2},|u|^{d+4})K^{2}(u)\;\mathrm{d}u<\infty$ $\quad\text{and}\quad\int|x|^{d+4}p^{*}(x)\;\mathrm{d}x<\infty.$
If $\sigma_{\min}\leq 1$, then there are $C_{1}$ and $C_{2}$ such that for any $\tilde{v}\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ and $v$ from (9),
$\mathbb{E}\big[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})\big]\leq C_{1}\sigma_{\min}^{1+\alpha}+\frac{C_{2}}{\sqrt{n\sigma_{\min}^{d}}}.$
Proof.
As in Theorem˜3.4,
$\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})$ $=\sup_{f\in\mathrm{Lip}(1)}\int f(x)(p^{*}(z)-(K_{\sigma_{\min}}\ast\mu_{n})(z))\;\mathrm{d}z$
$\leq\sup_{f\in\mathrm{Lip}(1)}\int f(z)(p^{*}(z)-(K_{\sigma_{\min}}\ast p^{*})(z))\;\mathrm{d}z$
$\quad+\sup_{f\in\mathrm{Lip}(1)}\int f(z)((K_{\sigma_{\min}}\ast p^{*})(z)-(K_{\sigma_{\min}}\ast\mu_{n})(z))\;\mathrm{d}z.$ (41)
For the first term, we obtain as in Theorem˜3.4
$\sup_{f\in\mathrm{Lip}(1)}\int f(z)(p^{*}(z)-(K_{\sigma_{\min}}\ast p^{*})(z))\;\mathrm{d}z\leq C_{1}\sigma_{\min}^{1+\alpha}.$
Second term: To bound the expectation of (41) we first use that by (Villani, 2008, Theorem 6.15)
$\sup_{f\in\mathrm{Lip}(1)}\int f(z)((K_{\sigma_{\min}}\ast p^{*})(z)-(K_{\sigma_{\min}}\ast\mu_{n})(z))\;\mathrm{d}z\leq\int|z||(K_{\sigma_{\min}}\ast p^{*})(z)-(K_{\sigma_{\min}}\ast\mu_{n})(z)|\;\mathrm{d}z.$
Then for $\rho>\frac{d}{2}$
$\mathbb{E}[\int$ $|z||(K_{\sigma_{\min}}\ast p^{*})(z)-(K_{\sigma_{\min}}\ast\mu_{n})(z)|\;\mathrm{d}z]$
$=\int(1+|z|)^{-\rho}|z|(1+|z|)^{\rho}\mathbb{E}\Big[\Big|\frac{1}{n}\sum_{i=1}^{n}K_{\sigma_{\min}}(z-X_{i})-\mathbb{E}[K_{\sigma_{\min}}(z-X_{i})]\Big|\Big]\;\mathrm{d}z$
$\leq\frac{1}{n}\sqrt{\int(1+|z|)^{-2\rho}}\;\mathrm{d}z\sqrt{\int|z|^{2}(1+|z|)^{2\rho}\mathbb{E}\Big[\Big(\sum_{i=1}^{n}K_{\sigma_{\min}}(z-X_{i})-\mathbb{E}[K_{\sigma_{\min}}(z-X_{i})]\Big)^{2}\Big]\;\mathrm{d}z},$ (42)
where the equality holds by Fubini’s theorem and the inequality holds by the Jensen inequality together with the Cauchy-Schwarz inequality.
Next we bound the expectation. Since $\mathbb{E}[K_{\sigma_{\min}}(z-X_{i})-\mathbb{E}[K_{\sigma_{\min}}(z-X_{i})]]=0$ and $X_{1},\dots,X_{n}$ are i.i.d. we have that
$\mathbb{E}\Big[\Big(\sum_{i=1}^{n}K_{\sigma_{\min}}(z-X_{i})-\mathbb{E}[K_{\sigma_{\min}}(z-X_{i})]\Big)^{2}\Big]$ $=n\operatorname{Var}\big(K_{\sigma_{\min}}(z-X_{1})\big)$
$\leq n\mathbb{E}[(K_{\sigma_{\min}}(z-X_{1}))^{2}].$
Therefore
$\int|z|^{2}(1+|z|)^{2\rho}$ $\mathbb{E}\Big[\Big(\sum_{i=1}^{n}K_{\sigma_{\min}}(z-X_{i})-\mathbb{E}[K_{\sigma_{\min}}(z-X_{i})]\Big)^{2}\Big]\;\mathrm{d}z$
$=n\int\mathbb{E}\Big[|X_{1}+\sigma_{\min}y|^{2}\big(1+|X_{1}+\sigma_{\min}y|^{2\rho}\big)^{2}K^{2}(y)\Big]\;\mathrm{d}y$
$\lesssim\frac{n}{\sigma_{\min}^{d}}\int\mathbb{E}\big[|X_{1}+\sigma_{\min}y|^{2}+|X_{1}+\sigma_{\min}y|^{2\rho+2}\big]K^{2}(y)\;\mathrm{d}y$
$\lesssim\frac{n}{\sigma_{\min}^{d}}\int\big(\mathbb{E}\big[|X_{1}|^{2\rho+2}\big]+\sigma_{\min}^{2}|y|^{2}+\sigma_{\min}^{2\rho+2}|y|^{2\rho+2}\big)K^{2}(y)\;\mathrm{d}y$
$\lesssim\frac{n}{\sigma_{\min}^{d}}(1+\sigma^{2}_{\min}+\sigma_{\min}^{2\rho+2})\int\max(|y|,|y|^{2\rho+2})K^{2}(y)\;\mathrm{d}y.$
Using that $\sigma_{\min}\leq 1$ we can bound (42) for $\rho>\frac{d}{2}$ further by
$\mathbb{E}[\int|z||(K_{\sigma_{\min}}\ast p^{*})(z)-(K_{\sigma_{\min}}\ast\mu_{n})(z)|\;\mathrm{d}z]$ $\lesssim\frac{1}{\sqrt{n\sigma_{\min}^{d}}}\int\max(|y|,|y|^{2\rho+2})K^{2}(y)\;\mathrm{d}y.$
For $p=\frac{d}{2}+1$ the last integral is finite and we thus obtain
$\mathbb{E}[\int|z||(K_{\sigma_{\min}}\ast p^{*})(z)$ $-(K_{\sigma_{\min}}\ast\mu_{n})(z)|\;\mathrm{d}z]\leq\frac{C_{2}}{\sqrt{n\sigma_{\min}^{d}}}.$
In the end, we get for (41) that
$\mathbb{E}\big[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})\big]\leq C_{1}\sigma_{\min}^{1+\alpha}+\frac{C_{2}\max(1,{\sigma_{\min}},\sigma_{\min}^{p+1})}{\sqrt{n\sigma_{\min}^{d}}}.\qed$
For comprehensiveness, we translate the following result from German including the proof.
Lemma A.2.
(Gabcke, 1979, Satz 4.4.3) Let $a\geq 1$ and $x>a.$ The upper incomplete gamma function
$\Gamma(a,x):=\int_{x}^{\infty}e^{-v}v^{a-1}\;\mathrm{d}v$ (43)
can be bounded by
$\Gamma(a,x)\leq ae^{-x}x^{a-1}.$
Proof.
Let $a\geq 0$. Substituting $v=1/w$ in (43) leads to
$\Gamma(a,x)=\int_{0}^{1/x}e^{-1/w}w^{-a-1}dw.$
Set
$h(w):=e^{-1/w}w^{-a-1},$
the first derivative of $h$ is
$h^{\prime}(w)=e^{-1/w}w^{-a-2}\left[\frac{1}{w}-(a+1)\right].$
If $x>a+1$, then $h^{\prime}(w)>0$ for all $w\in[0,\frac{1}{x}]$. Hence $h$ is strictly increasing and
$\Gamma(a,x)\leq\frac{1}{x}h\left(\frac{1}{x}\right)=e^{-x}x^{a}\quad(a\geq 0,x>a+1).$ (44)
For $a\geq 1$ integration by parts in (43) leads to
$\Gamma(a,x)=e^{-x}x^{a-1}+(a-1)\Gamma(a-1,x).$
Using (44) with a1a-1 instead of $a,$ we get
$\Gamma(a,x)\leq e^{-x}x^{a-1}+(a-1)e^{-x}x^{a-1}=ae^{-x}x^{a-1}\quad(a\geq 1,x>a).\qed$
Numerical illustration with smaller networks
(a) Mean distance of the generated samples to the manifold.
(b) Largest gap on the manifold between the projections of the generated samples.
(c) Mean distance of the generated samples to the manifold.
(d) Largest gap on the manifold between the projections of the generated samples.
Figure 5: Comparison between the Flow Matching estimator and the KDE in terms of distance to the manifold and the largest gap of the projection estimator on the manifold. Top row uses a three hidden layer SeLU net of width $32$ and bottom row uses a three hidden layer SeLU net of width 88.